A Bibliography of Publications about High-Performance Fortran

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

10 August 2023 Version 2.34

Title word cross-reference

+ [BMV03]. $24.95 [Eme94]. 2^{k} [AEG^{+}02]. 2^{k−1} [AEG^{+}02]. _{H}P F [OP99].

0-1 [BKK94]. 0-262-61094-9 [Eme94].

1.1 [BDPW98, OA02]. 100 [Str94]. 11 [Sul88]. 14.9 [SMSY02]. 186th [Ano95a]. 1992 [Ano92b]. 1993 [Hig94a]. 1994 [IEE94e]. 1st [Ano94b, Fer92, Kum94, Pra95].

2 [LC97, Sch97]. 2.0 [KK01]. 2/M [FK95]. 2000 [Ano00b]. 2003 [ACM03]. 2007 [ACM07]. 25-27 [Ano94l]. 25-28 [Ano93j]. 25th [Ano94h]. 2nd [DT94, FK95].

3 [KKS^{+}95]. 3rd [IEE96a].

4th [CKMU94, Iwa00].

5 [KBKT94]. 5th [Ban93, Fri94, IEE94a].

6th [BGNP94, HMPT94].

77 [MCL^{+}95]. 77toHPF [Van94b]. 7ET [Eme94]. 7th [PBG^{+}95].

860 [KR94, KR95]. 8th [Hua96].

9-10 [ACM07]. 90

[Ano92b, Ano93b, Ano94k, Ano96, BGvE^{+}97, Cou97, DL97c, EEV^{+}96, FSPC^{+}02, Met99a, Met99b, Met99c, Met99d, Met00c, Met00a, Met00b, Met01b, Met01c, Met01a, Met02a, Met02b, MCL^{+}95]. 90/95/HPF [Met99a, Met99b, Met99c, Met99d, Met00c, Met00a, Met00b, Met01b, Met01c, Met01a, Met02a, Met02b]. 90/HPF [FSPC^{+}02]. 90D [BCF^{+}94a, Ano94f, BCFH93, BCF^{+}93a, BCF^{+}93b, BCF^{+}93c, BCF^{+}94c, BCF^{+}94b, BCF^{+}94d, PHHF94a, PHHF94b, PHHF95, PH96a, PH96b, Pon94a, Pon94b]. 90D/High [BCF^{+}94a]. 90D/HPF [BCFH93, BCF^{+}93b, BCF^{+}93c, BCF^{+}94c, BCF^{+}94b, BCF^{+}94d, Pon94a, Pon94b]. ’92 [IEE92b]. ’93 [Ano93l, GGK^{+}93, IEE93a, IEE93c]. ’94 [BLT94, BGG^{+}94, CGS94, Fri94, HMPT94, IEE94b, IEE94e]. 94-VAPP [BV94]. ’95 [HAM95, Hua96, IEE95a, Ano93k, Ano94k, BGvE^{+}97, Cou97, MCL^{+}95, vWAH^{+}02]. ’96 [ACM96a, ACM96b, BDLS96]. 9th [ACM95, IEE95a].

abstracts [Sch93]. Academic [BGvE^{+}97]. accelerating [SIOS02]. Access [KNS95b, LP93, BxCW01, KNS95a]. accesses [DSvH94]. achieved [DPR94]. ACM [ACM97, ACM07, IEE02, Ano95b, EEV^{+}96, Kar95]. ACM/IEEE [ACM97, Kar95, ACM98]. ACPC [Vol93]. Activities [Ano99, Ano00a, Ano00b, Ano01a, Ano01b, Ano01c, Ano02b, Ano02a, Ano03]. Activity [Ano00b]. Adaptive [KK94, KK95b, AES^{+}96, CC94, HMS^{+}95, SPM^{+}94]. Adaptor [BZ94]. add [AEG^{+}02]. Address [SSC00, TR96]. Addresses [CGL^{+}95, CGL^{+}93]. ADDT [SR96]. Advance [EEV^{+}96]. Advanced [AMC01, Ben98, CZM94, CZVM94, Don95, MKF95, MCAB^{+}02, BLT94, BP97, Ben99b, BLZ99, CMZ94b, CMZ94a, FSPC^{+}02, CMZ95]. Advances [Nic91]. Aeroacoustic [NOL97]. aerospace [MZ00, MZ01]. Affine [SSC00]. age [HK95]. Airshed [SS00]. Alexandria [Ano94b]. algebra [ACIK97]. Algorithm [IK96, KNS95b, ADH95, KNS95a, SZG95]. Algorithms [Din98, HHK94, TCR96, BID95, Din99, FP92]. Align [HCLJ03]. Alignment [HCLJ03, CZM93, CMZ93, WI94]. alignments [vHKS94]. Alive [MCL^{+}95]. Allocatable [BGvE^{+}97]. alternative [CZM93, CMZ93]. amplifier [MIN^{+}95]. Analysis [KOM94, SNK06, SVD96, SDv98, DSvH94, GS97, KW94, KOM93, LPA95, SSG94]. ’Annai [CEF^{+}95]. Annapolis [IEE96b]. Anniversary [Ano93j]. Annual [Iwa00]. AP1000 [HDH^{+}94, HDH^{+}95, SIDH95]. Application [Fox94, GLPE97, PHHF94a, PHHF95, YFH97, PSG03]. Applications [ASS95, BGvE^{+}97, Ben98, BSSV98, Ben99a, Bra94a, BCC^{+}96a, BCC^{+}96b, CNBB96, CZM94, CZVM94, FGRT00, Fer92, FK95, Jou95, SN94, SN95, ASS93, BLT94, BP97, Ben99b, BLZ99, Ben00, Bra94b, BCC^{+}97b, BCC^{+}97a, BCC^{+}97c, BxCW01, BMV03, CMZ94b, CMZ94a, CMVZ94, CMZ95, DDcMR96, DSZ94, Don95, MM94, MZ00, MZ01, PD96, SM02a, SIOS02]. Applied [Lev94]. Approach [ASS95, BCFH93, CMMP98, BCF^{+}94c, KHJS94, KJEM12, Wag94, WW94]. Approaches [SM02b]. APR [Wag94]. April [CKMU94, DR94, Fri94, GH94a, GH94b, IEE95a, Sie94a, Sie94b, Ano96]. archetypal [HKM98]. Architectural [Ano94a, HDH^{+}94, HDH^{+}95]. Architecture [Ano93a, CMVZ94]. Architectures [HHK94, BZ99, CGS94, HMPT94, Sab95, ZCP95]. Arguments [MCL^{+}95]. Arles [Van95]. Arlington [IEE92a]. Array [BGvE^{+}97, BBZ94, HM96, HM98, HLJ01, KHS96, Ste95, SOG94, TCF94, TCR96, AW94, BBDR94, BBDR95, KHS95, KHRS95, KW95, LPA95, PQ94, WW94, WW95, WI94]. Arrays [BGvE^{+}97, vDSP96, BSCV95, DSvH94]. Articles [MCL^{+}95]. Ascona [DR94]. ASL [FGRT00]. aspects [Per94]. ASPLOS [Ano94a]. ASPLOS-VI [Ano94a]. Assignment [Ste95]. assimilation [HBD^{+}93]. Association [BGvE^{+}97, Hig94d]. Athens [HMPT94]. Atmospheric [HK93, PFS^{+}04]. Attribute [BGvE^{+}97]. August [Agr95, Ano00b, Ban93, BGNP94, CGS94, HAM95, Hua96, PBG^{+}95, Van95]. Austin [IEE94b]. Austria [BV94, Vol93]. Austrian [Fer92, FK95]. Austrian-Hungarian [Fer92, FK95]. Automatic [BKK94, BB96, CGSS94, CMKH03, CP94, FGL01, KK95a, KK98, NVG94, SR96, WI94, AGG^{+}97, GKH^{+}93]. Autonomous [NJ94]. Available [MCL^{+}95]. award [Str94]. Awards [Str94].

B [Eme94, UMM94]. Bad [BGvE^{+}97]. balanced [vHK00]. Baltimore [IEE02]. Bangalore [Kum94, Pra95]. Barbara [Ano95b, IEE95a]. Barcelona [ACM95]. Based [DCBC98, GLPE97, Ger98a, Ger98b, HZ94, MKS94, NOL97, OP98b, OP98c, SSG94, vHKS94, ZCFL98]. basic [SZG95]. Basis [TR96]. be [DPR94, MCL^{+}95]. beam [MKF95, QRH00]. Benchmark [SF02, DS02, HJJ^{+}00]. Benchmarks [AHOK02, MMY95b, CDD^{+}96, MMY95a]. Better [BBCR98, Str94]. Betty [BGvE^{+}97]. between [BID95]. big [Str94]. Binding [DCBC98, DCR99a]. Black [Nel96]. Blanks [BGvE^{+}97]. Block [ASS95, Van98, ASS93, HKMCS94, HC08, SZG95, VRT97, WO96]. block-cyclic [HKMCS94, HC08, VRT97, WO96]. block-sparse [SZG95]. body [MB95]. Book [Eme94, Rag95, UMM94]. Books [Ano92a, Met99d, Met02b, Met01b]. Boulder [Sch93]. breast [Str94]. Brian [BGvE^{+}97]. Brinch [Off98]. Bringing [FKKC96]. Budapest [FK95]. buffers [MR96]. bugs [DS02]. Builder [DL97c, Lef98].

C [BBG^{+}93, CC94, Chw97, EEV^{+}96, YGS^{+}94]. CA [Ano95b, BBG^{+}95, IEE95a, Kar95]. calculate [SHZ13]. California [ACM93, ACM97, ACM07, Ano94a, IEE93a]. call [YO95]. can [DPR94]. Canada [BGG^{+}94, CGS94, GGK^{+}93]. cancer [Str94]. Cancun [Sie94a, Sie94b]. Canonical [BV98]. Captured [BGvE^{+}97, EEV^{+}96, MCL^{+}95]. Carlo [MMV95]. cars [Str94]. CASCON [BGG^{+}94, GGK^{+}93]. Case [BF01, GLPE97, GS01, SS97, SN94, SN95, Bri00, DS97]. CC [Fri94]. Cell [ADHF96, CLiN^{+}02]. Cenju [KKS^{+}95]. Cenju-3 [KKS^{+}95]. Center [ACM98, IEE94b, Kar95, Wie94]. CERN [Van95]. Cetraro [Don95]. CG [SZG95]. CG-algorithm [SZG95]. Challenge [Sai95]. Challenges [Agr95, LL98, Ten93, BCM^{+}93]. Change [Mei94]. Chapter [AMC01, KK01]. Character [Cou97]. Charles [Eme94, Rag95, UMM94]. checks [KW95]. Chenies [Eme94]. Cholesky [LC97]. class [PQ94]. Client [Ano93j]. Client/Server [Ano93j]. climate [Str94]. Closed [TR96]. Cluster [CLiN^{+}02, BID95, SR95]. Clusters [Del98, BMV03, KJEM12]. CM [KBKT94, MM94, NOL97, Sai95]. CM-5 [KBKT94]. co [OP99, Ano94h]. co-ordination [OP99]. Coarse [BR96]. Coarse-Grained [BR96]. Code [AMC01, CLiN^{+}02, DCBC98, DCR99b, MB95, vWAH^{+}02, ACIK97, Eli98, Ogi02, SHZ13, vHK00]. Codes [Adv98, ADHF96, DL97c, Nel96, BSCV95, Sai95, Sta94, UZCZ96]. Colorado [Ano94h, Sch93]. COLT} [OP99]. Columbus [Hua96]. combined [MIN^{+}95]. Coming [HK95]. comments [Str94]. Committees [Ano99, Ano00a, Ano00b, Ano01a, Ano01b, Ano01c, Ano02b, Ano02a, Ano03, Ano93k, Ano94j, Ano94k]. Communicating [FKK^{+}96b, FKK96a]. Communication [BR98, BD96, CL97b, CGL^{+}95, KW94, MR96, PSC93b, SOG94, TRV96, VRT97, BBDR94, BBDR95, CGL^{+}93, GKH^{+}92, GB94, KHJS94]. Communication-buffers [MR96]. communication-efficient [KHJS94]. communication/computation [BBDR94, BBDR95]. communications [Ano94b, Coe94a]. comp [EEV^{+}96]. comp-fortran-90 [EEV^{+}96]. Comparing [BF01]. comparison [BID95, GS95, HKM98, SM02a]. Compatibility [SM02a]. Compcon [IEE93a]. Compilation [Adv98, BCFH93, CMMP98, Coe94a, Coe94b, CA96, FXAC94, HHKT96, HKMCS94, PSC93b, TBC94a, UZCZ97, ZCFL98, BCF^{+}94c, CGS94, MCH96, PSC^{+}95]. Compile [ASS95, GB94, PH96a, PH96b, SPM^{+}94]. Compile-Time [ASS95, PH96a, PH96b, GB94, SPM^{+}94]. Compiler [ASS93, BBZ94, BSSV98, BCF^{+}93c, BMN^{+}97, BD96, Bra00, Fri94, GMS^{+}95, HKT93a, HKT93b, IHKvW02, Ken94a, Nel96, SIDH95, SS97, TBC94b, ZCP95, AFMP95, ABC^{+}96, BCM^{+}93, BCF^{+}93a, BCF^{+}93b, BSCV95, CMT01, Chw97, DS97, HDH^{+}94, KKS^{+}95, Lov94, MCAB^{+}02, NVG94, SNK06, Spo94, WCC99]. Compilers [Ano93i, BB96, BCFH93, IK96, LZ97, Nak95b, SF02, Sch93, SS96, Ban93, BGNP94, BCF^{+}94c, DPR94, HDH^{+}95, Hua96, KW94, Met99b, Met00a, Met01a, Nak95a, Nic91, PBG^{+}95, Pon94a, Pon94b, SM02a, SNMC93]. Compiling [BZ99, BCF^{+}93a, BCF^{+}94b, BCF^{+}94d, BMMN94, BMN^{+}95, BMMN95, HBB^{+}95, HKT92, KHS96, RMCKB97, HMS^{+}95]. Compositional [KR94, KR95]. Computation [BD96, IEE94a, IEE96b, JB01, Sch96a, TR96, BBDR94, BBDR95, Eme94, HKM98, Vol93].

Computational

[BFHH94, BLW02, HF95, PSG03, Str94]. Computations [Ano94l, Bra00, MR95, CC94, KB94, MR96, PDS^{+}93, UZCZ95]. Computer [Ten93, vDSP96, Str94]. Computers [BCF^{+}93c, SS96, BCF^{+}93b, BCF^{+}94b, BCF^{+}94d, Duv92, Ger98a, Ger98b, KKS^{+}95, LP93, Wie94]. Computing [ACM97, ACM98, Ano93i, BBG^{+}95, Dow93, IEE94c, IEE96a, KNS95b, Kon00, LMR^{+}97, Ten93, Van95, Ano93l, Ano94b, Ban93, BGNP94, Bou95, Cel96, CDF^{+}93, DT94, Don95, GH94a, GH94b, HS95, Hua96, KNS95a, PBG^{+}95, Sab95]. Concepts [Ano93f]. Concerns [Off98]. Concurrent [BGMZ92, Bre92]. Conference [ACM94, ACM96a, ACM96b, ACM97, ACM98, ACM07, Ano94a, Ano94h, BBG^{+}95, BV94, CGS94, DSZ94, ERS95, ERS96, Fri94, GH94a, GH94b, HMPT94, HAM95, HS95, HS94, IEE94c, IEE94d, IEE95b, IEE96a, IEE96c, IEE02, Kar95, Vol93, ACM95, Ano93l, BLT94, BDLS96, CKMU94, DR94]. conformal [MKF95]. Confused [MCL^{+}95]. Connecticut [Ban93]. CONPAR [BV94]. Constructing [Ano93a]. construction [Fri94]. Constructor [BGvE^{+}97]. constructs [ABC^{+}96]. containing [BSCV95]. Contribution [BBCR98]. Controlled [NJ94]. Convention [ACM98, IEE94b, Kar95]. cooperation [Str94]. Coordinate [OP98b, OP98c, DRST03]. Coordination [OP98a, FKK96a]. Copies [Ano92b, Ano93k]. copy [GS97]. Core [TBC94b, TBC94a]. Corporation [AOL94a]. Cosmological [MB95]. costs [GB94, GS97]. County [ACM98]. Courses [Met01c, Met00c]. CPAR [CC94]. CPAR-language [CC94]. Craft [Chw97, SZG95, WCC99]. Craft-y [Chw97]. CRAY [PSG03, SZG95]. CRPC [Fox91]. CRPC-TR92225 [Fox91]. Curses [Ano93a]. cyclic [HKMCS94, HC08, VRT97, WO96].

D [KR95, HHKT96, HKT92, HKT93a, HKT93b, HKTW94, Ken94b, KR94, Wie94, vHKS94]. Dallas [Ano94g, IEE93b]. Dame [IEE96c]. DAQV [HM98]. d’Arc [BLT94]. DaReL [KN95a]. Data [BCC^{+}97b, Bra00, CFK^{+}94, CZM94, CGL^{+}95, Fox94, Guo01, HCLJ03, KK95a, KNS95b, KK98, Meh93a, Meh93b, MBFC98, PSC93b, PHD^{+}95, RSB97, SSC00, SR96, SR04, Ste95, TR96, UZCZ96, AW94, AFMP95, ABC^{+}96, AGG^{+}97, BKK94, BID95, BxCW01, CMZ94b, CMVZ94, CGL^{+}93, CP94, FKK^{+}96b, GKH^{+}92, GKH^{+}93, GHSJ94, GS97, GS95, HBD^{+}93, HC08, KN95a, KN95b, KHJS94, KNS95a, KGV97, Mar93, MBFC99, MR96, Off94, OPP00, PPW94, Per94, PD96, PSC^{+}95, SNK06, SV95, TBC94a, UZCZ95, WO96, YO95, BCC^{+}97c, BGvE^{+}97]. Data-Parallel [CZM94, CGL^{+}95, Guo01, KNS95b, PHD^{+}95, SSC00, Ste95, UZCZ96, AFMP95, CGL^{+}93, FKK^{+}96b, GS97, KNS95a, MR96, UZCZ95]. data-parallelism [PPW94]. Data-structure [BCC^{+}97b, BCC^{+}97c]. date [Din99]. David [Eme94, Rag95, UMM94]. DC [IEE94e]. DDT [AGG^{+}97]. Debugger [CH94, LL98, FSPC^{+}02]. debugging [SSG94, SSG97]. DEC [EEV^{+}96, Lov94]. December [HHK94, IEE92a, IEE93b, IEE96a, Kar95, Kum94, Pra95]. Decks [NOL97]. Decomposition [GLPE97, NVG94]. DECUS [EEV^{+}96]. Delaunay [CCW04]. Delft [DSZ94]. Denotational [Guo01]. densities [SHZ13]. dependent [SHZ13]. Deplore [EEV^{+}96]. descriptions [MKF95]. Design [BLLWW95, BDPW98, BCF^{+}93b, BCF^{+}93c, Coe96, PHHF94a, PHHF95, SOG94, ISKvW02, QRH00]. Designing [LL98]. Details [Cou97]. detect [Str94]. detection [CFMR95]. deterministic [CFMR95]. Developer [IEE96c]. developers [Str94]. developing [CDF^{+}93]. Development [Ano93g, NJ94, PHHF94a, PHHF95, DSZ94, KKMP95, Wie94]. dHPF [MCAB^{+}02]. Diego [ACM93, ACM07, Kar95]. difference [Str94]. Digest [IEE93a]. Digital [AOL94a, BLLWW95]. Dimensional [CLiN^{+}02, BSCV95, Eli98, Ogi02, SMSY02]. Direct [MMV95]. Directions [Bod94]. directives [BCF^{+}93a]. Discussing [Coe96]. Distributed [AW94, BR96, BCF^{+}93c, BMMN94, BMMN95, BMN^{+}97, CL97a, CL97b, CH94, HM96, HBB^{+}95, HLJ01, IEE92a, IEE93b, KHS96, KK98, RSB97, Sch93, vDSP96, BZ99, BCF^{+}93b, BCF^{+}94b, BCF^{+}94d, CEF^{+}95, DSvH94, DR94, GHSJ94, HKT92, HMS^{+}95, KN95a, KMR^{+}97, KHS95, KGV97, PZA93, SNK06, TBC94a, Wag94, WW95, WI94, YO95, ZA93, HM98]. Distributed-Memory [KHS96, KK98, HBB^{+}95, BZ99, GHSJ94, HKT92, HMS^{+}95, KN95a, KHS95, SNK06]. Distribution [CGSS94, SR96, ACIK97, AGG^{+}97, BCF^{+}93a, CZM93, CMZ93, CP94, LPA95]. Distributions [PHD^{+}95, BSCV95, GKH^{+}92, GKH^{+}93, HKMCS94, Pon94a, Pon94b, PSC^{+}95, VRT97, WO96, vHKS94]. Do [YWS^{+}94, BGvE^{+}97]. Domain [DRST03, GLPE97]. Domain-Decomposition [GLPE97]. Don’t [BGvE^{+}97]. Double [FKKC96]. down [Str94]. Downloadable [MCL^{+}95]. Draft [EEV^{+}96, Fox91, MCL^{+}95]. Driven [CMKH03, NJ94]. drugs [Str94]. Dusty [NOL97]. DVM [KKMP95]. Dybbuk [PSC^{+}93a, PDS^{+}93]. Dynamic [AMKS02, SR04, KB94]. Dynamics [BFHH94, DCBC98, DCR99a, DCR99b, HF95, QRH00].

Early [CL97a]. Earth [Ogi02, SMSY02].

Easy [Del98]. ECMWF [HK93, HK95]. Edinburgh [Fri94]. Edition [Lef98]. Editor [HKTW94, Hat94, Mei94]. Efficiency [BGvE^{+}97]. Efficient [BB02, CCW04, JB01, KHS96, KK94, KK95b, MCL^{+}95, SSC00, TCR96, TR96, CFPS94, KN95b, KHJS94, PPW94]. Eighth [ERS95, Sie94a, Sie94b]. Electromagnetic [CLiN^{+}02]. electromagnetics [PSG03]. Element [Ano94l, SM02b, KBKT94, OA02]. element-by-element [OA02]. ELF [MCL^{+}95]. Elimination [KW95]. Email [MCL^{+}95]. end [Lov94]. Engineering [LMR^{+}97, Str94, Ben99b, CKMU94, Eme94]. England [Eme94]. English [Met99d]. Enhanced [HCLJ03]. entitled [Wie94]. Enumeration [SVD96, SDv98]. Environment [BCC^{+}96b, DL97c, SS00, BCC^{+}97a, CEF^{+}95, HZ94, JA92, Vee94, BLLWW95]. Environments [DT94, Sch93, AES^{+}96, CDH^{+}94, DR94]. EPPP [NVG94]. Equation [LZ97]. Equipment [AOL94a]. Equivalent [BGvE^{+}97]. Erratum [KR95]. ESPRIT [CDH^{+}94, Hey94]. Essential [MCL^{+}95]. estimation [GB94]. etc [Met00c, Met01c, Met02a]. EURO [HAM95]. EURO-PAR [HAM95]. Europe [Ano93j, Ano93l, HMPT94]. European [BDLS96]. EuroPVM [BDLS96]. EUROSIM [DSZ94]. Evaluating [BBDR94, BBDR95]. Evaluation [AHOK02, BFHH94, BB96, Bra94a, Bra94b, Han98, LZ97, SF02, SM02b, SOG94, YFH97, BR98, Bou95, Din99, KKS^{+}95, KHRS95, MAH^{+}02, SM02a]. Evaluations [Din98]. Executing [BMMN94, BMMN95]. Execution [KMS^{+}95, RHH96, KHS95]. Exhibition [GH94a, GH94b, HS95, Ano93l]. Experience [CL97a]. Experiences [CNBB96, HKT93a, HKT93b, Sai95]. Experiments [Coe94b, SZG95]. Explained [Ano93a]. explicit [LC97]. exploitation [JA92]. Exploiting [HF95, OP98a, RSB97]. Expressing [MMV95]. Expressions [BGvE^{+}97, BBDR94, BBDR95, Mar93]. extended [Sch93]. Extending [CMZ94b, CZM94, MR95]. Extensions [AHOK02, UZCZ97, BCC^{+}97b, BCC^{+}97c, CC94, ISKvW02, SIOS02]. Extrinsic [Hig94c]. eyes [Str94].

F90 [WCC99]. F90/HPF [WCC99].

factorizations [LC97]. fall

[KKZ07, KKZ11, EEV^{+}96]. farming [Str94]. Farms [AOL94a, AOL94b, KC94]. Fast [AEG^{+}02, RHH96, GHSJ94]. FCRC [ACM96a]. Feasibility [KR94, KR95]. Features [Ano93b, Ano93h, UZCZ95, UZCZ96]. February [Ano93l, BBG^{+}95, IEE93a, IEE94a]. FEM [OA02]. Field [Ano94g]. Fifth [Ano95b, HK93, IEE93b, Ano93l]. File [Met99d, Met01c, Met02a, Met02b, BGMZ92, Bre92, CFPS94, Met99b, Met99c, Met00c, Met00a, Met00b, Met01b, Met01a]. Final [EEV^{+}96]. Financial [Jou95]. Find [MCL^{+}95]. Finding [DS02]. Fine [CL97a, CL97b]. Fine-Grain [CL97a, CL97b]. Finite [Ano94l, SM02b, KBKT94, PSG03]. finite-volume [PSG03]. Finland [Ano00b]. First [HAM95, Kum94]. Fitzroy [Eme94]. flexible [HC08]. Florida [ACM98]. flows [KBKT94]. Fluid [BFHH94, SMSY02, HF95]. Forecast [RHH96]. forecasting [GS95]. Form [BGvE^{+}97]. Formal [MCL^{+}95]. Forms [TR96]. Fortran [AMC01, Ano93a, Ano94k, Ano00b, DL97b, Eme94, Hig92, KK01, KKMP95, KR95, MB95, MCL^{+}95, Rag95, Sch97, UMM94, EEV^{+}96, BGMZ92, Bre92, Meh94, Pon94a, Pon94b, Sul88, Adv98, ADHF96, ACIK97, AOL94a, AOL94b, Ano92a, Ano92b, Ano92c, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93h, Ano93k, Ano94c, Ano94d, Ano94f, Ano94i, Ano94k, Ano96, Ano99, Ano00a, Ano00b, Ano01a, Ano01b, Ano01c, Ano02b, Ano02a, Ano03, AGG^{+}97, BCM^{+}93, BGvE^{+}97, BBZ94, Ben98, BZ99, Ben99b, BB02, BFHH94, BCFH93, BCF^{+}93a, BCF^{+}93b, BCF^{+}93c, BCF^{+}94c, BCF^{+}94b, BCF^{+}94d, BCF^{+}94a, BMN^{+}95, BMN^{+}97, Bra94a, Bra94b, BCC^{+}96a, BCC^{+}96b, BCC^{+}97b, BCC^{+}97a, BCC^{+}97c, CLiN^{+}02, Cel96, CMZ92, CZM93, CMZ93, CZVM94, CMZ94a, CMZ95, CCW04, CKZ93, Cou97, DDcMR96, DL97a, DS97]. Fortran [DL97c, DZ98, DBVS98, DCR99a, Din99, Eli98, EEV^{+}96, FSPC^{+}02, Hig93, FXAC94, Fox91, God93, GOS94, Hig94a, Hig94b, Hig94c, Hig94d, HM96, HHKT96, Han98, HBB^{+}95, Hat94, HF95, HKT92, HKT93a, HKT93b, HJT97, HJJ^{+}00, Iwa00, ISKvW02, KMR^{+}97, Ken94b, KK95a, KS02, KKZ07, KKZ11, KT00, KMBK96, KMS^{+}95, KC94, KOM93, KOM94, Koe92, Koe94, KGV97, KK94, KK95b, KR94, LMMW96, Lov93, Lov94, LC97, MM94, MMY95a, Meh93a, Meh93b, MVZ98a, MVZ98c, MVZ98b, MZ00, MZ01, Mei94, MH95, MCH96, Met95, Met99a, Met99b, Met99c, Met99d, Met00c, Met00a, Met00b, Met01b, Met01c, Met01a, Met02a, Met02b, MMV95, MMY95b, MCL^{+}95, MR95, NOL97, Off98, PFS^{+}04, PHHF94a, PHHF94b, PHHF95, PH96a, PH96b, Paz96, QRH00, RMCKB97, Sai95, SZM98, Sch96a, SNMC93, SHZ13]. Fortran [SIOS02, Ste93, SS97, SSG97, Tho93, Wag94, YGS^{+}94, YFH97, ZBC94, Zim02, Zos93, dSL98, van94a, vWAH^{+}02, vHKS94]. Fortran-based [NOL97]. Fortran-like [KGV97]. Fortran-Oriented [Ano96]. Fortran/HPF [UZCZ97]. Forum [Fox91, Mei94, MCL^{+}95, Str94]. foundations [PD96]. Fourier [GHSJ94]. Fourth [ACM93, IEE92a]. Framework [GOS94, RSB97, vDSP96, ACIK97, WCC99, vHK00, HM98]. France [BLT94, Van95].

Francisco [BBG^{+}95, IEE93a]. Free [Chw97, MCL^{+}95]. friendly [CFPS94]. front [Lov94]. Frontiers [IEE94a, IEE96b]. Frontiers’95 [IEE94a]. Frontiers’96 [IEE96b]. ftp [EEV^{+}96]. Fujitsu [AHOK02]. Function [BGvE^{+}97]. functions [vWAH^{+}02]. fused [AEG^{+}02]. Fusion [SMSY02]. Future [Ano92b, Ano92c, Ano93k, Zim02, MVZ98a, MVZ98c].

Generating

[CGL^{+}93, CGL^{+}95, SOG94, GKH^{+}92].

Generation

[AMC01, CMKH03, FGL01, KHS96, SSC00, GKH^{+}93, KHS95, TRV96, VRT97]. generator [vWAH^{+}02]. generators [AEG^{+}02]. Generic [BGvE^{+}97, BxCW01]. Genesis [Hey94]. Germany [BDLS96, GH94a, GH94b]. get [Str94, HDH^{+}94]. Getting [Chw97]. Give [DZ98]. global [ADH95, Ogi02, Str94]. Gmunden [Vol93]. Goals [Ano93c]. Grain [CL97a, CL97b]. Grained [BR96]. Grammars [MCL^{+}95]. Grand [Ten93]. graphic [Chw97]. Greece [HMPT94]. Grid [SR04]. Grids [BLW02]. Group [Iwa00, MCL^{+}95, Sch94]. Groups [Ano92b, Ano93k, EEV^{+}96]. Guest [Hat94]. Guide [BDPW98]. Guy [Eme94, Rag95, UMM94].

H [Eme94, UMM94]. Hague [Ano93j]. Halos [Ben99a, Bra00, Ben00]. Hampton [Wie94]. Handbook [Rag95, Koe94, UMM94, Eme94]. Handling [BBCR98, ABC^{+}96]. Hansen [Off98]. Haven [Ban93]. having [MIN^{+}95]. Hawaii [ERS95, ERS96, HS94]. HI [ERS96, HS94]. HICSS [ERS96]. HICSS-29 [ERS96]. hierarchical [BMV03]. High [ACM97, ACM98, AMC01, AOL94b, Ano94b, BBZ94, BCF^{+}94a, BCC^{+}96b, CMZ94a, DL97b, Don95, Dow93, Fos94, Fox91, GH94a, GH94b, Ger98a, Ger98b, God93, HS95, IEE94c, IEE96a, Lov93, Lov94, MCH96, Rag95, Sab95, Ten93, UMM94, Zos93, Ano93l, BCM^{+}93, BID95, BGMZ92, Bre92, CDF^{+}93, Duv92, Eme94, KC94, MKF95, OP99, SHZ13, SSG97, Adv98, ADHF96, ACIK97, AOL94a, Ano92a, Ano92c, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93h, Ano93k, Ano94c, Ano94d, Ano94i, Ano94k, AGG^{+}97, BGvE^{+}97, Ben98, BZ99, Ben99b, BB02, BFHH94, BMN^{+}95, BMN^{+}97, Bra94a, Bra94b, BCC^{+}96a, BCC^{+}97b, BCC^{+}97a, BCC^{+}97c, CLiN^{+}02, CZM93, CMZ93, CZVM94, CMZ95, CCW04, CKZ93, Cou97, DDcMR96, DL97a, DS97, DZ98, DBVS98, DCR99a, Din99, Eli98, FXAC94, GOS94, Hig92, HM96]. High [Han98, HBB^{+}95, Hat94, HF95, HJT97, HJJ^{+}00, Iwa00, KMR^{+}97, Ken94b, KK95a, KK01, KS02, KKZ07, KKZ11, KT00, KMBK96, KMS^{+}95, KOM93, KOM94, Koe92, Koe94, KGV97, KK94, KK95b, LMMW96, LC97, MB95, MMY95a, Meh93a, Meh93b, Meh94, MVZ98a, MVZ98c, MVZ98b, MZ00, MZ01, MH95, Met95, MMV95, MMY95b, MCL^{+}95, MR95, NOL97, Off98, PFS^{+}04, Paz96, RMCKB97, SZM98, Sch96a, Sch97, SNMC93, SIOS02, Ste93, SS97, Tho93, Wag94, YGS^{+}94, YFH97, Zim02, dSL98, van94a]. High-level [Ger98a, Ger98b, OP99]. High-Performance [Fos94, GH94a, HS95, IEE94c, GH94b, Ano93l, SHZ13, BFHH94, Bra94a, CZVM94]. Highly [HJT97]. HiPPI [JA92]. HIRLAM [GS95]. historical [KKZ07]. History [ACM07, MVZ98a, Zim02, MVZ98c]. HiWEP [Zim02]. Holberton [BGvE^{+}97]. Home [EEV^{+}96]. honor [Str94]. Honors [BGvE^{+}97]. HOPL [ACM07]. HOPL-III [ACM07]. House [Eme94]. HPC [Fox94]. HPF [ABC^{+}96, Ano94e, Ano94f, Ano96, Ano99, Ano00a, Ano00b, Ano01a, Ano01b, Ano01c, Ano02b, Ano02a, Ano03, AMKS02, AHOK02, BP97, Ben98, BSSV98, Ben99b, BLZ99, Ben99a, Ben00, BP98, BF01, BDPW98, Bod94, BV98, BB96, BR98, Bou95, BCFH93, BCF^{+}93b, BCF^{+}93c, BCF^{+}94c, BCF^{+}94b, BCF^{+}94d, BMMN94, BMMN95, BID95, BZ94, BD96, BG96, BCC^{+}97b, BCC^{+}97c, Bra00, BSCV95, BxCW01, BLW02, Bri00, BMV03, CNBB96, CMMP98, CMT01, CL97a, CL97b, CMZ94b, CZM94, CM98, CGSS94, Chw97, CDD^{+}96, Coe94a, Coe94b, Coe96, CA96, Cou97, DL97c, Del98, DS02, DCBC98, DCR99b, DRST03, Din98, EGKU99, EGKU02, EEV^{+}96, FGL01, FGRT00, FSPC^{+}02, FKK^{+}96b, FKKC96, FKK96a, Fox94, GLPE97, GS01, Guo01, GMS^{+}95, HKM98, HLJ01, HCLJ03, IK96, IHKvW02, ISKvW02, JB01, Jou95, KKS^{+}95]. HPF [KHS96, KMS^{+}95, LL98, Lef98, LZ97, MM94, MBFC98, MBFC99, Met99a, Met99b, Met99c, Met99d, Met00c, Met00a, Met00b, Met01b, Met01c, Met01a, Met02a, Met02b, MAH^{+}02, Nak95b, Nak95a, Nel96, NJ94, Ogi02, OA02, OP98a, OP98b, OP98c, OP99, OPP00, PSG03, PHHF94a, PHHF94b, PHHF95, PH96a, PH96b, PD96, Pon94a, Pon94b, Sai95, SM02a, SF02, SMSY02, SNK06, Sch96b, SZG95, SIOS02, SIDH95, SM02b, SVD96, SDv98, Smi95, Spo94, Str94, SS00, SN94, SN95, TBC94b, TCF94, TRV96, UZCZ97, Van98, Van94b, Vee94, WSL94, WCC99, ZCFL98, Zim99, vDSP96, vWAH^{+}02]. HPF-Builder [DL97c, Lef98]. HPF-combined [MIN^{+}95]. HPF-Like [Guo01, CMT01]. HPF-MPI [BP98]. HPF/Fortran [Ano94f, PHHF94a, PHHF94b, PHHF95, PH96a, PH96b]. HPF/JA [AHOK02, ISKvW02, Ogi02, SIOS02]. HPF/SX [MAH^{+}02]. HPF2 [BBCR98]. HPFBench [HJJ^{+}00]. HPFF [Ano92b, Mei94]. HPFIT [BCC^{+}96a, BCC^{+}96b, BCC^{+}97b, BCC^{+}97a, BCC^{+}97c]. HPFORTRAN [Sul88]. HUG2000 [Iwa00]. Hungarian [Fer92, FK95]. Hungary [Fer92, FK95]. hybrid [KJEM12].

I/O [BLW02, Coe94a]. I/Os [CFPS94]. IBM [GMS^{+}95, LC97, Sai95, Van98]. ICIP [IEE94b]. ICPP [Agr95]. Idaho [Str94]. Idea [BGvE^{+}97]. IEC [Ano00b]. IEEE [ACM97, ACM98, Kar95]. IFIP [CGS94, DR94]. Igniting [ACM03]. II [Ano94c, BCC^{+}97b, BCC^{+}97c, Hig94b, Hig94c, Hig94d]. III [ACM07, Ano94d, Hig94a]. image [MKS94]. Impact [Hat94]. impacts [Str94]. Implementation [Ano92b, BP98, BCF^{+}93c, CZVM94, KBKT94, KK94, KK95b, LZ97, MAH^{+}02, SOG94, vDSP96, BCF^{+}93b, BMV03, CMZ94a, CMZ95, CFMR95, DCR99a, KKS^{+}95, PSG03, QRH00]. implementations [CCW04, HKM98]. Implementing [BBG^{+}93, BD96, DL97a, GHSJ94]. Implementor [CKZ93]. implicit [KBKT94]. Implied [BGvE^{+}97]. Including [Cou97]. incompressible [KBKT94]. Incremental [KHS95]. independent [Ken94a]. Index [KHS96, KHS95]. India [IEE96a, Kum94, Pra95]. indirect [DSvH94]. Industrial [Ben98, Kon00, BLT94, BLZ99]. Information [Ano92b, Ano92c, Ano93k, Ano94j, Ano94k, Ano96, Met99d, Met01c, Met02a, Met02b, Met99b, Met99c, Met00c, Met00a, Met00b, Met01b, Met01a]. Innovation [ACM03]. Insight [IEE02]. Inspection [NJ94]. Installation [BDPW98]. Institute [Ano94l]. integer [BKK94]. Integrated [ASS95, BGG^{+}94, BCC^{+}96a, BCC^{+}96b, CFK^{+}94, BCC^{+}97b, BCC^{+}97a, BCC^{+}97c]. Integrates [FXAC94]. Integrating [CM98, CMVZ94, YWS^{+}94]. Integration [CMMP98, DCBC98]. Intel [KR94, KR95, SZG95]. interaction [DRST03, Eli98]. interactive [HKTW94, MKF95]. Interface [BV98, BG96, FKKC96, YGS^{+}94, BDPW98, BxCW01, HDH^{+}94]. Interfaces [BBZ94]. Interfacing [LMMW96]. International [ACM94, ACM95, ACM96a, Ano92c, Ano93i, Ano94a, Ano94g, Ano99, Ano00a, Ano00b, Ano01a, Ano01b, Ano01c, Ano02b, Ano02a, Ano03, BV94, CKMU94, ERS95, ERS96, Fri94, GH94a, GH94b, HMPT94, HAM95, HS95, HS94, HHK94, IEE95a, IEE96a, Kum94, PBG^{+}95, Pra95, Sie94a, Sie94b, Vol93, Ano93l, Ban93, BGNP94, BLT94, Hua96]. Internetworking [Ano93a]. Interpretations [Ano92b, Ano93k, Ano94k]. Interpreting [Ano94f, PHHF94b]. Interprocedural [HHKT96]. interstage [MIN^{+}95]. Intrinsic [Hig94b]. Introduction [Hat94, MH95, Sch96b, Zim99]. Invited [Meh93b, Zim02]. IO [Sni92]. IP [Ano93a, JA92]. IPPS [IEE95a]. iPSC [KR94, KR95]. iPSC/860 [KR94, KR95]. Irregular [BSSV98, Ben99a, BBCR98, Bra00, DL97a, DL97b, GLPE97, HJT97, PHD^{+}95, Sch96a, UZCZ97, Ben00, BCC^{+}97b, BCC^{+}97c, BSCV95, CC94, HMS^{+}95, MR96, Pon94a, Pon94b, PSC^{+}95, SPM^{+}94, dSL98]. ISBN [Eme94]. ISO [Ano00b]. ISO/IEC [Ano00b]. ISPAN [HHK94]. Issue [Ano94i, Hig94b, Hig94c, Hig94d, KS02]. Issues [Coe96, FGL01, Nak95b, CMT01, Nak95a]. Italy [DR94, Don95, HS95]. Iterative [DL97a, dSL98]. Ithaca [PBG^{+}95]. IWPP [Kum94]. IWPP-94 [Kum94]. IWWP [Kum94].

JA [AHOK02, ISKvW02, Ogi02, SIOS02]. January [ERS96, Eme94, HS94]. Japan [CKMU94, HHK94, Iwa00]. Japanese [SM02a, SF02]. Joint [BV94]. Jose [ACM97, Ano94a]. Journal [Ano93g, MCL^{+}95]. Jr [Eme94, Rag95, UMM94]. JTC1 [Ano00b].

JTC1/SC22/WG5 [Ano00b]. July [ACM95, Ano95b, HMPT94, IEE96c]. June [ACM07, Ano94h, Ano95a, DSZ94, Don95, Wie94]. Jupiter [Str94].

Kanazawa [HHK94]. KeLP [MBFC98, MBFC99]. Kemari [KMR^{+}97]. Kernels [YFH97]. KFKI [FK95]. KFKI-1995-2 [FK95]. KFKI-1995-2/M [FK95]. Knoxville [IEE94c]. Koelbel [Eme94, Rag95, UMM94].

L [Eme94, UMM94]. Lab [Str94]. Lahey [MCL^{+}95]. Langley [Wie94]. Language [Ano93e, Ano93k, Ano94c, Ano94d, Ano94i, Ano94k, BFHH94, Hig93, Fox91, Guo01, Hig92, Hig94a, Hig94b, Hig94c, Hig94d, Cel96, HMS^{+}95, KKMP95, UZCZ95, UZCZ96, dLD85, CC94]. Languages [ACM07, Ano93i, Ano94a, Ban93, BGNP94, CZVM94, CMZ95, CMKH03, Fos94, Fox94, Hua96, Mar93, PZA93, PBG^{+}95, PHD^{+}95, Sch93, SS96, ZA93, ASM^{+}94, CMZ94a, HMPT94, KGV97, Nic91]. LARC [Wie94]. large [BxCW01]. laser [Eli98]. lasers [Ano94h]. Latest [MCL^{+}95]. Lattice [Ano94g]. Layer [OP98a, FKK96a, PFS^{+}04]. Layout [KK95a, KK98, BKK94]. LCPC [Hua96]. Lecture [Meh93b]. Length [Cou97]. lesson [KKZ07]. level [Ger98a, Ger98b, MKF95, OP99, SNK06, SHZ13]. Libraries [BGvE^{+}97, DCBC98, EGKU99, IEE94d, IEE95b, BCF^{+}94a]. Library [Bra00, CMKH03, Hig94b, Chw97, FKK^{+}96b, HKM98, KN95a, PQ94, vWAH^{+}02]. Life [NOL97, Str94]. Lightweight [IHKvW02]. Like [Guo01, CMT01, KGV97]. Limitations [Meh93b, Meh93a]. Line [MCL^{+}95]. Linear [ACIK97, KNS95b, LZ97, KNS95a]. Linear-Time [KNS95b, KNS95a]. Linked [BGvE^{+}97]. Linker [MCL^{+}95]. Linking [MCL^{+}95]. Linux [Ano96, Del98]. Linz [BV94]. List [EEV^{+}96]. Lists [BGvE^{+}97].

Living [MCL^{+}95]. Local [CGL^{+}95, EGKU02, SVD96, CGL^{+}93, SDv98]. locality [ADH95]. localization [HC08]. Lockheed [Str94]. London [Eme94, MCL^{+}95]. Loop [FGL01, IK96, LPA95, LP93]. Loops [BGvE^{+}97, YWS^{+}94]. lossy [MIN^{+}95]. Loveman [Eme94, Rag95, UMM94]. Low [Str94]. Low-life [Str94]. Lowering [BV98]. LPF [MIN^{+}95]. LPF/HPF [MIN^{+}95]. LPF/HPF-combined [MIN^{+}95]. LU [LC97].

M [FK95, FXAC94]. M.I.T [Eme94]. machine [BDLS96, Ken94a]. machine-independent [Ken94a]. Machines [BR96, BMN^{+}97, HLJ01, KHS96, KK98, BBDR94, BBDR95, HKT92, HMS^{+}95, KN95a, KHS95, SNK06, TBC94a, WSL94]. Macintosh [Ano96]. magnetohydrodynamic [KT00]. magnetosphere [Ogi02]. Mailbase [EEV^{+}96]. MAKE [MCL^{+}95]. makes [Str94]. management [AW94]. managing [Off94]. mapping [HC08, KN95b, MKF95, SNK06, SV95]. Mappings [Lef98]. March [Ano94b]. Mary [Eme94, Rag95, UMM94]. Maryland [IEE96b]. Massively [DSZ94, IEE94a, IEE96b, ASM^{+}94, BBDR94, BBDR95, DR94, Ger98a, Ger98b, Sta94]. match [MIN^{+}95]. Mathematically [BGvE^{+}97]. Matrices [MCL^{+}95, dSL98]. Matrix [DL97a, DL97b, UZCZ95]. May [ACM93, ACM96a, DT94, Hig94a, Hig94b, Hig94c, Hig94d, HS95, IEE94c, SS96]. McLean [IEE94a]. MD [IEE02]. measurements [SZG95]. Media [Ano93a]. Meek [BGvE^{+}97]. Meeting [Ano93j, Ano95a, Iwa00, Mei94, MCL^{+}95, Ano00b]. Meetings [Ano92b, Ano92c, Ano93k, Ano99, Ano00a, Ano00b, Ano01a, Ano01b, Ano01c, Ano02b, Ano02a, Ano03]. Memory [BR96, BCF^{+}93c, BMMN94, BMMN95, BMN^{+}97, CL97a, CL97b, HLJ01, KHS96, KNS95b, KK98, RSB97, Sch93, BZ99, BB02, Bod94, BCF^{+}93b, BCF^{+}94b, BCF^{+}94d, Ger98a, Ger98b, GHSJ94, GS97, HBB^{+}95, HKT92, HMS^{+}95, KN95a, KMR^{+}97, KHS95, KNS95a, PZA93, SNK06, TBC94a, Wag94, WW95, WI94, ZA93]. Merge [YWS^{+}94]. Message [FKKC96, vDSP96, GS95, Hem96, LC97]. Message-Passing [vDSP96, Hem96]. Metcalf [Ano96]. Meteorology [HK93, HK95]. methodology [CDF^{+}93, GKH^{+}92]. Methods [CMK00, MMV95, Don95, KBKT94, PSC^{+}95]. Mexico [Sie94a, Sie94b]. MHD [Ogi02]. Michael [Ano96]. Microsoft [Sul88]. Migrating [MM94]. Milan [HS95]. MIMD [BCF^{+}93b, BCF^{+}93c, BCF^{+}94b, BCF^{+}94d, HKT92]. Mind [Ano93k]. mining [WW94]. Minneapolis [Ano94l, IEE92b]. Minnesota [IEE92b]. Mississippi [IEE94d, IEE95b]. Mixed [OPP00]. MN [Ano94l]. Model [Guo01, HCLJ03, MKS94, NOL97, BMV03, CZM93, CMZ93, GS95, HBD^{+}93, KB94, PFS^{+}04, PD96, Str94]. Model-based [MKS94]. Modeling [FGRT00, SS00, KHRS95]. Models [CMK00, RHH96, CDF^{+}93, Duv92, YO95]. Modern [Cel96]. Module [BGvE^{+}97, Cou97, MCL^{+}95]. Modules [BGvE^{+}97, MCL^{+}95]. modulus [AEG^{+}02]. Molecular [DCBC98, DCR99a, DCR99b]. monolithic [MIN^{+}95]. Monte [MMV95]. Montreal [CGS94]. MPI [BP98, BF01, CFMR95, CDD^{+}96, DZ98, FKK^{+}96b, FKK96a, Hem96, IEE96c, LZ97, OP98b, OP98c, SM02b, WO96]. MPI-based [OP98b, OP98c]. Multi [KHRS95, BSCV95, PFS^{+}04]. multi-dimensional [BSCV95]. multi-layer [PFS^{+}04]. Multi-phase [KHRS95]. multicomputer [KW94]. Multicomputers [RSB97, WW95, WI94]. Multiconference [Ten93]. multidisciplinary [CMVZ94]. Multiple [MBFC98, MBFC99]. multiply [AEG^{+}02]. Multiprocessor [SR04, Wag94]. MultiProcessors [BMV03, Sch93, AW94, GHSJ94]. Munich [BDLS96, GH94a, GH94b].

N [FK95, DCR99a, MB95]. N-body [MB95]. NAS [AHOK02, CDD^{+}96, Sai95]. National [Ano92b, Ano00a, Ano00b, Ano01b, Ano01c, Ano02b, Ano02a, Ano03, Str94, Ano93k]. NC [Agr95]. near [CCW04]. Nested [EEV^{+}96, PPW94]. Net [EEV^{+}96, BGvE^{+}97, MCL^{+}95]. Netherlands [Ano93j, Ano93l, DSZ94]. nets [Str94]. Network [Coe94b, BID95, MIN^{+}95]. Networking [ACM97, ACM98, GH94a, GH94b, HS95]. Networks [HHK94]. Neural [Str94]. News [Str94]. Ninth [ERS96]. non [KB94]. non-uniform [KB94]. normalization [LP93]. Note [GS01]. Notes [EEV^{+}96]. Notre [IEE96c]. Nov [Ano92b]. November [ACM96b, ACM97, ACM98, ACM03, BGG^{+}94, Fox91, HK95, IEE92b, IEE93c, IEE94b, IEE94e, IEE02]. nuclear [SHZ13]. NUMA [AW94, LP93]. number [AEG^{+}02]. Numerical [EGKU99, FP92]. NY [PBG^{+}95, SS96].

O [BLW02, Coe94a]. obituary [BGvE^{+}97]. Object [BGvE^{+}97, Bod94, KKZ07, QRH00]. Object-Oriented [BGvE^{+}97, QRH00]. Objects [MCL^{+}95]. October [Ano93j, Ano94a, Ano94g, Ano94l, BDLS96, BGG^{+}94, Fer92, FK95, GGK^{+}93, IEE94d, IEE95b, IEE96b, Sch93, Vol93, Iwa00]. Offshore [CKMU94]. Ohio [Hua96]. Oil [KR94, KR95]. Ontario [BGG^{+}94, GGK^{+}93]. OpenMP [BF01, Bri00, BMV03, CM98, KJEM12]. Operating [Ano94a]. Operation [HLJ01]. Operational [RHH96]. operations [BGMZ92, Bre92, Off94]. optical [Chw97].

Optimal [CA96, SV95]. Optimization [AMKS02, BGvE^{+}97, OA02, WW95, ADH95, MCAB^{+}02, TRV96]. Optimizations [Nel96, WCC99]. Optimize [HLJ01, GKH^{+}92]. Optimized [DCBC98]. Optimizing [BP97, Ben99a, Ben00, BMN^{+}97, CL97b, EGKU02]. Orange [ACM98]. ordination [OP99]. Oregon [BGNP94, IEE93c]. Organization [Ano93k, Ano94j, Ano94k, Ano99, Ano00a, Ano00b, Ano01a, Ano01b, Ano01c, Ano02b, Ano02a, Ano03]. Oriented [Ano96, BGvE^{+}97, QRH00]. Origin2000 [Bri00]. Orlando [ACM98]. Osaka [CKMU94]. Oulu [Ano00b]. Out-of-Core [TBC94b, TBC94a]. overlap [BBDR94, BBDR95]. Overview [Koe92, KK94, KK95b, dLD85, Hey94, ZCP95, Zos93]. Ownership [JB01].

P [PQ94]. PA [ACM96b, Ano95a]. Pacific [Van94b]. PACK [BR96]. PACK/UNPACK [BR96]. PACT [CGS94]. Page [EEV^{+}96]. Pandore [AFMP95]. Paperback [Eme94]. papers [Ano93l, IEE93a]. Paradigms [CM98]. Paragon [SZG95]. Parallel [ACM93, Agr95, Ano93i, Ano94l, Ano95b, AHOK02, BR96, BBG^{+}95, BP98, BDLS96, BMMN94, BMMN95, Bra00, BLW02, BV94, CZM94, CGL^{+}95, CH94, CGS94, DL97b, DBVS98, DCBC98, Fox94, FP92, Guo01, HK93, HK95, HHK94, HCLJ03, IEE92a, IEE93b, IEE94a, IEE94d, IEE95b, IEE96b, Ken94b, KNS95b, Kon00, Kum94, Lev94, LZ97, LMR^{+}97, Meh93b, PHD^{+}95, Pra95, SZM98, SSC00, Sie94a, Sie94b, Ste95, Str94, TR96, Vol93, YGS^{+}94, dSL98, vDSP96, AES^{+}96, ASM^{+}94, AFMP95, ABC^{+}96, Ban93, BGNP94, BB02, BBG^{+}93, Bod94, BBDR94, BBDR95, BID95, BxCW01, Cel96, CMZ94b, CGL^{+}93, CC94, CCW04, CEF^{+}95, CDD^{+}96, CFPS94, CDH^{+}94, DR94, DSZ94, DT94, Duv92, FKK^{+}96b, Ger98a, Ger98b, GS97, GS95, HMPT94, HAM95, HZ94, HKTW94, Hua96, IEE95a, KKS^{+}95, KMR^{+}97, Kas93, Ken94a]. parallel [KNS95a, KB94, KKMP95, LPA95, LC97, Meh93a, MKS94, MR96, Nic91, Off94, PSG03, PQ94, Per94, PD96, PBG^{+}95, QRH00, Sab95, Sta94, SV95, TBC94a, UZCZ95, UZCZ96, YWS^{+}94, YO95, ZCP95, BLLWW95]. Parallelisation [HBD^{+}93]. parallelise [PFS^{+}04]. Parallelism [CFK^{+}94, Fos94, FKKC96, GOS94, MBFC98, OP98a, RSB97, ADH95, Bod94, CMVZ94, Mar93, MBFC99, OPP00, PPW94, PQ94]. Parallelization [BB96, BCC^{+}96a, BCC^{+}96b, DCR99b, Eli98, IK96, Van98, BCC^{+}97b, BCC^{+}97a, BCC^{+}97c, CDD^{+}96, DDcMR96]. Parallelizing [Adv98, ASS95, BSSV98, DPR94, RHH96, SR95, HDH^{+}94, HDH^{+}95]. parameters [YO95]. parity [SHZ13]. parity-dependent [SHZ13]. PARLE [HMPT94]. Part [Ano93e, Ano94c, Ano94d, Hig94a, Met02a, Met02b, BCC^{+}97b, BCC^{+}97c, BCC^{+}96b, BCC^{+}97a, Hig94b, Hig94c, Hig94d, Met99b, Met99c, Met99d, Met00c, Met00a, Met00b, Met01b, Met01c, Met01a]. Participants [MCL^{+}95]. Particle [ADHF96, CLiN^{+}02]. Particle-in-Cell [ADHF96, CLiN^{+}02]. Partitioning [PSC93b, LPA95]. partners [Str94]. Pass [MCL^{+}95]. Passing [FKKC96, vDSP96, GS95, Hem96, LC97]. patterns [DRST03]. PC [Ano96, CLiN^{+}02, Bod94]. PCRC [ZCFL98]. PCRC-based [ZCFL98]. PCTE [HZ94]. PDS [HKM98]. Pennsylvania [ACM96a]. Performance [ACM97, ACM98, Adv98, AMC01, ADHF96, ACIK97, AOL94a, Ano92a, Ano92c, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93h, Ano93k, Ano94c, Ano94d, Ano94f, Ano94i, Ano94k, AGG^{+}97, BGvE^{+}97, BBZ94, Ben98, BZ99, Ben99b, BB02, Bou95, BCF^{+}93c, BCF^{+}94a, BMN^{+}95, BMN^{+}97, Bra94b, BCC^{+}96a, BCC^{+}96b, BCC^{+}97b, BCC^{+}97a, BCC^{+}97c, CLiN^{+}02, CMMP98, CMT01, CZM93, CMZ93, CMZ95, CCW04, CKZ93, Cou97, DDcMR96, DL97a, DL97b, DS97, DZ98, DBVS98, DCR99a, Din99, Dow93, EGKU02, Eli98, FGRT00, FXAC94, Fos94, Fox91, GH94a, GOS94, Hig92, HM96, Han98, HBB^{+}95, Hat94, HF95, HS95, HJT97, HJJ^{+}00, IEE94c, IEE96a, Iwa00, KMR^{+}97, Ken94b, KK95a, KK01, KS02, KKZ07, KKZ11, KT00, KMBK96, KMS^{+}95, KOM93, KOM94, Koe92, Koe94, KGV97, KK94, KK95b, LMMW96]. Performance [Lov94, LC97, MB95, MMY95a, Meh93a, Meh93b, Meh94, MVZ98a, MVZ98c, MVZ98b, MZ00, MZ01, MH95, MCH96, Met95, MMV95, MMY95b, MCL^{+}95, MR95, NOL97, Off98, PFS^{+}04, PHHF94b, PH96a, PH96b, Paz96, Rag95, RMCKB97, SF02, SZM98, Sch96a, Sch97, SNMC93, SIOS02, SM02b, Ste93, SS97, SSG94, Ten93, Tho93, UMM94, Wag94, YGS^{+}94, YFH97, Zim02, dSL98, van94a, AOL94b, Ano93l, Ano94b, BCM^{+}93, BCF^{+}93b, BGMZ92, Bre92, BMV03, CMZ94a, CDF^{+}93, DS02, Don95, Duv92, Eme94, GH94b, God93, KKS^{+}95, KC94, Lov93, Sab95, SM02a, SZG95, SHZ13, SSG97, Zos93, BFHH94, Bra94a, CZVM94]. performance-prediction [BMV03]. Performances [DCR99b]. PerfVisS [KC94]. Periodicals [Ano92a, Ano96]. Perspective [Fox94, Smi95]. PETSc [HKM98]. PGHPF [BMN^{+}97, Sch94]. phase [KHRS95]. Philadelphia [ACM96a]. Phoenix [ACM03]. PIC [BMV03]. Pipeline [SR04]. Pipelined [BD96]. Pittsburgh [ACM96b, Ano95a]. Placement [EEV^{+}96, vHK00]. planning [MKF95]. Plans [Ano94k, WSL94]. Plasma [DBVS98, Eli98]. Plasmadynamics [Ano94h]. PM [MB95]. Pointer [BGvE^{+}97]. Pointers [BGvE^{+}97]. Poisson [BP98]. polar [CKMU94]. Pollution [SS00]. Portable [CH94, CDH^{+}94, RHH96, Sta94, AFMP95, CEF^{+}95, HZ94, KN95a, KMR^{+}97, KKMP95]. Porting [NOL97, SR96, SN94, SN95, Sai95]. Portland [BGNP94, IEE93c, Sch94]. poster [Sch93]. Power [Sai95, MIN^{+}95, Ano96]. PPOPP [ACM93, Ano95b]. PPPE [CDH^{+}94]. PPTran [KMBK96]. Practical [Din98, KOM94, Din99, KOM93]. Practice [ACM93, Ano95b]. Pre [BR98]. Pre-evaluation [BR98]. Predictable [Ano93a]. Prediction [PH96a, PH96b, BMV03, CDF^{+}93]. Preliminary [BFHH94, HKT93a, HKT93b]. PREPARE [BBZ94, BSCV95, Vee94]. presentations [Sch93]. presented [Ano00b]. Press [Eme94]. Price [Eme94]. Principles [ACM93, Ano93a, Ano95b]. print [BGvE^{+}97]. private [Str94]. problem [Sab95]. Problems [BBCR98, DL97a, DL97b, HJT97, UZCZ97, BCC^{+}97b, BCC^{+}97c, SPM^{+}94]. Procedure [BBZ94, BV98, YO95]. Procedures [Hig94b, Hig94c]. Proceedings [ACM93, ACM94, ACM96b, ACM97, ACM98, Agr95, Ano93i, Ano93j, BBG^{+}95, BGG^{+}94, ERS95, ERS96, Fer92, FK95, GGK^{+}93, HS94, HK93, IEE92a, IEE92b, IEE93b, IEE93c, IEE94c, IEE94d, IEE95b, IEE96c, IEE02, Kar95, Sie94a, Sie94b, Ten93, ACM96a, ACM07, Ano94a, Ano94g, Ban93, BGNP94, BLT94, BDLS96, BV94, CGS94, DSZ94, Fri94, GH94a, GH94b, HMPT94, HAM95, HS95, HK95, HHK94, Hua96, IEE94a, IEE94b, IEE96b, IEE96a, Kum94, PBG^{+}95, Van95, Vol93, IEE94e]. Processing [Agr95, BBG^{+}95, BBZ94, BSCV95, IEE92a, IEE93b, Sie94a, Sie94b, ASM^{+}94, BV94, Chw97, DSZ94, HAM95, IEE95a, Kas93, Kum94, MKS94, Nic91, Pra95]. Processor [KN95b, HC08]. Processors [HK93, HK95, KMR^{+}97, Sta94]. Products [Chw97]. program [KKMP95]. Programmers [Ano93a]. Programming [ACM93, ACM07, Ano94a, Ano95b, BF01, CMK00, CMZ92, DR94, Del98, EEV^{+}96, Ken94b, Meh93b, SZM98, SS97, AES^{+}96, ASM^{+}94, BB02, BKK94, Bod94, CEF^{+}95, CDH^{+}94, DS97, Duv92, Ger98a, Ger98b, GS95, HZ94, HKTW94, Ken94a, KB94, Meh93a, MKS94, Per94, PD96, PZA93, Vee94, ZA93, dLD85]. Programs [AMKS02, BMMN94, BMMN95, BZ94, CL97b, CGL^{+}95, CH94, FGL01, HLJ01, HCLJ03, KNS95b, KMS^{+}95, SSC00, SR96, TBC94b, TCF94, TR96, CGL^{+}93, FSPC^{+}02, GB94, GS97, HMS^{+}95, KNS95a, KC94, SR95, SSG97, TBC94a, YO95]. Progress [Fei94]. Project [Zim99, BLZ99, CDH^{+}94, Hey94]. Promise [Ano93j]. Promises [Meh93b, Meh93a]. Proposal [Sni92]. Prospects [MVZ98b]. Protocols [Ano93a]. pseudorandom [AEG^{+}02]. pseudorandom-number [AEG^{+}02]. Public [Str94]. Public-private [Str94]. Publications [Ano92a]. PUT [HDH^{+}94]. PUT/GET [HDH^{+}94]. PVM [BDLS96, BID95, Hem96, OPP00, SR96, Str94].

QCD [Sta94]. QR [LC97]. Query [HM96, HM98].

R [Wie94]. R&D [Str94]. R&D-100 [Str94]. race [CFMR95]. Raleigh [Agr95]. RAM [Chw97]. range [KW95]. read [BGvE^{+}97]. Reading [HK95]. Ready [DZ98]. Real [Ano93a, Bra94a, Chw97, Bra94b, SIOS02, Ano94k]. Real-Time [Ano93a, Chw97, Ano94k]. real-world [SIOS02]. Realistic [SN94, SN95]. Reality [Ano93j]. Realization [BG96, PD96]. Reasoning [Ste95]. recognition [PQ94]. Recompilation [AMKS02]. Red [Nel96]. Red-Black [Nel96]. redefinition [Sul88]. Redistribution [TCF94, TCR96, WO96, HC08, KN95a, KN95b, KHJS94, KHRS95, WW94, WW95].

Redistributions [BG96, GHSJ94]. reduce [GS97]. reductions [YWS^{+}94]. redundant [KW95]. relation [Hem96]. Remappings [CA96]. remote [GS97]. Replacement [BGvE^{+}97]. replay [CFMR95]. Report [DZ98, Iwa00, Ste93, Ano00b]. Reports [Ano94e]. Representation [BV98]. Research [ASM^{+}94, GS01, AGG^{+}97, BLT94, Wie94, Lev94]. Reservoir [KR94, KR95]. Resolution [Str94]. restructuring [DDcMR96, LP93]. Results [BGvE^{+}97, BCF^{+}93c, MMY95b, NOL97, BCF^{+}93b, MMY95a]. Retargetable [BCM^{+}93, SNMC93]. Reuse [PSC93b]. Review [Eme94, Rag95, UMM94]. revived [Cel96]. Rice [MCAB^{+}02]. rise [KKZ07, KKZ11]. Robert [Eme94, Rag95, UMM94]. Robot [NJ94]. robust [KB94]. Role [Wie94]. Royale [BLT94]. Run [OP98b, OP98c, PQ94, Sch93, SPM^{+}94, SS96, OP99, Str94]. Run-Time [OP98b, OP98c, Sch93, SS96, PQ94, SPM^{+}94, OP99]. Runtime [ASS95, AES^{+}96, HMS^{+}95, PSC93b, PSC^{+}95, TBC94b, TCF94, ASS93, AFMP95, BBG^{+}93, PSC^{+}93a, PDS^{+}93].

S [Eme94, UMM94]. SAC [GS01]. Saline [BLT94]. San [ACM93, ACM97, ACM07, Ano94a, BBG^{+}95, IEE93a, Kar95]. Santa [Ano95b, IEE95a]. save [BGvE^{+}97]. SC2002 [IEE02]. SC2003 [ACM03]. SC22 [Ano00b]. SC97 [ACM97, ACM97]. SC98 [ACM98, ACM98]. scalability [SSG94]. Scalable [BCF^{+}94a, Fox94, IEE94c, IEE94d, IEE95b, SS96, BB02, BBG^{+}93, PSG03, ZCP95]. ScaLAPACK [BDPW98, BG96, LMMW96]. scale [CDF^{+}93]. Schedule [Mei94, PSC93b]. Scheduling [KK94, KK95b]. scheme [HBD^{+}93]. Schemes [SVD96, SDv98]. School [Van95]. Schreiber [Eme94, Rag95, UMM94]. Science [HK93, SMSY02]. Sciences [ERS95, ERS96, HS94]. Scientific [Adv98, BBG^{+}95, CNBB96, DS97, Din98, Eme94, LMR^{+}97, SS97, Ben99b, Bou95, Cel96, Din99, DT94, KB94, PD96, SM02a]. scientists [Str94]. Scope [Ano93c]. Second [IEE96c, Vol93]. sees [Chw97]. Segmented [HCLJ03]. Semantics [Guo01]. Sempa [LMR^{+}97]. Senans [BLT94]. September [BLT94, BV94, FK95, Sch93, Van95]. Sequence [Hig94d, KNS95b, KNS95a]. Sequences [TR96, SV95]. sequential [SR95]. series [Eme94]. Server [Ano93j]. Session [EEV^{+}96]. Set [BCC^{+}96a, BCC^{+}96b, KHS96, BCC^{+}97b, BCC^{+}97a, BCC^{+}97c]. Sets [CGL^{+}95, JB01, BxCW01, CGL^{+}93, KHS95]. Seventh [BBG^{+}95, HS94]. Several [MMY95b, MMY95a]. SGI [Sai95]. SHARE [Ano93j]. Shared [BMMN94, BMMN95, CL97a, CL97b, BB02, Bod94, Ger98a, Ger98b]. Short [BGvE^{+}97, ZCP95]. SHPF [MCH96]. SIAM [BBG^{+}95]. Sierra [Van94b]. Significant [BGvE^{+}97]. SIGPLAN [ACM07, ACM93, Ano95b]. Simulation [ADHF96, DBVS98, KR94, KR95, MMV95, SMSY02, Ten93, DCR99a, Ogi02]. Simulations [MB95, SM02b, KT00, QRH00]. Simulator [SMSY02]. Sixth [Ano94a, HK95, IEE96b]. SLHPF [BDPW98]. Slicing [DSvH94]. Smithsonian [Str94]. Snobs [BGvE^{+}97]. SofTech [Spo94]. Software [Ano96, BLLWW95, HS94, LMR^{+}97, CMVZ94, Kas93, Str94]. Solution [DL97b]. Solutions [BGG^{+}94]. Solver [BP98, LZ97, Van98]. Solvers [DL97a, dSL98]. solving [Sab95]. Some [Bra94a, Per94, Bra94b]. Sopron [Fer92]. Source [KMBK96]. SP [LC97]. SP-2 [LC97]. SP2 [GMS^{+}95, Sai95, Van98]. Spain [ACM95]. Sparse [DL97a, DL97b, UZCZ97, SZG95, UZCZ95, UZCZ96, dSL98]. Speak [Ano93k]. Special [Ano94i, Hig94b, Hig94c, Hig94d, KS02, SF02]. Specification [Ano93e, Ano93k, Ano94c, Ano94d, Ano94i, Ano94k, Hig93, Fox91, Hig94a, Hig94b, Hig94c, Hig94d, Hig92]. specified [PSC^{+}95]. spectral [Eli98, GS95]. speed [BID95]. SPiDER [FSPC^{+}02]. spin [SHZ13]. spin- [SHZ13]. spring [IEE93a]. Springs [Ano94h]. standard [Hem96, Ano92b, Ano93k, Ano94k, Ano02a, Ano03, EEV^{+}96]. Standardize [BGvE^{+}97]. Standards [Ano92b, Ano92c, Ano93k, Ano02b, Ano02a, Ano03, Fei94, FKKC96, Ano93k, Ano94j, Ano94k, Ano99, Ano00a, Ano00b, Ano01a, Ano01b, Ano01c, Ano02b, Ano02a, Ano03]. State [IEE94d, IEE95b]. statement [KHS95]. Statements [BGvE^{+}97, BBZ94, KHS96, SOG94]. Static [GS97, ACIK97]. Status [BGvE^{+}97, DZ98, MVZ98b, Nak95b, Nak95a, Zim02, Hem96, MVZ98a, MVZ98c, Ste93]. Steele [Eme94, Rag95, UMM94]. Stencils [RMCKB97]. Stockholm [HAM95]. Storage [Hig94d, MCL^{+}95, SVD96, SDv98]. storm [CDF^{+}93]. storm-scale [CDF^{+}93]. Strategies [BB96, DCR99b, MCAB^{+}02]. Street [Eme94]. Strength [Kon00]. String [MCL^{+}95]. Strings [Cou97]. strip [WW94]. structure [BCC^{+}97b, BCC^{+}97c, Off94]. Structured [ASS95, ASS93]. structures [KGV97]. Study [BF01, GLPE97, GS01, KR94, KR95, SS97, SN94, SN95, Bri00, DS97]. Style [SS00]. subscript [KW95]. Subscripts [SSC00]. Subset [Ano93b, BGvE^{+}97, MCH96]. Suite [SF02, DS02, HJJ^{+}00]. Summation [EEV^{+}96]. SUPERB [ZBC94]. Supercomputer [Ano94l, Str94]. Supercomputing [ACM94, ACM95, ACM96a, ACM96b, Ano93l, HK93, IEE92b, IEE93c, IEE94e, Kar95]. Support [Ano94a, Bra00, BLW02, CFK^{+}94, Ken94b, MR95, OP98b, OP98c, Sch96a, TBC94b, ASS93, AES^{+}96, HDH^{+}94, HDH^{+}95, HMS^{+}95, OP99, PSC^{+}95, SPM^{+}94]. supported [CDD^{+}96]. Supporting [BLZ99, Pon94a, Pon94b, PHD^{+}95]. Survey [Paz96]. Sweden [HAM95]. Switch [EEV^{+}96]. SX [MAH^{+}02]. Symbol [EEV^{+}96]. symbolic [FSPC^{+}02]. Symmetric [BMV03]. Symposium [ACM93, Ano94b, Ano94g, Ano94l, Ano95b, HHK94, IEE92a, IEE93b, IEE94a, IEE96b, Sie94a, Sie94b, Ten93, IEE95a]. synchronization [GS97]. Synthesis [HLJ01, Per94]. System [Ano93a, CMMP98, ERS95, ERS96, FXAC94, HS94, Kas93, BBG^{+}93, CFPS94, KMR^{+}97, Lev94, MCH96, PSG03, PSC^{+}93a, PDS^{+}93, dLD85]. Systematical [NJ94]. Systems [Ano94a, BMMN94, BMMN95, HBB^{+}95, SS96, vDSP96, BB02, BBG^{+}93, DR94, PZA93, Wag94, ZA93].

T3D [SZG95]. T3E [Din98, PSG03]. Talk [Zim02]. talks [Sch93]. Targets [BGvE^{+}97]. Task [CFK^{+}94, Fos94, FKKC96, Fox94, GOS94, OP98a, RSB97, CMVZ94, OPP00, PQ94]. Tasks [OP98b, OP98c, DRST03, FKK^{+}96b, FKK96a, OP99, SV95]. TBSCM [BP98]. TCP [Ano93a, JA92]. TCP/IP [Ano93a, JA92]. Teach [BGvE^{+}97]. Technique [AMKS02, SR04, HC08]. Techniques [Adv98, BMMN94, BMMN95, PSC93b, CGS94, HKMCS94, KN95b, MKF95]. Technology [HS94, ABC^{+}96, Don95, Ken94a, ZCP95]. Telescoping [CMKH03]. template [Chw97]. templates [CZM93, CMZ93]. Templex [Chw97]. Tennessee [IEE94c]. Terabytes [IEE02]. Terms [Ano93f]. Texas [Ano94g, IEE92a, IEE93b, IEE94b]. Textbook [MCL^{+}95]. TFLOPS [SMSY02]. Their [CZVM94, UZCZ97, CMZ94a, CMZ95]. Theory [Ano94g]. therapy [MKF95]. Thinking [WSL94]. Third [ACM07, BDLS96, BV94]. Three [CLiN^{+}02, Ogi02, SMSY02, Eli98]. Three-Dimensional [CLiN^{+}02, Ogi02, SMSY02, Eli98]. Tight [DCBC98, DCR99a]. Tight-Binding [DCBC98, DCR99a]. Time [ASS95, Ano93a, KNS95b, OP98b, OP98c, PH96a, PH96b, Sch93, SS96, Chw97, GB94, KNS95a, OP99, PQ94, SPM^{+}94, Ano94k]. TN [DT94]. TNO [DS02]. Tokyo [Iwa00]. Tool [BZ94, CMMP98, SR96, AGG^{+}97, DDcMR96, HKTW94, SSG97]. Toolkit [PHHF94a, PHHF95]. Tools [BCC^{+}96a, BCC^{+}96b, CDD^{+}96, Met02a, Paz96, BCC^{+}97b, BCC^{+}97a, BCC^{+}97c, DT94, Met00b]. Tools-supported [CDD^{+}96]. Toronto [BGG^{+}94, GGK^{+}93]. Townsend [DT94]. TPL [dLD85]. TR92225 [Fox91]. Transfer [SR04]. Transform [DL97c]. Transformation [BZ94]. transformational [vWAH^{+}02]. transforms [GHSJ94]. translation [Sul88]. Translator [KMBK96]. transport [PFS^{+}04]. transportable [dLD85]. Transputer [Fer92, FK95]. Transputers [BLT94]. TransTOOL [BCC^{+}97a, DDcMR96, BCC^{+}96b]. Trends [Duv92]. Triangular [MCL^{+}95]. triangulation [CCW04]. Tridiagonal [Van98]. Trivandrum [IEE96a]. Troy [SS96]. Tuning [Ano93a]. Twenty [ERS95, ERS96, HS94]. Twenty-Eighth [ERS95]. Twenty-Ninth [ERS96]. Twenty-Seventh [HS94]. Two [CM98, SNK06]. two-level [SNK06]. Type [CMKH03]. Type-Driven [CMKH03].

UK [Ano00b, Fri94, HK95]. Undefined [BGvE^{+}97]. unified [HBD^{+}93]. uniform [KB94]. University [IEE95b]. Unix [Ano93a]. UNPACK [BR96]. Unstructured [MR95, PDS^{+}93, SM02b]. up-to-date [Din99]. Update [Ano92b]. USA [ACM07, ACM96a, ACM97, ACM98, Agr95, Ano94g, Ano94l, BBG^{+}95, Ban93, BGNP94, ERS96, HS94, Hua96, IEE94c, IEE96c, IEE02, Kar95, PBG^{+}95, SS96]. Use [HK93, HK95, Sul88]. User [CFPS94, EEV^{+}96, Iwa00, Smi95, PSC^{+}95]. User-friendly [CFPS94]. user-specified [PSC^{+}95]. Users [CKZ93]. Using [AHOK02, Ben99a, BCC^{+}96a, BCC^{+}96b, CLiN^{+}02, DL97b, KT00, LZ97, PFS^{+}04, PHD^{+}95, SZM98, SS97, TR96, AEG^{+}02, Ben00, BKK94, BID95, BCC^{+}97b, BCC^{+}97a, BCC^{+}97c, CFPS94, DDcMR96, DS97, Eli98, GHSJ94, Ogi02, SM02a, WO96]. utilities [Met99c]. Utrecht [Ano93l].

V2 [MAH^{+}02]. VA [Ano94b, Wie94]. Validation [BMV03]. Value [vHKS94]. Value-based [vHKS94]. Values [MCL^{+}95]. VAPP [BV94]. variable [BGvE^{+}97, Cou97]. Variable-Length [Cou97]. Variables [BGvE^{+}97, MCL^{+}95]. Varying [MCL^{+}95]. VAST [Van94b]. VAST-HPF [Van94b]. VAST/77toHPF [Van94b]. VAX [Sul88]. VAX-11 [Sul88]. Vector [BV94, Sab95]. Vectors [TR96]. Version [Sch97, Str94, Hig92, Hig94a, Hig94b, Hig94c, Hig94d]. versus [LC97]. Vesta [CFPS94]. VFC [BSSV98]. VI [Ano94a, BV94]. Via [FKKC96, EEV^{+}96, KK94, KK95b]. Vienna [BSSV98, CMZ92, UZCZ97, ZBC94]. Vienna-Fortran [UZCZ97]. Vienna-Fortran/HPF [UZCZ97]. Virginia [IEE94a]. virtual [BDLS96, Bod94, Ger98a, Ger98b]. viruses [Str94]. Visibility [BGvE^{+}97]. vision [Chw97]. Visual [DL97c, Lef98]. Visualization [HM96, HM98, KGV97, BCC^{+}97b, BCC^{+}97c]. visualizer [KC94]. Visualizing [KMS^{+}95]. Vol.II [HS94]. Volatile [BGvE^{+}97]. Volume [Ano93a, PSG03]. volumetric [MKF95]. VPP [AHOK02, ISKvW02]. vs [GS01].

Wailea [ERS96, HS94]. Wants [MCL^{+}95].

Washington [IEE94e]. Way [Del98].

WC1E [Eme94]. Weather

[RHH96, CDF^{+}93, GS95]. Welcomes [MCL^{+}95, Str94]. Well [MCL^{+}95]. WG10.3 [CGS94, DR94]. WG5 [Ano00b]. Where [MCL^{+}95]. wideband [MIN^{+}95]. within [PQ94]. without [CZM93, CMZ93, MCL^{+}95]. Work [PPW94]. Work-efficient [PPW94]. Working [Ano92b, Ano93k, CGS94, DR94]. Workshop [Agr95, Ano92b, Ano93i, CKZ93, DT94, Fer92, FK95, HK93, HK95, Kum94, PBG^{+}95, Sch93, Smi95, Wie94, Ban93, BGNP94, Don95, Hua96, Pra95]. Workstation [AOL94a, AOL94b, KC94]. Workstations [Coe94b, BID95, SR95]. world [SIOS02]. WWW [EEV^{+}96].

xHPF [DS97, Lev94, SS97].

Zosel [Eme94, UMM94, Rag95].

References

Andre:1996:NCT

[ABC^{+}96] F. Andre, P. Brezany, O. Cheron, W. Denissen, J.-L. Pazat, and K. Sanjari. A new compiler technology for handling HPF data parallel constructs. In Szymanski and Sinharoy [SS96], pages 279–282. ISBN 0-7923-9635-9. LCCN QA76.58.L37 1996.

Ancourt:1997:LAF

[ACIK97] Corinne Ancourt, Fabien Coelho, François Irigoin, and Ronan Keryell. Linear algebra framework for static High Performance Fortran code distribution. Scientific Programming, 6(1):3–27, Spring 1997. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

ACM:1993:PFA

[ACM93] ACM, editor. Proceedings of the Fourth ACM SIGPLAN Symposium on Principles and Practice of Parallel Programming, PPOPP: San Diego, California, May 19–22, 1993, volume 28(7) of ACM SIGPLAN Notices. ACM Press, New York, NY 10036, USA, July 1993. CO-DEN SINODQ. ISBN 0-89791-589-5. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.7 .S54 v.28:7.

ACM:1994:CPI

[ACM94] ACM, editor. Conference Proceedings. 1994 International Conference on Supercomputing. ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-665-4. LCCN ????

ACM:1995:SIC

[ACM95] ACM, editor. Supercomputing:

9th International conference — July 1995, Barcelona, Spain, Conference Proceedings of the International Conference on Supercomputing. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-728-6. LCCN ????

ACM:1996:FCP

[ACM96a] ACM, editor. FCRC ’96:

Conference proceedings of the 1996 International Conference on Supercomputing: Philadelphia, Pennsylvania, USA, May 25–28, 1996. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

ACM:1996:SCP

[ACM96b] ACM, editor. Supercomputing ’96 Conference Proceedings: November 17–22, Pittsburgh, PA. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-89791-854-1. LCCN ????

URL http://www.supercomp.
org/sc96/proceedings/. ACM

Order Number: 415962, IEEE Computer Society Press Order Number: RS00126.

ACM:1997:SHP

[ACM97] ACM, editor. SC’97: High Performance Networking and Computing: Proceedings of the 1997 ACM/IEEE SC97 Conference: November 15–21, 1997, San Jose, California, USA. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-89791-985-8. LCCN ????

URL http://www.supercomp.
org/sc97/proceedings/. ACM

SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

ACM:1998:SHP

[ACM98] ACM, editor. SC’98: High Performance Networking and Computing: Proceedings of the 1998 ACM/IEEE SC98 Conference: Orange County Convention Center, Orlando, Florida, USA, November 7–13, 1998. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/papers/.

ACM:2003:SII

[ACM03] ACM, editor. SC2003: Igniting Innovation. Phoenix, AZ, November 15–21, 2003. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 1-58113-695-1. LCCN ????

ACM:2007:TAS

[ACM07]  ACM, editor.  The Third

ACM SIGPLAN History of Programming Languages Conference (HOPL-III): proceedings: San Diego, California, USA, 9-10 June 2007. ACM Press, New York, NY 10036, USA, 2007. ISBN 1-59593-766-8. LCCN QA76.7 .H56 2007. URL

http://portal.acm.org/toc.
cfm?id=1238844.

Appelbe:1995:NAG

[ADH95] B. Appelbe, S. Doddapaneni, and C. Hardnett. A new algorithm for global optimization for parallelism and locality. In Pingali et al. [PBG^{+}95], pages 125–140. ISBN 3-540-58868-X. LCCN QA76.58 .W656 1994.

Akarsu:1996:PCS

[ADHF96] Erol Akarsu, Kivanc Dincer, Tomasz Haupt, and Geoffrey C. Fox. Particle-in-cell simulation codes in High Performance Fortran. In ACM [ACM96b], page ?? ISBN 0-89791-854-1. LCCN ???? URL

                                        http://www.supercomp.org/
                                        sc96/proceedings/SC96PROC/
                                        AKARSU/INDEX.HTM. ACM Or-

der Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Adve:1998:HPF

[Adv98] Vikram S. Adve. High Performance Fortran compilation techniques for parallelizing scientific codes. In ACM [ACM98], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        supercomp.org/sc98/papers/.

Agarwal:2002:FPN

[AEG^{+}02] R. C. Agarwal, R. F. Enenkel, F. G. Gustavson, A. Kothari, and M. Zubair. Fast pseudorandomnumber generators with modulus 2^{k} or 2^{k−1} using fused multiply–add. IBM Journal of Research and Development, 46 (1):97–116, January 2002. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

                                        tronic).  URL http://www.
                                        research.ibm.com/journal/
                                        rd/461/agarwal.html;   http:
                                        //www.research.ibm.com/journal/
                                        rd/461/agarwal.pdf.

Agrawal:1996:RSP

[AES^{+}96] G. Agrawal, G. Edjlali, A. Sussman, J. Humphries, and J. Saltz. Runtime support for programming in adaptive parallel environments. In Szymanski and Sinharoy [SS96], pages 241–252. ISBN 0-7923-9635-9. LCCN QA76.58.L37 1996.

Andre:1995:PDC

[AFMP95] F. Andre, M. Le Fur, Y. Maheo, and J.-L. Pazat. The Pandore data-parallel compiler and its portable runtime. In Hertzberger and Serazzi [HS95], pages 176–183. ISBN 3-540-59393-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Ayguade:1997:DRT

[AGG^{+}97] Eduard Ayguade, Jordi Garcia, Merce Girones, M. Luz Grande, and Jesus Labarta. DDT: a research tool for automatic data distribution in High Performance Fortran. Scientific Programming, 6(1):73–94, Spring 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Agrawal:1995:PIW

[Agr95] D. P. Agrawal, editor. Proceedings of the 1995 ICPP Workshop on Challenges for Parallel Processing, August 14, 1995, Raleigh, NC, USA. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1995. ISBN 0-8493-2618-4. LCCN QA76.58.I34 1995.

Asaoka:2002:EHJ

[AHOK02] Kae Asaoka, Akio Hirano, Yasuo Okabe, and Masanori Kanazawa. Evaluation of the HPF/JA extensions on Fujitsu VPP using the NAS parallel benchmarks. Lecture Notes in Computer Science, 2327:503– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/
service/series/0558/bibs/
2327/23270503.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/
2327/23270503.pdf.

Adve:2001:CAC

[AMC01] Vikram Adve and John Mellor-Crummey. Chapter 16. advanced code generation for high performance fortran. Lecture Notes in Computer Science, 1808:553–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-
ny.com/link/service/series/
0558/bibs/1808/18080553.htm;
 http://link.springer-ny.
com/link/service/series/0558/
papers/1808/18080553.pdf.

Araki:2002:OHP

[AMKS02] Takuya Araki, Hitoshi Murai, Tsunehiko Kamachi, and Yoshiki Seo. Optimization of HPF programs with dynamic recompilation technique.

Lecture Notes in Computer Science, 2327:551–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/
service/series/0558/bibs/
2327/23270551.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/
2327/23270551.pdf.

Anonymous:1992:PHP

[Ano92a] Anonymous. Publications: High Performance Fortran, books, periodicals. ACM Fortran Forum, 11(3):5–6, September 1992. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1992:SIC

[Ano92b]  Anonymous.  Standards in-

formation: Copies of Fortran Standard; interpretations; future meetings; national Fortran working groups; HPFF update; Fortran 90 implementation workshop, Nov 1992. ACM Fortran Forum, 12(2):2–5, June 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1992:SIFb

[Ano92c] Anonymous. Standards: Information; future meetings; international meetings; High Performance Fortran. ACM Fortran Forum, 11(2):2–5, June 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1993:CPR

[Ano93a] Anonymous. Constructing Predictable Real-Time System; Unix for Fortran Programmers; Unix Curses Explained; Internetworking with TCP/IP, Volume I: Principles, Protocols, and Architecture; System Performance Tuning; New Media. IEEE Software, 10(3):106–111, May 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:FFS

[Ano93b] Anonymous. Fortran 90 features in subset High Performance Fortran. ACM Fortran Forum, 12 (4):151–??, December 1, 1993. ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1993:GSH

[Ano93c] Anonymous. Goals and scope of High Performance Fortran. ACM Fortran Forum, 12(4):1– ??, December 1, 1993. ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1993:HPFa

[Ano93d] Anonymous. High Performance Fortran. Scientific Programming, 2(1-2):1–170, Spring– Summer 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:1993:HPFb

[Ano93e] Anonymous. High Performance Fortran language specification (part I). ACM Fortran Forum, 12(4):1–86, December 1, 1993. ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1993:HPFc

[Ano93f] Anonymous. High Performance Fortran terms and concepts. ACM Fortran Forum, 12(4): 169–??, December 1, 1993. ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1993:JD

[Ano93g] Anonymous. Journal of development. Scientific Programming, 2(1-2):A1–44, Spring–Summer 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:1993:NFH

[Ano93h]  Anonymous.  New features

in High Performance Fortran. ACM Fortran Forum, 12(4):3– ??, December 1, 1993. ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1993:PIW

[Ano93i] Anonymous, editor. Proceedings of International Workshop on Languages and Compilers for Parallel Computing. Min. Sci. and Technol, Jerusalem, Israel, 1993. ISBN ???? LCCN ????

Anonymous:1993:PSE

[Ano93j] Anonymous, editor. Proceedings. SHARE Europe Anniversary Meeting Client/Server — the Promise and the Reality: October 25-28, 1993, the Hague, the Netherlands. SHARE Europe, Carouge/Geneva, Switzerland, 1993. ISBN ???? ISSN 0254-6213. LCCN ????

Anonymous:1993:SIS

[Ano93k] Anonymous. Standards information: Speak your mind on Fortran 95; interpretations; organization of Standards Committees; future meetings; copies of Fortran 90 Standard; National Fortran working groups; High Performance Fortran language specification. ACM Fortran Forum, 12(3):1–5, September 1993. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1993:SEC

[Ano93l] Anonymous, editor. Supercomputing Europe ’93 conference papers: fifth international exhibition and conference highperformance computing, 22–24 February, 1993, Utrecht, The Netherlands. Royal Dutch Fairs, Utrecht, Netherlands, 1993. ISBN ???? LCCN ????

Anonymous:1994:AVP

[Ano94a] Anonymous, editor. ASPLOS-VI proceedings / Sixth International Conference on Architectural Support for Programming Languages and Operating Systems, San Jose, California, October 4–7, 1994, volume 29(11) of ACM SIGPLAN Notices. ACM Press, New York, NY 10036, USA, November 1994. CODEN SINODQ. ISBN 0-89791-660-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.9.A73 I558 1994.

Anonymous:1994:HPC

[Ano94b] Anonymous, editor. High performance computing and communications 1st Symposium High performance computing and communications symposium: – March 1994, Alexandria, VA. ARPA-CSTO, Arlington, VA, USA, 1994.

Anonymous:1994:HPFa

[Ano94c] Anonymous. High Performance Fortran language specification (part II). ACM Fortran Forum, 13(2):87–150, June 1, 1994. ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:HPFb

[Ano94d] Anonymous. High Performance Fortran language specification (part III). ACM Fortran Forum, 13(3):22–55, September 1, 1994. ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:HR

[Ano94e]  Anonymous.  HPF reports.

IEEE parallel and distributed technology: systems and applications, 2(3):71–??, Fall 1994. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Anonymous:1994:IPH

[Ano94f] Anonymous. Interpreting the performance of HPF/Fortran 90D. In IEEE [IEE94e], pages 743–752. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994.

URL http://sc94.ameslab.
gov/AP/contents.html. IEEE

catalog number 94CH34819.

Anonymous:1994:ISL

[Ano94g] Anonymous, editor. Lattice 93: proceedings of the International Symposium on Lattice Field Theory, Dallas Texas, USA, 12– 16 October 1993, volume 34 of Nuclear Physics B, Proceedings Supplements. North-Hol-land, Amsterdam, The Netherlands, April 1994. CODEN NPBSE7. ISBN ???? ISSN 0920-5632 (print), 1873-3832 (electronic). LCCN ????

Anonymous:1994:PLC

[Ano94h] Anonymous, editor. Plasmadynamics and lasers: 25th Conference — June 1994, Colorado Springs, CO, number 94 in Papers — American Institute of Aeronautics and Astronautics 1994; 2262. American Institute of Aeronautics and Astronautics, 370 L’Enfant Promenade SW, Washington, DC 20024– 2518, 1994.

Anonymous:1994:SIH

[Ano94i] Anonymous. Special issue: High Performance Fortran language specification. ACM Fortran Forum, 13(2), June 1, 1994. ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:SIO

[Ano94j] Anonymous. Standards information. organization of Standards Committees. ACM Fortran Forum, 13(1):4–??, March 1, 1994. ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:SIOa

[Ano94k] Anonymous. Standards information. organization of Standards Committees; Real-Time Fortran Standard; Fortran 90 interpretations; plans for Fortran 95; High Performance Fortran language specification. ACM Fortran Forum, 13(1):4–6, March 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:SPF

[Ano94l] Anonymous, editor. Symposium on Parallel Finite Element Computations: October 25-27, 1993, Supercomputer Institute, Minneapolis, MN, USA, volume 119(1–2) of Computer Methods in Applied Mechanics and Engineering. Elsevier, Amsterdam, The Netherlands, November 1994. CODEN CMMECC. ISBN ???? ISSN 0045-7825, 0374-2830. LCCN ????

Anonymous:1995:MJP

[Ano95a] Anonymous, editor. 186th Meeting: — June 1995, Pittsburgh, PA, volume 27(2) of Bulletin — American Astronomical Society. American Institute of Physics, Woodbury, NY, USA, 1995. ISSN 0002-7537.

Anonymous:1995:FAS

[Ano95b] Anonymous, editor. Fifth ACM SIGPLAN Symposium on Principles and Practice of Parallel Programming (PPoPP), Santa Barbara, CA, July 19–21, 1995, volume 30(8) of ACM SIGPLAN Notices. ACM Press, New York, NY 10036, USA, August 1995. CODEN SINODQ. ISBN ???? ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN ????

Anonymous:1996:SFP

[Ano96] Anonymous. Software: Fortran 90 for Power Macintosh; Fortran 90 and HPF for Linux (PC); Fortran-oriented periodicals; Fortran 90 information (April) from Michael Metcalf. ACM Fortran Forum, 15(1):15– 19, April 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1999:FSA

[Ano99]  Anonymous.  Fortran Stan-

dards activities: Organization of Standards committees; international meetings; the Fortran and HPF Standards. ACM Fortran Forum, 19(1):30–31, December 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2000:FSOa

[Ano00a] Anonymous. Fortran Standards: Organization of Standards committees; international meetings; national Fortran activities; the Fortran and HPF Standards. ACM Fortran Forum, 19(2):31– 32, August 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2000:FSOb

[Ano00b] Anonymous. Fortran Standards:

Organization of Standards committees; international meetings; national Fortran activities; the Fortran and HPF Standards; UK National Activity Report presented to the meeting of ISO/ IEC JTC1/SC22/WG5 (Fortran), Oulu, Finland, August 14–18, 2000. ACM Fortran Forum, 19(3):29–31, December 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2001:FSA

[Ano01a] Anonymous. Fortran Standards activities: Organization of Standards committees; international meetings; the Fortran and HPF Standards. ACM Fortran Forum, 20(2):8–9, August 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2001:FSOa

[Ano01b] Anonymous. Fortran Standards:

Organization of Standards committees; international meetings; national Fortran activities; the Fortran and HPF Standards. ACM Fortran Forum, 20(1):31– 32, April 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2001:FSOb

[Ano01c] Anonymous. Fortran Standards: Organization of Standards committees; international meetings; national Fortran activities; the Fortran and HPF Standards. ACM Fortran Forum, 20(3):31– 32, December 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2002:FSAb

[Ano02a]  Anonymous.  Fortran Stan-

dard activities: Organization of Standards committees; international meetings; national Fortran activities; the Fortran and HPF standards. ACM Fortran Forum, 21(3):30–31, December 2002. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2002:FSAa

[Ano02b] Anonymous. Fortran Standards activities: Organization of Standards committees; international meetings; national Fortran activities; the Fortran and HPF standards. ACM Fortran Forum, 21(1):31–32, April 2002. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

                                        URL ftp://softlib.rice.
                                        edu/pub/HPF/hpf-v20.ps.gz;
                                         http://www.crpc.rice.edu/
                                        HPFF/; http://www.fortran.
                                        com/fortran/;   http://www.
                                        iso.ch/; http://www.j3-fortran.
                                        org/;   http://www.nag.co.
                                        uk/sc22wg5/.

Anonymous:2003:FSA

[Ano03] Anonymous. Fortran Standard activities: Organization of Standards committees; international meetings; national Fortran activities; the Fortran and HPF standards. ACM Fortran Forum, 22(1):27–28, April 2003. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Annaratone:1994:DEC

[AOL94a] M. Annaratone, C. D. Offner, and D. B. Loveman. Digital Equipment Corporation — High Performance Fortran on workstation farms. In Siegel [Sie94b], pages 664–669. ISBN 0-8186-5602-6, 0-8186-5601-8. ISSN 1063-7133. LCCN QA 76.58 I56 1994.

Annaratone:1994:HPF

[AOL94b] M. Annaratone, C. D. Offner, and D. B. Loveman. High performance Fortran on workstation farms. In Siegal [Sie94a], pages 664–669. ISBN 0-8186-5602-6. LCCN QA76.58.I58 1994. IEEE catalog no. 94CH34819.

Amamiya:1994:RPL

[ASM^{+}94] M. Amamiya, M. Satoh, A. Makinouchi, K. Hagiwara, T. Yuasa, H. Aida, K. Ueda, K. Araki, T. Ida, and T. Baba. Research on programming languages for massively parallel processing. In Horiguchi et al. [HHK94], pages 443–450. ISBN 0-8186-6507-6. LCCN QA76.58 .I5673 1994. IEEE catalog no. 94TH0697-3.

Agrawal:1993:CRS

[ASS93] G. Agrawal, A. Sussman, and J. Saltz. Compiler and runtime support for structured and block structured applications. In IEEE [IEE93c], pages 578–587. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Agrawal:1995:IRC

[ASS95] Gagan Agrawal, Alan Sussman, and Joel Saltz. An integrated runtime and compile-time approach for parallelizing structured and block structured applications. IEEE Transactions on Parallel and Distributed Systems, 6(7):747–754, July 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.
computer.org/tpds/td1995/
l0747abs.htm.

Abdelrahman:1994:DAD

[AW94] T. S. Abdelrahman and T. N. Wong. Distributed array data management on NUMA multiprocessors. In IEEE [IEE94c], pages 551–559. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog no. 94TH0637-9.

Banerjee:1993:LCP

[Ban93] Utpal Banerjee, editor. Languages and compilers for parallel computing: 5th international workshop, New Haven, Connecticut, USA, August 3– 5, 1992: proceedings, volume 757 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-57502-2, 0-387-57502-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .L36 1993.

Boulet:1996:EAP

[BB96] P. Boulet and T. Brandes. Evaluation of automatic parallelization strategies for HPF compilers. Lecture Notes in Computer Science, 1067:778–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benkner:2002:EPP

[BB02] Siegfried Benkner and Thomas Brandes. Efficient parallel programming on scalable shared memory systems with High Performance Fortran. Concurrency and Computation: Practice and Experience, 14(8–9): 789–803, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/95016125/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=95016125{\&}PLACEBO=IE.
pdf.

Brandes:1998:CBH

[BBCR98] T. Brandes, F. Bregier, M. C.

Counilh, and J. Roman. Contribution to better handling of irregular problems in HPF2. Lecture Notes in Computer Science, 1470:639–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouchitte:1994:EAE

[BBDR94] V. Bouchitte, P. Boulet, A. Darte, and Y. Robert. Evaluating array expressions on massively parallel machines with communication/computation overlap. In Buchberger and Volkert [BV94], pages 713–724. ISBN 3-540-58430-7 (Berlin), 0-387-58430-7 (New York). LCCN QA76.58 .J65 1994.

Bouchitte:1995:EAE

[BBDR95] V. Bouchitte, P. Boulet, A. Darte, and Y. Robert. Evaluating array expressions on massively parallel machines with communication/computation overlap. International Journal of Supercomputer Applications and High Performance Computing, 9(3): 205–219, Fall 1995. CODEN IJSCFG. ISSN 0890-2720.

Bodin:1993:IPC

[BBG^{+}93] F. Bodin, P. Beckman, D. Gannon, S. Yang, S. Kesavan, A. Malony, and B. Mohr. Implementing a parallel C++ runtime system for scalable parallel systems. In IEEE [IEE93c], pages 588–597. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Bailey:1995:PSS

[BBG^{+}95] D. H. Bailey, P. E. Bjorstad, J. R. Gilbert, M. V. Mascagni, R. S. Schreiber, H. D. Simon, V. J. Torczon, and L. T. Watson, editors. Proceedings of the Seventh SIAM Conference on Parallel Processing for Scientific Computing (15–17 February, 1995, San Francisco, CA, USA). SIAM Press, Philadelphia, PA, USA, 1995. ISBN 0-89871-344-7. LCCN QA76.58.S55 1995.

Benkner:1994:PAS

[BBZ94] S. Benkner, P. Brezany, and H. Zima. Processing array statements and procedure interfaces in the PREPARE High Performance Fortran compiler. In Fritzson [Fri94], pages 324–338. ISBN 0-387-57877-3 (New York), 3-540-57877-3 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C65 I52 1994.

Brandes:1996:HSIa

[BCC^{+}96a] T. Brandes, S. Chaumette, M.- C. Counilh, A. Darte, J. Mignot, F. Desprez, and J. Roman. HP-FIT: A set of integrated tools for the parallelization of applications using High Performance Fortran. Rapport de recherche 96-28, LIP ENS, Lyon, France, 1996.

Brandes:1996:HSIb

[BCC^{+}96b] T. Brandes, S. Chaumette, M.- C. Counilh, A. Darte, J. Mignot, F. Desprez, and J. Roman. HPFIT: A set of integrated tools for the parallelization of applications using high performance Fortran: Part I: HP-FIT and the Transtool environment’. In J. Dongarra and B. Tourancheau, editors, Third Workshop on Environments and Tools for Parallel Scientific Computing. SIAM Press, Philadelphia, PA, USA, 1996. ISBN ???? LCCN ????

Brandes:1997:HSIa

[BCC^{+}97a] T. Brandes, S. Chaumette, M. C. Counilh, J. Roman, A. Darte, F. Desprez, and J. C. Mignot. HPFIT: A set of integrated tools for the parallelization of applications using High Performance Fortran. Part I: HPFIT and the TransTOOL environment. Parallel Computing, 23(1–2):71–87, April 16, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/parco/cas_sub/
browse/browse.cgi?year=1997&
volume=23&issue=1-2&aid=1148.

Brandes:1997:HSI

[BCC^{+}97b] T. Brandes, S. Chaumette, M. C. Counilh, J. Roman, F. Desprez, and J. C. Mignot. HP-FIT: A set of integrated tools for the parallelization of applications using High Performance Fortran. PART II: Datastructure visualization and HPF extensions for irregular problems. Parallel Computing, 23 (1):89–105, April 1997. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Brandes:1997:HSIb

[BCC^{+}97c] T. Brandes, S. Chaumette, M. C. Counilh, J. Roman, F. Desprez, and J. C. Mignot. HP-FIT: A set of integrated tools for the parallelization of applications using High Performance Fortran. PART II: Datastructure visualization and HPF extensions for irregular problems. Parallel Computing, 23 (1–2):89–105, April 16, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/parco/cas_sub/
browse/browse.cgi?year=1997&
volume=23&issue=1-2&aid=1149.

Bozkus:1993:CDD

[BCF^{+}93a] Z. Bozkus, A. Choudhary, G. Fox, T. Haupt, and S. Ranka. Compiling distribution directives in a Fortran 90D compiler. In IEEE [IEE93b], pages 617–620. ISBN 0-8186-4222-X. LCCN QA 76.58 I42 1993. IEEE catalog no. 93TH0584-3.

Bozkus:1993:FCD

[BCF^{+}93b] Z. Bozkus, A. Choudhary, G. Fox, T. Haupt, and S. Ranka. Fortran 90D/HPF compiler for distributed memory MIMD computers: Design, implementation, and performance results. In IEEE [IEE93c], pages 351–360. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Bozkus:1993:FHC

[BCF^{+}93c] Zeki Bozkus, Alok Choudhary, Geoffrey Fox, Tomasz Haupt, and Sanjay Ranka. Fortran 90D/HPF compiler for distributed memory MIMD computers: Design, implementation, and performance results. In IEEE [IEE93c], pages 351–360. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Bozkus:1994:SLF

[BCF^{+}94a] Z. Bozkus, A. Choudhary, G. Fox, T. Haupt, S. Ranka, R. Thakur, and Jhy-Chun Wang. Scalable libraries for Fortran 90D/High Performance Fortran. In IEEE [IEE94d], pages 67–76. ISBN 0-8186-4980-1. LCCN QA76.58.S34 1993.

Bozkus:1994:CFD

[BCF^{+}94b] Z. Bozkus, A. Choudhary, G. Fox, T. Haupt, S. Ranka, and Min-You Wu. Compiling Fortran 90D/HPF for distributed memory MIMD computers. Journal of Parallel and Distributed Computing, 21 (1):15–26, April 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bozkus:1994:CAF

[BCF^{+}94c] Zeki Bozkus, Alok Choudhary, Geoffrey Fox, Tomasz Haupt, and Sanjay Ranka. A compilation approach for Fortran 90D/ HPF compilers. In Banerjee et al. [BGNP94], pages 200–

215. ISBN 3-540-57659-2. LCCN QA76.58 .W656 1993.

Bozkus:1994:CFH

[BCF^{+}94d] Zeki Bozkus, Alok Choudhary, Geoffrey Fox, Tomasz Haupt, Sanjay Ranka, and Min-You Wu. Compiling Fortran 90D/ HPF for distributed memory MIMD computers. Journal of Parallel and Distributed Computing, 21(1):15–26, April 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic).  URL http://www.
idealibrary.com/links/doi/
10.1006/jpdc.1994.1039/production;
http://www.idealibrary.com/
links/doi/10.1006/jpdc.1994.
1039/production/pdf.

Bozkus:1993:CAF

[BCFH93] Z. Bozkus, A. Choudhary, G. Fox, and T. Haupt. A compilation approach for Fortran 90D/HPF compilers. In Banerjee [Ban93], pages 200–

215. ISBN 3-540-57502-2, 0-387-57502-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .L36 1993.

Babb:1993:RHP

[BCM^{+}93] R. Babb II, A. Choudhary, L. Meadows, S. Nakamoto, and V. J. Schuster. Retargetable high performance Fortran compiler challenges. In IEEE [IEE93a], pages 137–146. ISBN 0-8186-3400-6. LCCN QA75.5.C58 1993. IEEE catalog no. 93CH3251-6.

Brandes:1996:IPC

[BD96] T. Brandes and F. Desprez. Implementing pipelined computation and communication in an HPF compiler. Lecture Notes in Computer Science, 1123:459–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bode:1996:PVM

[BDLS96] Arndt Bode, Jack Dongarra, T. Ludwig, and V. Sunderam, editors. Parallel virtual machine, EuroPVM ’96: third European PVM conference, Munich, Germany, October 7–9, 1996: proceedings, volume 1156 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Blackford:1998:IGD

[BDPW98] L. S. Blackford, J. J. Dongarra, C. A. Papadopoulos, and R. C. Whaley. Installation guide and design of the HPF 1.1 interface to ScaLAPACK, SLHPF. LAPACK Working Note 137, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, August 1998.

URL http://www.netlib.org/
lapack/lawns/lawn137.ps;
http://www.netlib.org/lapack/
lawnspdf/lawn137.pdf. UT-

CS-98-396, August 1998.

Benkner:1998:HHP

[Ben98] S. Benkner. HPF+: High Performance Fortran for advanced industrial applications. Lecture Notes in Computer Science, 1401:797–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benkner:1999:OIH

[Ben99a] S. Benkner. Optimizing irregular HPF applications using halos. Lecture Notes in Computer Science, 1586:1015– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benkner:1999:HHP

[Ben99b] Siegfried Benkner. HPF+: High Performance Fortran for advanced scientific and engineering applications. Future Generation Computer Systems, 15(3): 381–??, ???? 1999. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Benkner:2000:OIH

[Ben00] Siegfried Benkner. Optimizing irregular HPF applications using halos. Concurrency: Practice and Experience, 12 (2-3):137–155, February–March 2000. CODEN CPEXEI. ISSN 1040-3108. URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract/72504938/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=72504938&PLACEBO=IE.pdf.

Berthou:2001:COH

[BF01] Jean-Yves Berthou and Eric Fayolle. Comparing OpenMP, HPF, and MPI programming: a study case. The International Journal of High Performance Computing Applications, 15(3):297–309, Fall 2001. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:/

                                        /journals.sagepub.com/doi/
                                        pdf/10.1177/109434200101500307.

Bogucz:1994:PEH

[BFHH94] E. A. Bogucz, G. C. Fox, T. Haupt, and K. A. Hawick. Preliminary evaluation of High-Performance Fortran as a language for computational fluid dynamics. In Anonymous [Ano94h], page ALL.

Brandes:1996:RHI

[BG96] T. Brandes and D. Greco. Realization of an HPF interface to ScaLAPACK with redistributions. Lecture Notes in Computer Science, 1067:834– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Botsford:1994:PCI

[BGG^{+}94] J. Botsford, A. Gawman, M. Gentleman, E. Kidd, K. Lyons, and J. Slonim, editors. Proceedings. CASCON ’94. Integrated Solutions: Toronto, Ontario, Canada, 31 October–3 November 1994. Nat. Res. Council Canada, Ottawa, Ont., Canada, 1994. ISBN ???? LCCN ????

Brezany:1992:CFOa

[BGMZ92] Peter Brezany, Michael Gerndt, Piyush Mehrotra, and Hans Zima. Concurrent file operations in a high performance FOR-TRAN. In IEEE [IEE92b], pages 230–237. ISBN 0-8186-2630-5. LCCN QA76.5 .S894 1992. IEEE catalog no. 92CH3216-9.

Banerjee:1994:LCP

[BGNP94] Uptal Banerjee, D. Gelernter, A. Nicolau, and D. Padua, editors. Languages and compilers for parallel computing: 6th international workshop, Portland, Oregon, USA, August 12–14, 1993: proceedings, volume 768 of Lecture notes in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-57659-2. LCCN QA76.58 .W656 1993.

Barkley:1997:CNB

[BGvE^{+}97] John Barkley, Roman Grzonka, Axel vom Endt, David Muxworthy, Tim Zeisloft, Michael Metcalf, Van Snyder, Robert Corbett, Lawrie Schonfelder, Dave Bailey, Pierre Hugonnet, Charles Knoebel, Clifford Blair, Loren P. Meissner, Phillip Helbig, Dirk Thorsten Vogel, Ronald Sverdlove, Dick Hendrickson, Malcom Cohen, Roger Young, James Giles, Larry Rolison, Richard Maine, Henry Zongarro, Petros Dafniotis, Jing Guo, David C. P. LaFrance-Linden, Harvey Richardson, Walt Brainerd, Patrice Lignelet, Jean Vezina, David Vowles, Clive Page, Viggo Norum, Neil N. Carlson, John Reid, Mike Lijewski, William F. Mitchell, and Robert Ferrell. Captured on the Net: Brian Meek (obituary); honors for Betty Holberton; significant new applications for Fortran 90 and Fortran 95?; why don’t academic snobs teach Fortran? is object-oriented Fortran a bad idea? status of High Performance Fortran; replacement for data statements; array of pointers; blanks in short form read and print; undefined pointer association status; visibility of implied do variable in array constructor; efficiency of loops; generic linked lists; module variables and the save attribute; modules and libraries; standardize Subset Fortran?; “mathematically equivalent” expressions, “volatile” variables, and optimization; pointers, targets, and allocatable arrays as function results. ACM Fortran Forum, 16(2):5–17, August 1997. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Branca:1995:CBH

[BID95] A. Branca, M. Ianigro, and A. Distante. A comparison between HPF and PVM for data parallel algorithms on a cluster of workstations using a high speed network. In Hertzberger and Serazzi [HS95], pages 930– 931. ISBN 3-540-59393-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Bixby:1994:ADL

[BKK94] R. Bixby, K. Kennedy, and U. Kremer. Automatic data layout using 0-1 integer programming. In Cosnard et al. [CGS94], pages 111–122. CO-DEN ITATEC. ISBN 0-444-81926-6. ISSN 0926-5473. LCCN QA76.58 .I46 1994.

Benson:1995:DDP

[BLLWW95] Edward G. Benson, David C. P.

LaFrance-Linden, Richard A. Warren, and Santa Wiryaman. Design of Digital’s Parallel Software Environment. Digital Technical Journal, 7(3): 24–38, Fall 1995. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.com/
pub/Digital/info/DTJ/v7n3/
Design_of_Digitals_Parallel_
S_02jan1996DTJJ02P8.ps;
http://www.digital.com:80/
info/DTJJ02/DTJJ02AH.HTM;
 http://www.digital.com:80/
info/DTJJ02/DTJJ02P8.PS;
http://www.digital.com:80/
info/DTJJ02/DTJJ02PF.PDF;

                                         http://www.digital.com:80/
                                        info/DTJJ02/DTJJ02SC.TXT.

Becker:1994:TPI

[BLT94] M. Becker, L. Litzler, and M. Tehel, editors. Transputers ’94: [advanced research and industrial applications]: proceedings of the international conference, September 21–23, 1994, Saline Royale d’Arc et Senans, France. IOS Press, Postal Drawer 10558, Burke, VA 2209-0558, USA, 1994. ISBN 90-5199-179-7 (IOS Press), 4-274-90003-7 (Ohmsha). LCCN ????

Brezany:2002:PSH

[BLW02] Peter Brezany, Jonghyun Lee, and Marianne Winslett. Parallel I/O support for HPF on computational grids. Lecture Notes in Computer Science, 2327:539– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/series/0558/bibs/
                                        2327/23270539.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/series/0558/papers/
                                        2327/23270539.pdf.

Benkner:1999:HPS

[BLZ99] S. Benkner, G. Lonsdale, and H. P. Zima. The HPF+ project: Supporting HPF for advanced industrial applications. Lecture Notes in Computer Science, 1685:??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bozkus:1994:TCE

[BMMN94] Z. Bozkus, L. Meadows, D. Miles, and S. Nakamoto. Techniques for compiling and executing HPF programs on shared memory and distributed memory parallel systems. In Kumar [Kum94], pages 515–520. ISBN 0-07-462332-X. LCCN QA 76.58 I587 1994.

Bozkus:1995:TCE

[BMMN95] Z. Bozkus, L. Meadows, D. Miles, and S. Nakamoto. Techniques for compiling and executing HPF programs on shared memory and distributed memory parallel systems. In Prasanna [Pra95], pages 515–520. ISBN 0-07-462332-X. LCCN ????

Bozkus:1995:CHP

[BMN^{+}95] Z. Bozkus, L. Meadows, S. Nakamoto, V. Schuster, and M. Young. Compiling High Performance Fortran. In Bailey et al. [BBG^{+}95], pages 704–709. ISBN 0-89871-344-7. LCCN QA76.58.S55 1995.

Bozkus:1997:POH

[BMN^{+}97] Zeki Bozkus, Larry Meadows, Steven Nakamoto, Vincent Schuster, and Mark Young. PGHPF — an optimizing High Performance Fortran compiler for distributed memory machines. Scientific Programming, 6(1):29–40, Spring 1997. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Briguglio:2003:PPM

[BMV03] Sergio Briguglio, Beniamino Di Martino, and Gregorio Vlad. A performance-prediction model for PIC applications on clusters of symmetric multiprocessors: Validation with hierarchical HPF + OpenMP implementation. Scientific Programming, 11(2):159–176, 2003. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bodin:1994:DPP

[Bod94]  François Bodin.  Directions

in parallel programming HPF, shared virtual memory and object parallelism in pC++. NASA contractor report NASA CR-194943; ICASE report no. 94-54, Institute for Computer Applications in Science and Engineering, NASA Langley Research Center, Hampton, VA, USA, 1994.

Boulter:1995:PEH

[Bou95] B. Boulter. Performance evaluation of HPF for scientific computing. In Hertzberger and Serazzi [HS95], pages 652–657. ISBN 3-540-59393-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Benkner:1997:OHA

[BP97] S. Benkner and M. Pantano. Optimizing HPF for advanced applications. Supercomputer, 13 (2):31–43, ???? 1997. CODEN SPCOEL. ISSN 0168-7875.

Berthou:1998:PHM

[BP98] J.-Y. Berthou and L. Plagne.

Parallel HPF-MPI implementation of the TBSCM Poisson solver. Lecture Notes in Computer Science, 1401:252– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bae:1996:PUC

[BR96] Seungjo Bae and Sanjay Ranka.

PACK/UNPACK on coarsegrained distributed memory parallel machines. Journal of Parallel and Distributed Computing, 38(2):204–216, November 1, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

idealibrary.com/links/doi/
10.1006/jpdc.1996.0141/production;
http://www.idealibrary.com/
links/doi/10.1006/jpdc.1996.
0141/production/pdf.

Boulet:1998:CPH

[BR98] P. Boulet and X. Redon. Communication pre-evaluation in HPF. Lecture Notes in Computer Science, 1470:263–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandes:1994:EHF

[Bra94a] T. Brandes. Evaluation of High-Performance Fortran on some real applications. Lecture Notes in Computer Science, 797:417– 422, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandes:1994:EHP

[Bra94b] T. Brandes. Evaluation of High Performance Fortran on some real applications. In Gentzsch and Harms [GH94a], pages 417–422. ISBN 3-540-57980-X (Berlin), 0-387-57980-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1–2 (c1994).

Brandes:2000:HLC

[Bra00] T. Brandes. HPF library and compiler support for halos in data parallel irregular computations. Parallel Processing Letters, 10(2/3):189–??, September 2000. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic). URL http://

ejournals.wspc.com.sg/ppl/
10/1002_03/S0129626400000196.
html.

Brezany:1992:CFOb

[Bre92]  Peter Brezany.  Concurrent

file operations in a high performance FORTRAN. NASA contractor report CR-189711; ICASE report no. 92-46, National Aeronautics and Space Administration, Langley Research Center; National Technical Information Service, distributor, Hampton, VA, USA, 1992. i + 15 pp.

Brieger:2000:HOO

[Bri00] Leesa Brieger. HPF to OpenMP on the Origin2000: a case study. Concurrency: Practice and Experience, 12(12):1147–1154, October 2000. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/76500351/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76500351&PLACEBO=IE.pdf.

Brezany:1995:PIC

[BSCV95] P. Brezany, K. Sanjari, O. Cheron, and E. Van Konijnenburg. Processing irregular codes containing arrays with multidimensional distributions by the PREPARE HPF compiler. In Hertzberger and Serazzi [HS95], pages 526–531. ISBN 3-540-59393-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Benkner:1998:PIA

[BSSV98] S. Benkner, K. Sanjari, V. Sipkova, and B. Velkov. Parallelizing irregular applications with the Vienna HPF+ compiler VFC. Lecture Notes in Computer Science, 1401:816– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buchberger:1994:PPC

[BV94] Bruno Buchberger and Jens Volkert, editors. Parallel processing: CONPAR 94-VAPP VI: third Joint International Conference on Vector and Parallel Processing, Linz, Austria, September 6–8, 1994: proceedings. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994.

ISBN 3-540-58430-7 (Berlin), 0-387-58430-7 (New York). LCCN QA76.58 .J65 1994.

Borowiec:1998:LHP

[BV98] J. Borowiec and A. Veen. Lowering HPF procedure interface to a canonical representation. Lecture Notes in Computer Science, 1366:192–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brezany:2001:GIP

[BxCW01] P. Brezany, P. x. Czerwinski, and M. Winslett. A generic interface for parallel access to large data sets from HPF applications. Future Generation Computer Systems, 17(8):977– 985, June 2001. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Brandes:1994:ATT

[BZ94] T. Brandes and F. Zimmermann. Adaptor — A transformation tool for HPF programs. In Decker and Rehmann [DR94], pages 91–96. ISBN 0-8176-5090-3. LCCN QA76.58 .P767 1994.

Benkner:1999:CHP

[BZ99] S. Benkner and H. Zima. Compiling High Performance Fortran for distributed-memory architectures. Parallel Computing, 25(13):1785–1825, December 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Coelho:1996:OCH

[CA96] Fabien Coelho and Corinne Ancourt. Optimal compilation of HPF remappings. Journal of Parallel and Distributed Computing, 38(2):229– 236, November 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic).  URL http://www.
idealibrary.com/links/doi/
10.1006/jpdc.1996.0143/production;
http://www.idealibrary.com/
links/doi/10.1006/jpdc.1996.
0143/production/pdf.

Chen:1994:CEC

[CC94] Dong-Yuan Chen and M. C.

Chen. CPAR-language extensions to C for irregular and adaptive parallel computations. In IEEE [IEE94c], pages 501–

508. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog no. 94TH0637-9.

Chen:2004:EPI

[CCW04] Min-Bin Chen, Tyng-Ruey Chuang, and Jan-Jan Wu. Efficient parallel implementations of near Delaunay triangulation with High Performance Fortran. Concurrency and Computation: Practice and Experience, 16(12): 1143–1159, October 2004. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Clemencon:1996:THM

[CDD^{+}96] C. Clemencon, K. M. Decker, V. R. Deshpande, A. Endo, J. Fritscher, P. A. R. Lorenzo, N. Masuda, A. Muller, R. Ruhl, W. Sawyer, B. J. N. Wylie, and F. Zimmermann. Toolssupported HPF and MPI parallelization of the NAS parallel benchmarks. In IEEE [IEE96b], pages 309–318. ISBN 0-8186-7551-9. LCCN QA76.58 .S95 1996. IEEE catalog number 96TB100062.

Chrisochoides:1993:MDH

[CDF^{+}93] N. Chrisochoides, K. Droegemeier, G. Fox, K. Mills, and Ming Xue. A methodology for developing high performance computing models: storm-scale weather prediction. In Tentner [Ten93], pages 82–89. ISBN 1-56555-052-8. LCCN ????

Cownie:1994:PPP

[CDH^{+}94] J. Cownie, A. Dunlop, S. Hellberg, A. J. G. Hey, and D. Pritchard. Portable parallel programming environments — the ESPRIT PPPE project. In Dekker et al. [DSZ94], pages 135–142. ISBN 0-444-81784-0. LCCN QA76.58.E98 1994.

Clemencon:1995:AEP

[CEF^{+}95] C. Clemencon, A. Endo, J. Fritscher, A. Muller, R. Ruhl, and B. J. N. Wylie. The ’annai’ environment for portable distributed parallel programming. In El-Rewini and Shriver [ERS95], pages 242–251 (vol. 2). ISBN 0-8186-6935-7. LCCN ????

Celmaster:1996:MFR

[Cel96] William N. Celmaster. Modern Fortran revived as the language of scientific parallel computing. Digital Technical Journal, 8(3): 39–45, December 1996. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.com/
pub/Digital/info/DTJ/v8n3/
Modern_Fortran_Revived_as_
the_07jan1997DTJN03P8.ps;
http://www.digital.com:80/
info/DTJN03/DTJN03AH.HTM;
http://www.digital.com:80/
info/DTJN03/DTJN03HM.HTM;
http://www.digital.com:80/
info/DTJN03/DTJN03P8.PS;
http://www.digital.com:80/
info/DTJN03/DTJN03PF.PDF;
http://www.digital.com:80/
info/DTJN03/DTJN03SC.TXT.

Chandy:1994:IST

[CFK^{+}94] Mani Chandy, Ian Foster, Ken Kennedy, Charles Koelbel, and Chau-Wen Tseng. Integrated support for task and data parallelism. International Journal of Supercomputer Applications, 8(2):80–98, Summer 1994. CO-DEN IJSAE9. ISSN 0890-2720.

Clemencon:1995:IRD

[CFMR95] C. Clemencon, J. Fritscher, M. J. Meehan, and R. Ruhl. An implementation of race detection and deterministic replay with MPI. In Haridi et al. [HAM95], pages 155–166. ISBN 3-540-60247-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I553 1995.

Corbett:1994:UEP

[CFPS94] P. F. Corbett, D. G. Feitelson, J.-P. Prost, and M. Snir.

User-friendly and efficient parallel I/Os using the Vesta parallel file system. In Becker et al. [BLT94], pages 23–38. ISBN 90-5199-179-7 (IOS Press), 4-274-90003-7 (Ohmsha). LCCN ????

Chatterjee:1993:GLA

[CGL^{+}93] Siddhartha Chatterjee, John R. Gilbert, Fred J. E. Long, Robert Schreiber, and Shang-Hua Teng. Generating local addresses and communication sets for dataparallel programs. ACM SIG-PLAN Notices, 28(7):149–158, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chatterjee:1995:GLA

[CGL^{+}95] Siddhartha Chatterjee, John R. Gilbert, Fred J. E. Long, Robert Schreiber, and Shang-Hua H. Teng. Generating local addresses and communication sets for data-parallel programs. Journal of Parallel and Distributed Computing, 26(1): 72–84, April 1, 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic).  URL http://www.
idealibrary.com/links/doi/
10.1006/jpdc.1995.1049/production;
 http://www.idealibrary.com/
links/doi/10.1006/jpdc.1995.
1049/production/pdf.

Cosnard:1994:PAC

[CGS94] Michel Cosnard, Guang R. Gao, and Gabriel M. Silberman, editors. Parallel architectures and compilation techniques: proceedings of the IFIP WG10.3 Working Conference on Parallel Architectures and Compilation Techniques, PACT ’94, Montreal, Canada, 24–26 August, 1994, volume A-50 of IFIP Transactions. A. Computer Science and Technology. North-Hol-land, Amsterdam, The Netherlands, 1994. CODEN ITATEC. ISBN 0-444-81926-6. ISSN 0926-5473. LCCN QA76.58 .I46 1994.

Chatterjee:1994:ADH

[CGSS94] S. Chatterjee, J. R. Gilbert, R. Schreiber, and T. J. Sheffler. Automatic distribution in HPF. In Dongarra and Tourancheau [DT94], pages 11– 18. ISBN 0-89871-343-9. LCCN QA76.58.I568 1994.

Cheng:1994:PDP

[CH94] Doreen Cheng and Robert Hood. A portable debugger for parallel and distributed programs. In IEEE [IEE94e], pages 723–732. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994.

URL http://sc94.ameslab.
gov/AP/contents.html. IEEE

catalog number 94CH34819.

Chweh:1997:NPF

[Chw97] Crystal Chweh. New products: Free C++ template library; templex sees vision of optical RAM; craft-y HPF compiler; getting graphic; realtime vision processing. IEEE Concurrency, 5(3):81–82, July/ September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

                                        URL http://dlib.computer.
                                        org/pd/books/pd1997/pdf/p3081.
                                        pdf.

Chung:1994:OPE

[CKMU94] J. S. Chung, K. Karal, H. Maeda, and Y. Ueda, editors. Offshore and polar engineering: 4th International conference — April 1994, Osaka, Japan. ISOPE, Golden, 1994. ISBN 1-880653-12-5, 1-880653-10-9. LCCN ????

Choudhary:1993:HPF

[CKZ93] Alok Choudhary, Charles Koelbel, and Mary Zosel. High Performance Fortran: Implementor and users workshop. In IEEE [IEE93c], pages 610–614. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Chandra:1997:HFG

[CL97a] S. Chandra and J. R. Larus. HPF on fine-grain distributed shared memory: Early experience. Lecture Notes in Computer Science, 1239:450–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chandra:1997:OCH

[CL97b] Satish Chandra and James R. Larus. Optimizing communication in HPF programs on finegrain distributed shared memory. ACM SIGPLAN Notices, 32(7):100–111, July 1997. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cai:2002:TDE

[CLiN^{+}02] DongSheng Cai, Yaoting Li, Ken ichi Nishikawa, Chiejie Xiao, and Xiaoyan Yan. Threedimensional electromagnetic particle-in-cell code using High Performance Fortran on PC cluster. Lecture Notes in Computer Science, 2327:515– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/
service/series/0558/bibs/
2327/23270515.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/
2327/23270515.pdf.

Chapman:1998:OHI

[CM98] B. Chapman and P. Mehrotra.

OpenMP and HPF: Integrating two paradigms. Lecture Notes in Computer Science, 1470:650– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cahir:2000:PMM

[CMK00] Margaret Cahir, Robert Moench, and Alice E. Koniges. Programming models and methods. In Koniges [Kon00], chapter 3, pages 27–54. ISBN 1-55860-540-1. LCCN QA76.58 .I483 2000. Discusses PVM, MPI, SHMEM, High-Performance Fortran, and POSIX threads.

Chauhan:2003:ATD

[CMKH03] Arun Chauhan, Cheryl McCosh, Ken Kennedy, and Richard Hanson. Automatic type-driven library generation for telescoping languages. In ACM [ACM03], page ?? ISBN 1-58113-695-1.

                                        LCCN ???? URL http://www.
                                        sc-conference.org/sc2003/
                                        inter_cal/inter_cal_detail.
                                        php?eventid=10692#1; http:
                                        //www.sc-conference.org/sc2003/
                                        paperpdfs/pap296.pdf.

Calzarossa:1998:ICS

[CMMP98] M. Calzarossa, L. Massari, A. Merlo, and M. Pantano. Integration of a compilation system and a performance tool: The HPF+ approach. Lecture Notes in Computer Science, 1401:809– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calzarossa:2001:PIH

[CMT01] M. Calzarossa, L. Massari, and D. Tessera. Performance issues of an HPF-like compiler. Future Generation Computer Systems, 18(1):147–156, September 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chapman:1994:SAM

[CMVZ94] B. Chapman, P. Mehrotra, J. Van Rosendale, and H. Zima. A software architecture for multidisciplinary applications: integrating task and data parallelism. In Buchberger and Volkert [BV94], pages 664–676. ISBN 3-540-58430-7 (Berlin), 0-387-58430-7 (New York). LCCN QA76.58 .J65 1994.

Chapman:1992:PVF

[CMZ92] Barbara M. Chapman, Piyush Mehrotra, and Hans P. Zima. Programming in Vienna Fortran. Scientific Programming, 1(1):31–50, Fall 1992. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Chapman:1993:HPFb

[CMZ93] B. M. Chapman, P. Mehrotra, and H. P. Zima. High Performance Fortran without templates: an alternative model for distribution and alignment. In ACM [ACM93], pages 92– 101. CODEN SINODQ. ISBN 0-89791-589-5. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.7 .S54 v.28:7.

Chapman:1994:HPF

[CMZ94a] B. Chapman, P. Mehrotra, and H. Zima. High performance Fortran languages: advanced applications and their implementation. In Gentzsch and Harms [GH94a], pages 407–416. ISBN 3-540-57980-X (Berlin), 0-387-57980-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1–2 (c1994).

Chapman:1994:EHAa

[CMZ94b] Barbara Chapman, Piyush Mehrotra, and Hans Zima. Extending HPF for advanced data parallel applications. NASA contractor report NASA CR-194913; ICASE report no. 94-34, Institute for Computer Applications in Science and Engineering, NASA Langley Research Center, Hampton, VA, USA, 1994.

Chapman:1995:HPF

[CMZ95] Barbara Chapman, Piyush Mehrotra, and Hans Zima. High Performance Fortran languages: Advanced applications and their implementation. Future Generation Computer Systems, 11(4– 5):401–407, August 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cabitza:1996:EHS

[CNBB96] G. Cabitza, C. Nardone, C. Bagaini, and A. Balzano. Experiences with HPF for scientific applications. Lecture Notes in Computer Science, 1067:290– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coelho:1994:CIC

[Coe94a] F. Coelho. Compilation of I/O communications for HPF. In IEEE [IEE94a], pages 102–109. ISBN 0-8186-6965-9. LCCN QA76.58.S95 1994. IEEE catalog no. 95TH8024.

Coelho:1994:EHC

[Coe94b] F. Coelho. Experiments with HPF compilation for a network of workstations. In Gentzsch and Harms [GH94a], pages 423–428. ISBN 3-540-57980-X (Berlin), 0-387-57980-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1–2 (c1994).

Coelho:1996:DHD

[Coe96] F. Coelho. Discussing HPF design issues. Lecture Notes in Computer Science, 1123:571–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Counihan:1997:FIF

[Cou97] Martin Counihan. Fortran 95:

Including Fortran 90, Details of High Performance Fortran (HPF), and the Fortran Module for Variable-Length Character Strings. UCL Press, January 1997. ISBN 1-85728-367-8. US$37.95. URL http:/

/www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
1857283678.

Crooks:1994:ADD

[CP94] P. Crooks and R. H. Perrott. Automatic data distribution. In Gentzsch and Harms [GH94a], pages 463–468. ISBN 3-540-57980-X (Berlin), 0-387-57980-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1–2 (c1994).

Chapman:1993:HPFa

[CZM93] Barbara Chapman, Hans Zima, and Piyush Mehrotra. High Performance Fortran without templates an alternative model for distribution and alignment. NASA contractor report NASA CR-191451; ICASE report 93-17, National Aeronautics and Space Administration, Langley Research Center; National Technical Information Service, distributor, Hampton, VA, USA, 1993. ???? pp.

Chapman:1994:EHAb

[CZM94] Barbara Chapman, Hans Zima, and Piyush Mehrotra. Extending HPF for advanced dataparallel applications. IEEE parallel and distributed technology: systems and applications, 2(3):59–70, Fall 1994. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Chapman:1994:HFL

[CZVM94] B. Chapman, H. Zima, Vienna, and P. Mehrotra. High-Performance Fortran languages: Advanced applications and their implementation. Lecture Notes in Computer Science, 797:407– 416, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiMartino:1998:PPS

[DBVS98] B. Di Martino, S. Briguglio, G. Vlad, and P. Sguazzero. Parallel plasma simulation in High Performance Fortran. Lecture Notes in Computer Science, 1401:203–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiMartino:1998:PTB

[DCBC98] B. Di Martino, M. Celino, M. Briscolini, and L. Colombo. Parallel tight-binding molecular dynamics code based on integration of HPF and optimized parallel libraries. Lecture Notes in Computer Science, 1541:104– 111, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiMartino:1999:HPF

[DCR99a] B. Di Martino, M. Celino, and V. Rosato. A[n] High Performance Fortran implementation of a tight-binding molecular dynamics simulation. Computer Physics Communications, 120 (2):255–268, August 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

DiMartino:1999:HPM

[DCR99b] B. Di Martino, M. Celino, and V. Rosato. HPF parallelization of a molecular dynamics code: Strategies and performances. Lecture Notes in Computer Science, 1593:535– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Darte:1996:TRT

[DDcMR96] A. Darte, F. Desprez, J. c. Mignot, and Y. Robert. TransTOOL: A restructuring tool for the parallelization of applications using High Performance Fortran. Journal of the Brazilian Computer Society, 3(2):5– 12, ???? 1996. ISSN 0104-6500 (print), 1678-4804 (elec-

                                        tronic). URL http://www.cs.
                                        utk.edu/~yrobert/journals/
                                        transtool.ps.gz.

Delves:1998:HPL

[Del98] Mike Delves. HPF: Programming Linux clusters the easy way. Linux Journal, 45:??, January 1998. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL

                                        ftp://ftp.ssc.com/pub/lj/
                                        listings/issue45/2432.tgz.

Ding:1998:EHP

[Din98] C. H. Q. Ding. Evaluations of HPF for practical scientific algorithms on T3E. Lecture Notes in Computer Science, 1401:223– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ding:1999:HPF

[Din99] Chris H. Q. Ding. High Performance Fortran for practical scientific algorithms: An up-to-date evaluation. Future Generation Computer Systems, 15(3): 343–352, April 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

DeSturler:1997:IIS

[DL97a] E. De Sturler and D. Loher. Implementing iterative solvers for irregular sparse matrix problems in High Performance Fortran. Lecture Notes in Computer Science, 1336:293–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeSturler:1997:PSI

[DL97b] E. De Sturler and D. Loher. Parallel solution of irregular, sparse matrix problems using high performance fortran. Lecture Notes in Computer Science, 1225:360– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dekeyser:1997:HBV

[DL97c] J.-L. Dekeyser and C. Lefebvre.

HPF-Builder: A visual environment to transform Fortran 90 codes to HPF. International Journal of Supercomputer Applications and High Performance Computing, 11(2):95–102, Summer 1997. CODEN IJSCFG. ISSN 1078-3482.

deMaine:1985:TPL

[dLD85] P. A. D. de Maine, S. Leong, and C. G. Davis. A transportable programming language (TPL) system. I overview. International Journal of Computer and Information Sciences, 14(3): 161–182, June 1985. CODEN IJ-CIAH. ISSN 0091-7036.

Dongarra:1995:HPC

[Don95] J. J. Dongarra, editor. High performance computing: technology, methods and applications: Advanced workshop — June 1994, Cetraro, Italy, volume 10 of Advances in Parallel Computing — Amsterdam. Elsevier, Amsterdam, The Netherlands, 1995. ISBN 0-444-82163-5. ISSN 0927-5452. LCCN QA76.88 .H54 1995.

Dowd:1993:HPC

[Dow93] Kevin Dowd. High Performance Computing. RISC architectures, optimization and benchmarks; A Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1993. ISBN 1-56592-032-5. xxv + 371 pp. LCCN QA76.88 .D6 1993; QA76.9.A73 D68 1993. US$25.95. URL http:

//www.oreilly.com/catalog/
9781565920323.

Dion:1994:PCW

[DPR94] M. Dion, J.-L. Philippe, and Y. Robert. Parallelizing compilers: what can be achieved? In Gentzsch and Harms [GH94a], pages 447–456. ISBN 3-540-57980-X (Berlin), 0-387-57980-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1–2 (c1994).

Decker:1994:PEM

[DR94] K. M. Decker and R. M. Rehmann, editors. Programming environments for massively parallel distributed systems: working conference of the IFIP WG10.3, April 25–29, 1994, Ascona, Italy. Birkhäuser Boston Inc., Cambridge, MA, USA, 1994. ISBN 0-8176-5090-3. LCCN QA76.58 .P767 1994.

Diaz:2003:DIP

[DRST03] Manuel Díaz, Bartolomé Rubio, Enrique Soler, and José M. Troya. Domain interaction patterns to coordinate HPF tasks. Parallel Computing, 29(7):925– 951, July 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

DeSturler:1997:SPH

[DS97] Eric De Sturler and Volker Strumpen. Scientific programming with High Performance Fortran: a case study using the xHPF compiler. Scientific Programming, 6(1):127–152, Spring 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Denissen:2002:FPB

[DS02] Will Denissen and Henk J.

Sips. Finding performance bugs with the TNO HPF benchmark suite. Concurrency and Computation: Practice and Experience, 14(8–9):691–712, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/95016124/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=95016124{\&}PLACEBO=IE.
pdf.

deSturler:1998:PIS

[dSL98] E. de Sturler and D. Loher. Parallel iterative solvers for irregular sparse matrices in High Performance Fortran. Future Generation Computer Systems, 13 (4):315–325, March 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Das:1994:SAI

[DSvH94] R. Das, J. Saltz, and R. von Hanxleden. Slicing analysis and indirect accesses to distributed arrays. In Banerjee et al. [BGNP94], pages 152–168. ISBN 3-540-57659-2. LCCN QA76.58 .W656 1993.

Dekker:1994:MPP

[DSZ94] L. (Leendert) Dekker, W. Smit, and J. C. Zuidervaart, editors. Massively parallel processing applications and development: proceedings of the 1994 EUROSIM Conference on Massively Parallel Processing Applications and Development, Delft, The Netherlands, 21–23 June 1994. Elsevier, Amsterdam, The Netherlands, 1994. ISBN 0-444-81784-0. LCCN QA76.58.E98 1994.

Dongarra:1994:ETP

[DT94] J. J. Dongarra and B. Tourancheau, editors. Environments and tools for parallel scientific computing: 2nd Workshop — May 1994, Townsend, TN, Proceedings of the Workshop on Environments and Tools for Parallel Scientific Computing. SIAM Press, Philadelphia, PA, USA, 1994. ISBN 0-89871-343-9. LCCN QA76.58.I568 1994.

Duval:1992:TPP

[Duv92] D. Duval. Trends in parallel programming models for high performance computers. In Ferenczi [Fer92], page 33. ISBN ???? LCCN ???? KFKI preprint/report series: KFKI-1992-34/M,N.

Delves:1998:HPF

[DZ98] M. Delves and H. Zima. High Performance Fortran: A status report or: Are we ready to give up MPI? Lecture Notes in Computer Science, 1497:161– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ellis:1996:CNF

[EEV^{+}96] Miles Ellis, Bernd Eggen, Arne Vajhoej, Emili Besalu, and Steve Lionel. Captured on the net: Final Fortran 95 Draft Stan-

dard via ftp; comp-fortran-90

mailbase list now on WWW; HPF home page on WWW; user notes on Fortran programming on WWW; nested summation symbol; session notes for Fall 1995 DECUS DEC Fortran 90 session; ACM groups deplore switch to C++ for advance placement. ACM Fortran Forum, 15(1):2–3, April 1996. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Ehold:1999:HNL

[EGKU99] H. J. Ehold, W. N. Gansterer, D. F. Kvasnicka, and C. W. Ueberhuber. HPF and numerical libraries. Lecture Notes in Computer Science, 1557:140– 152, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehold:2002:OLP

[EGKU02] H. J. Ehold, W. N. Gansterer, D. F. Kvasnicka, and C. W. Ueberhuber. Optimizing local performance in HPF. Parallel Computing, 28(3):415–432, March 2002. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Elisseev:1998:PTD

[Eli98] V. V. Elisseev. Parallelization of three-dimensional spectral laser–plasma interaction code using High Performance Fortran. Computers in Physics, 12 (2):173–??, March 1998. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.
org/doi/10.1063/1.168630.

Emerson:1994:BRH

[Eme94] D. R. Emerson. Book review: The high performance Fortran handbook, Scientific and engineering computation series: Charles H. Koelbel, David B. Loveman, Robert S. Schreiber, Guy L. Steele Jr. and Mary E. Zosel, M.I.T. Press, Fitzroy House, 11, Chenies Street, London WC1E 7ET, England, ISBN 0-262-61094-9, January 1994, Price: $24.95 (Paperback). Journal of Computational and Applied Mathematics, 54(3):N4, October 20, 1994.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0377042794902631.

El-Rewini:1995:PTH

[ERS95] H. El-Rewini and B. D. Shriver, editors. Proceedings of the Twenty-Eighth Hawaii International Conference on System Sciences. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-6935-7. LCCN ????

El-Rewini:1996:PTN

[ERS96] Hesham El-Rewini and Bruce D.

Shriver, editors. Proceedings of the Twenty-Ninth Hawaii International Conference on System Sciences (HICSS-29): Wailea, HI, USA, 3–6 January 1996. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7324-9. ISSN 1060-3425. LCCN ???? Five volumes.

Feibus:1994:SP

[Fei94] Andy Feibus. Standards in progress. Open Systems Today, ??(163):20–??, November 1994. ISSN 1061-0839.

Ferenczi:1992:AWT

[Fer92] S. Ferenczi, editor. Proceedings of the 1st Austrian-Hungarian Workshop on Transputer Applications: October 8–10, 1992, Sopron, Hungary. Hungarian Academy of Sciences Central Research Institute for Physics, Budapest, Hungary, 1992. ISBN ???? LCCN ???? KFKI preprint/report series: KFKI-1992-34/M,N.

Faber:2001:IAG

[FGL01] Peter Faber, Martin Griebl, and Christian Lengauer. Issues of the automatic generation of HPF loop programs. Lecture Notes in Computer Science, 2017:359– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/
service/series/0558/bibs/
2017/20170359.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/
2017/20170359.pdf.

Fahringer:2000:PMH

[FGRT00] Thomas Fahringer, Michael Gerndt, Graham Riley, and Jesper Larsson Träff. On performance modeling for HPF applications with ASL. Lecture Notes in Computer Science, 1940:191– ??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/
service/series/0558/bibs/
1940/19400191.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/
1940/19400191.pdf.

Ferenczi:1995:PAW

[FK95] Szabolcs Ferenczi and Peter Kacsuk, editors. Proceedings of the 2nd Austrian-Hungarian Workshop on Transputer Applications (KFKI-1995-2/M, N), September 29–October 1, 1994, Budapest, Hungary, volume 2 of Hungarian Academy of Sciences Central Research Institute for Physics — Publications, KFKI //M,N. Hungarian Acad. Sci, Budapest, Hungary, 1995. ISBN ???? LCCN ????

Foster:1996:MCL

[FKK96a] I. T. Foster, D. R. Kohr, Jr., and R. Krishnaiyer. MPI as a coordination layer for communicating HPF tasks. In IEEE [IEE96c], pages 68–78. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Foster:1996:CDT

[FKK^{+}96b] I. T. Foster, D. R. Kohr, Jr., R. Krishnaiyer, Choudhary, and A. Communicating data-parallel tasks: an MPI library for HPF. In IEEE [IEE96a], pages 433– 438. ISBN 0-8186-7557-8. LCCN QA76.88.I575 1996. IEEE catalog number 96TB100074.

Foster:1996:DSB

[FKKC96] Ian Foster, David R. Kohr, Jr., Rakesh Krishnaiyer, and Alok Choudhary. Double standards: Bringing task parallelism to HPF via the message passing interface. In ACM [ACM96b], page ?? ISBN 0-89791-854-1. LCCN ???? URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
FOSTER2/INDEX.HTM. ACM Or-

der Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Foster:1994:TPH

[Fos94] Ian Foster. Task parallelism and high-performance languages. IEEE parallel and distributed technology: systems and applications, 2(3):27–36, Fall 1994. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Fox:1991:DHP

[Fox91] G. Fox. Draft High Performance Fortran Language Specification: High Performance Fortran Forum, CRPC-TR92225, November 1992. Computer Information Technical, June 1991. ISBN 99922-4-636-7. US$50.00. URL http:/

                                        /www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        9992246367.

Fox:1994:APS

[Fox94] G. Fox. The application perspective for scalable data and task parallel languages HPF and HPC++. In Anonymous [Ano94b], pages 445–457.

Freeman:1992:PNA

[FP92] T. L. (Len) Freeman and C. (Christopher) Phillips. Parallel numerical algorithms. Prentice Hall International Series in Computer Science. Prentice-Hall International, Englewood Cliffs, NJ 07632, USA, 1992. ISBN 0-13-651597-5. xii + 315 pp. LCCN QA76.9.A43 F74 1992. US$40.00. Chapter 5 discusses HPF and PVM.

Fritzson:1994:CCI

[Fri94] Peter A. Fritzson, editor. Compiler construction: 5th International Conference, CC ’94, Edinburgh, UK, April 7–9, 1994: proceedings, volume 786 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 0-387-57877-3 (New York), 3-540-57877-3 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C65 I52 1994.

Fahringer:2002:SAS

[FSPC^{+}02] T. Fahringer, K. Sowa-Piekło, P. Czerwiński, P. Brezany, M. Bubak, R. Koppler, and R. Wismüller. SPiDER — An advanced symbolic debugger for Fortran 90/HPF programs. Concurrency and Computation: Practice and Experience, 14(2):103–136, February 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/91513538/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=91513538&PLACEBO=IE.pdf.

Foster:1994:CSI

[FXAC94] I. Foster, Ming Xu, B. Avalani, and A. Choudhary. A compilation system that integrates High Performance Fortran and Fortran M. In IEEE [IEE94c], pages 293–300. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog no. 94TH0637-9.

Gupta:1994:CEC

[GB94] M. Gupta and P. Banerjee. Compile-time estimation of communication costs of programs. Journal of Programming Languages, 2(3):191–225, September 1994. CODEN JPLAER. ISSN 0963-9306.

Gerndt:1998:HLP

[Ger98a] Michael Gerndt. High-level programming of massively parallel computers based on shared virtual memory. Parallel Computing, 24(3–4):383–400, May 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/parco/sub/1998/24/3-
                                        4/1295.pdf.

Gerndt:1998:HPM

[Ger98b] Michael Gerndt. High-level programming of massively parallel computers based on shared virtual memory. Parallel Computing, 24(3-4):383–400, May 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gawman:1993:PC

[GGK^{+}93] Ann Gawman, W. Morven Gentleman, E. Kidd, Per-Åke Larson, and J. Slonim, editors. Proceedings CASCON ’93: Toronto, Ontario, Canada, 24–28 October 1993. Nat. Res. Council of Canada, Ottawa, Ont., Canada, 1993. ISBN ???? LCCN QA76.76.S64 C378 1993 v.1-2. Two volumes.

Gentzsch:1994:HCNa

[GH94a] Wolfgang Gentzsch and Uwe Harms, editors. High-Perfor-mance computing and networking: International Conference and Exhibition, Munich, Germany, April 18–20, 1994: proceedings, volume 796 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-57980-X (Berlin), 0-387-57980-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1–2 (c1994).

Gentzsch:1994:HCNb

[GH94b] Wolfgang Gentzsch and Uwe Harms, editors. High-performance computing and networking: International Conference and Exhibition, Munich, Germany, April 18–20, 1994: proceedings, volume 797 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-57981-8 (Berlin: vol. 2: paperback), 0-387-57981-8 (New York: vol. 2: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1–2 (c1994). DM96.00. Two volumes.

Gupta:1994:IFF

[GHSJ94] S. K. S. Gupta, C.-H. Huang, P. Sadayappan, and R. W. Johnson. Implementing fast Fourier transforms on distributed-memory multiprocessors using data redistributions. Parallel Processing Letters, 4(4):477–488, December 1994. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Gupta:1992:MGD

[GKH^{+}92] S. K. S. Gupta, S. D. Kaushik, C.-H. Huang, J. R. Johnson, R. W. Johnson, and P. Sadayappan. A methodology for generating data distributions to optimize communication. In IEEE [IEE92a], pages 436–441. ISBN 0-8186-3200-3. LCCN QA 76.58 I42 1992. IEEE catalog no. 92TH0492-9.

Gupta:1993:AGD

[GKH^{+}93] S. K. S. Gupta, S. D. Kaushik, C.-H. Huang, J. R. Johnson, R. W. Johnson, and P. Sadayappan. On the automatic generation of data distributions. In Schnabel [Sch93], page 82. CO-DEN SINODQ. ISBN ???? ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.7 .S54 v.28:1.

Germain:1997:HCS

[GLPE97] C. Germain, J. Laminie, M. Pallud, and D. Etiemble. An HPF case study of a domaindecomposition based irregular application. Lecture Notes in Computer Science, 1277:201– ??, 1997. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1995:HCI

[GMS^{+}95] Manish Gupta, Sam Midkiff, Edith Schonberg, Ven Seshadri, David Shields, Ko-Yang Wang, Wai-Mee Ching, and Ton Ngo. An HPF compiler for the IBM SP2. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL

http://www.supercomp.org/
sc95/proceedings/417_SAMM/

SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Goda:1993:HPF

[God93] O. Goda. High performance Fortran. Joho-Shori (J. Information Processing Soc. Japan), 34 (9):1179–1186, September 1993. CODEN JOSHA4. ISSN 0447-8053.

Gross:1994:TPH

[GOS94] Thomas Gross, David R.

O’Hallaron, and Jaspal Subhlok. Task parallelism in a High Performance Fortran framework. IEEE parallel and distributed technology: systems and applications, 2(3):16–26, Fall 1994. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Gustafsson:1995:PSH

[GS95] N. Gustafsson and D. Salmond. A parallel spectral HIRLAM with the data parallel programming model and with message passing — a comparison (weather forecasting). In Hoffmann and Kreitz [HK95], pages 32–48. ISBN 981-02-2211-4. LCCN QC866.E26 1994.

Gupta:1997:SAR

[GS97] M. Gupta and E. Schonberg. Static analysis to reduce synchronization costs of dataparallel programs with remote memory copy. Parallel Processing Letters, 7(2):145–156, June 1997. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Grelck:2001:HVS

[GS01] Clemens Grelck and Sven-Bodo Scholz. HPF vs. SAC – A case study (research note). Lecture Notes in Computer Science, 1900:620–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/series/0558/bibs/
                                        1900/19000620.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/series/0558/papers/
                                        1900/19000620.pdf.

Guo:2001:DSH

[Guo01] M. Guo. Denotational semantics of an HPF-like data-parallel language model. Parallel Processing Letters, 11(2–3):363–??, 2001. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Haridi:1995:EPP

[HAM95] Seif Haridi, Khayri Ali, and Peter Magnusson, editors. EURO-PAR ’95 parallel processing: First International EURO PAR Conference, Stockholm, Sweden, August 29–31, 1995: proceedings, number 966 in Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-60247-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I553 1995.

Hansen:1998:EHP

[Han98] Per Brinch Hansen. An evaluation of High Performance Fortran. ACM SIGPLAN Notices, 33(3):57–64, March 1998. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). The author strongly criticizes HPF, and remarks “HPF has severe conceptual and practical limitations.”.

Hatcher:1994:GEI

[Hat94] Philip J. Hatcher. Guest Editor’s introduction: The impact of High Performance Fortran. IEEE parallel and distributed technology: systems and applications, 2(3):13–??, Fall 1994. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Harris:1995:CHP

[HBB^{+}95] Jonathan Harris, John A. Bircsak, M. Regina Bolduc, Jill Ann Diewald, Israel Gale, Neil W. Johnson, Shin Lee, C. Alexander Nelson, and Carl D. Offner. Compiling High Performance Fortran for distributed-memory systems. Digital Technical Journal, 7 (3):5–23, Fall 1995. CODEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.com/
                                        pub/Digital/info/DTJ/v7n3_
                                        03_FORTRAN.pdf.

Hawick:1993:PUM

[HBD^{+}93] K. A. Hawick, R. S. Bell, A. Dickinson, P. D. Surry, and B. J. N. Wylie. Parallelisation of the unified model data assimilation scheme. In Hoffmann and Kauranne [HK93], pages 188– 203. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Huang:2008:FPM

[HC08] Jih-Woei Huang and Chih-Ping Chu. A flexible processor mapping technique toward data localization for block-cyclic data redistribution. The Journal of Supercomputing, 45(2): 151–172, August 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/openurl.
                                        asp?genre=article&issn=0920-
                                        8542&volume=45&issue=2&spage=

Hwang:2003:SAE

[HCLJ03] Gwan-Hwan Hwang, Cheng-Wei Chen, Jenq Kuen Lee, and Roy Dz-Ching Ju. Segmented alignment: An enhanced model to align data parallel programs of HPF. The Journal of Supercomputing, 25(1):17–41, May 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//ipsapp009.kluweronline.
com/content/getfile/5189/
43/2/abstract.htm;   http:
//ipsapp009.kluweronline.
com/content/getfile/5189/
43/2/fulltext.pdf.

Hayashi:1994:AAS

[HDH^{+}94] Kenichi Hayashi, Tsunehisa Doi, Takeshi Horie, Yoichi Koyanagi, Osamu Shiraki, Nobutaka Imamura, Toshiyuki Shimizu, Hiroaki Ishihata, and Tatsuya Shindo. AP1000+: architectural support of PUT/GET interface for parallelizing compiler. In Anonymous [Ano94a], pages 196–207. CODEN SIN-ODQ. ISBN 0-89791-660-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.9.A73 I558 1994. URL http://www.

acm.org:80/pubs/citations/
proceedings/asplos/195473/
p196-hayashi/.

Hayashi:1995:AAS

[HDH^{+}95] K. Hayashi, T. Doi, T. Horie, Y. Koyanagi, O. Shiraki, N. Imamura, T. Shimizu, H. Ishihata, and T. Shindo. AP1000+: architectural support for parallelizing compilers. Transactions of the Information Processing Society of Japan, 36(7):1680–1690, July 1995. CODEN JSGRD5. ISSN 0387-5806.

Hempel:1996:SMM

[Hem96] R. Hempel. The status of the MPI message-passing standard and its relation to PVM. In Bode et al. [BDLS96], pages 14–21. ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Hey:1994:GEP

[Hey94] Tony Hey. The Genesis Esprit project — An overview. Parallel Computing, 20(10–11):1605– 1612, November 3, 1994. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/parco/cas_sub/
                                        browse/browse.cgi?year=1994&
                                        volume=20&issue=10-11&aid=
                                        916.

Hawick:1995:EHP

[HF95] K. A. Hawick and G. C. Fox. Exploiting High Performance Fortran for computational fluid dynamics. In Hertzberger and Serazzi [HS95], pages 413–419. ISBN 3-540-59393-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Horiguchi:1994:ISP

[HHK94] S. Horiguchi, D. F. Hsu, and M. Kimura, editors. International Symposium on Parallel Architectures, Algorithms, and Networks (ISPAN): proceedings of the 1994, December 14–16, 1994, Kanazawa, Japan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6507-6. LCCN QA76.58 .I5673 1994. IEEE catalog no. 94TH0697-3.

Hall:1996:ICF

[HHKT96] Mary W. Hall, Seema Hiranandani, Ken Kennedy, and Chau-Wen Tseng. Interprocedural compilation of Fortran D. Journal of Parallel and Distributed Computing, 38(2):114– 129, November 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic).  URL http://www.
idealibrary.com/links/doi/
10.1006/jpdc.1996.0135/production;
http://www.idealibrary.com/
links/doi/10.1006/jpdc.1996.
0135/production/pdf.

HPFF:1992:HPF

[Hig92] High Performance Fortran Forum, ???? High Performance Fortran Language Specification, Version 0.4, 1992.

Forum:1993:HPF

[Hig93] High Performance Fortran Forum. High Performance Fortran language specification. Scientific Programming, 2(1–2):1– 170, Spring–Summer 1993. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

HPPF:1994:HPF

[Hig94a] High Performance Fortran Forum. High Performance Fortran language specification (part III), Version 1.0, May 1993. ACM Fortran Forum, 13(3):22– 55, September 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

HPPF:1994:SIHa

[Hig94b] High Performance Fortran Forum. Special issue: High Performance Fortran language specification, Version 1.0 (May 3, 1993): Part II. 5. intrinsic and library procedures. ACM Fortran Forum, 13(2):87–138, June 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

HPPF:1994:SIHb

[Hig94c] High Performance Fortran Forum. Special issue: High Performance Fortran language specification, Version 1.0 (May 3, 1993): Part II. 6. extrinsic procedures. ACM Fortran Forum, 13(2):139–142, June 1994. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

HPPF:1994:SIHc

[Hig94d] High Performance Fortran Forum. Special issue: High Performance Fortran language specification, Version 1.0 (May 3, 1993): Part II. 7. storage and sequence association. ACM Fortran Forum, 13(2):143–150, June 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Hu:2000:HHP

[HJJ^{+}00] Y. Charlie Hu, Guohua Jin, S. Lennart Johnsson, Dimitris Kehagias, and Nadia Shalaby. HPFBench: a High Performance Fortran benchmark suite. ACM Transactions on Mathematical Software, 26(1): 99–149, March 2000. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/
pubs/citations/journals/toms/
2000-26-1/p99-hu/;   http:
//www.acm.org/pubs/citations/
journals/toms/2000-26-1/p99-
hu/p99-hu.pdf.

Hu:1997:HPF

[HJT97] Y. Charlie Hu, S. Lennart Johnsson, and Shang-Hua Teng. High Performance Fortran for highly irregular problems. ACM SIGPLAN Notices, 32(7):13–24, July 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hoffmann:1993:PFE

[HK93] Geerd-R. Hoffmann and Tuomo Kauranne, editors. Proceedings of the Fifth ECMWF Workshop on the Use of Parallel Processors in Meteorology. Parallel Supercomputing in Atmospheric Science. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1993. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Hoffmann:1995:CAP

[HK95] Geerd-R. Hoffmann and Norbert Kreitz, editors. Coming of age: proceedings of the Sixth ECMWF Workshop on the Use of Parallel Processors in Meteorology, Reading, UK, November 21–25, 1994. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1995. ISBN 981-02-2211-4. LCCN QC866.E26 1994.

Hayder:1998:CPL

[HKM98] M. E. Hayder, D. E. Keyes, and P. Mehrotra. A comparison of PETSc library and HPF implementations of an archetypal PDS computation. Advances in Engineering Software, 29(3):415–423, April 1998. CO-DEN AESODT. ISSN 0045-7949 (print), 1879-2243 (electronic).

Hiranandani:1994:CTB

[HKMCS94] S. Hiranandani, K. Kennedy, J. Mellor-Crummey, and A. Sethi. Compilation techniques for block-cyclic distributions. In ACM [ACM94], pages 392–403. ISBN 0-89791-665-4. LCCN ????

Hiranandani:1992:CFD

[HKT92] Seema Hiranandani, Ken Kennedy, and Chau-Wen Tseng. Compiling Fortran D for MIMD distributed-memory machines.

Communications of the ACM, 35 (8):66–80, August 1992. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.
org/pubs/toc/Abstracts/0001-
0782/135230.html.

Hiranandani:1993:PEF

[HKT93a] Seema Hiranandani, Ken Kennedy, and Chau-Wen Tseng. Preliminary experiences with the Fortran D compiler. In IEEE [IEE93c], pages 338–350. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Hiranandani:1993:PEFb

[HKT93b] Seema Hiranandani, Ken Kennedy, and Chau-Wen Tseng. Preliminary experiences with the Fortran D compiler. In IEEE [IEE93c], pages 338–350. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Hiranandani:1994:DEN

[HKTW94] S. Hiranandani, K. Kennedy, Chau-Wen Tseng, and S. Warren. The D editor: a new interactive parallel programming tool. In IEEE [IEE94e], pages 733–742, 817. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994.

                                        URL http://sc94.ameslab.
                                        gov/AP/contents.html. IEEE

catalog number 94CH34819.

Hwang:2001:AOS

[HLJ01] G-H. Hwang, J. K. Lee, and R. D. C. Ju. Array operation synthesis to optimize HPF programs on distributed memory machines. Journal of Parallel and Distributed Computing, 61 (4):467–500, April 2001. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hackstadt:1996:DAQ

[HM96] S. T. Hackstadt and A. D. Malony. Distributed array query and visualization for High Performance Fortran. Lecture Notes in Computer Science, 1123:55–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hackstadt:1998:DDA

[HM98] Steven T. Hackstadt and Allen D. Malony. DAQV: Distributed Array Query and Visualization Framework. Theoretical Computer Science, 196 (1–2):289–317, April 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/tcs/sub/1998/196/1-
                                        2/2741.pdf.

Halatsis:1994:PPA

[HMPT94] C. Halatsis, D. Maritsas, G. Philokyprou, and S. Theodoridis, editors. PARLE ’94: parallel architectures and languages Europe: 6th International PARLE Conference, Athens, Greece, July 4–8, 1994: proceedings. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-58184-7. LCCN QA76.58 .I564 1994.

Hwang:1995:RLS

[HMS^{+}95] Yuan-Shin Hwang, Bongki Moon, S. D. Sharma, R. Ponnusamy, R. Das, and J. H. Saltz. Runtime and language support for compiling adaptive irregular programs on distributedmemory machines. Software —Practice and Experience, 25 (6):597–621, June 1995. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hesham:1994:PTH

[HS94] E.-R. Hesham and B. D. Shriver, editors. Proceedings of the Twenty-Seventh Hawaii International Conference on System Sciences. Vol.II: Software Technology, January 4–7, 1994, Wailea, HI, USA, volume 27. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5060-5. ISSN 1060-3425. LCCN ???? IEEE catalog no. 94TH0607-2.

Hertzberger:1995:HCN

[HS95] Bob Hertzberger and Giuseppe Serazzi, editors. High-Perfor-mance computing and networking: International Conference and Exhibition, Milan, Italy, May 3–5, 1995: proceedings, volume 919 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-59393-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Huang:1996:LCP

[Hua96] C.-H. Huang, editor. Languages and compilers for parallel computing: 8th international workshop, LCPC ’95, Columbus, Ohio, USA, August 10–12, 1995: proceedings, number 1033 in Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-60765-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .W656 1995.

Hellberg:1994:PPP

[HZ94] S. A. Hellberg and E. Zaluska. A portable parallel programming environment based around PCTE. Information and Software Technology, 36(7):419–425, July 1994. CODEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (electronic).

IEEE:1992:PFI

[IEE92a] IEEE, editor. Proceedings of the Fourth IEEE Symposium on Parallel and Distributed Processing, December 1–4, 1992, Arlington, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-3200-3. LCCN QA 76.58 I42 1992. IEEE catalog no. 92TH0492-9.

IEEE:1992:PSM

[IEE92b] IEEE, editor. Proceedings, Supercomputing ’92: Minneapolis, Minnesota, November 16– 20, 1992. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-2630-5. LCCN QA76.5 .S894 1992. IEEE catalog no. 92CH3216-9.

IEEE:1993:DPC

[IEE93a] IEEE, editor. Digest of papers:

Compcon spring ’93, San Francisco, California, February 22– 26, 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-3400-6. LCCN QA75.5.C58 1993. IEEE catalog no. 93CH3251-6.

IEEE:1993:PFI

[IEE93b] IEEE, editor. Proceedings of the Fifth IEEE Symposium on Parallel and Distributed Processing: December 1–4, 1993, Dallas, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-4222-X. LCCN QA 76.58 I42 1993. IEEE catalog no. 93TH0584-3.

IEEE:1993:PSP

[IEE93c] IEEE, editor. Proceedings, Supercomputing ’93: Portland, Oregon, November 15–19, 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

IEEE:1994:FSF

[IEE94a] IEEE, editor. Frontiers’95, the 5th Symposium on the Frontiers of Massively Parallel Computation: proceedings, February 6–9, 1995, McLean, Virginia. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6965-9. LCCN QA76.58.S95 1994. IEEE catalog no. 95TH8024.

IEEE:1994:IPN

[IEE94b]  IEEE, editor.  ICIP ’94:

proceedings, November 13–16, 1994, Austin Convention Center, Austin, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6952-7 (casebound), 0-8186-6950-0 (paperback), 0-8186-6951-9 (microfiche). LCCN TA1637.I25 1994. Three volumes. IEEE catalog no. 94CH35708.

IEEE:1994:PSH

[IEE94c] IEEE, editor. Proceedings of the Scalable High-Performance Computing Conference, May 23–25, 1994, Knoxville, Tennessee, USA, Scalable High Performance Computing Conference. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog no. 94TH0637-9.

IEEE:1994:PSP

[IEE94d] IEEE, editor. Proceedings of the Scalable Parallel Libraries Conference , October 6–8, 1993, Mississippi State, Mississippi. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-4980-1. LCCN QA76.58.S34 1993.

IEEE:1994:PSW

[IEE94e]  IEEE, editor.  Proceedings,

Supercomputing ’94: Washington, DC, November 14–18, 1994, Supercomputing. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994.

URL http://sc94.ameslab.
gov/AP/contents.html. IEEE

catalog number 94CH34819.

IEEE:1995:IIP

[IEE95a] IEEE, editor. IPPS ’95: 9th International parallel processing symposium — April 25– 28, 1995, Santa Barbara, CA, International Parallel Processing Symposium. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7074-6. ISSN 1063-7133. LCCN QA76.58 .I58 1995.

IEEE:1995:PSP

[IEE95b] IEEE, editor. Proceedings of the 1994 Scalable Parallel Libraries Conference: October 12– 14, 1994, Mississippi State University, Mississippi. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-6895-4. LCCN QA76.58 .S34 1994.

IEEE:1996:ICH

[IEE96a] IEEE, editor. 3rd International Conference on High Performance Computing: proceedings, December 19–22, 1996, Trivandrum, India. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7557-8. LCCN QA76.88.I575 1996. IEEE catalog number 96TB100074.

IEEE:1996:FSS

[IEE96b] IEEE, editor. Frontiers’96, the Sixth Symposium on the Frontiers of Massively Parallel Computation: October 27–31, 1996, Annapolis, Maryland: proceedings. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7551-9. LCCN QA76.58 .S95 1996. IEEE catalog number 96TB100062.

IEEE:1996:PSM

[IEE96c] IEEE, editor. Proceedings. Second MPI Developer’s Conference: Notre Dame, IN, USA, 1– 2 July 1996. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

IEEE:2002:STI

[IEE02] IEEE, editor. SC2002: From Terabytes to Insight. Proceedings of the IEEE ACM SC 2002 Conference, November 16–22, 2002, Baltimore, MD, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7695-1524-X. LCCN ????

Iwashita:2002:TLH

[IHKvW02] Hidetoshi Iwashita, Kohichiro Hotta, Sachio Kamiya, and Matthijs van Waveren. Towards a lightweight HPF compiler. Lecture Notes in Computer Science, 2327:526–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/
service/series/0558/bibs/

                                        2327/23270526.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/series/0558/papers/
                                        2327/23270526.pdf.

Ishizaki:1996:LPA

[IK96] K. Ishizaki and H. Komatsu. A loop parallelization algorithm for HPF compilers. Lecture Notes in Computer Science, 1033:176–190, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iwashita:2002:VFD

[ISKvW02] Hidetoshi Iwashita, Naoki Sueyasu, Sachio Kamiya, and Matthijs van Waveren. VPP Fortran and the design of HPF/JA extensions. Concurrency and Computation: Practice and Experience, 14(8–9):575–588, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract/95016131/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=95016131{\&}PLACEBO=IE.
                                        pdf.

Iwashita:2000:RHA

[Iwa00] Hidetoshi Iwashita. Report of HUG2000 [4th annual High Performance Fortran user group meeting, October 19–20, 2000, Tokyo, Japan]. ACM Fortran Forum, 19(3):21–22, December 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Jung:1992:HET

[JA92] C. C.-D. Jung and D. Advani. HiPPI exploitation in TCP/IP environment. In IEEE [IEE92b], pages 610–615. ISBN 0-8186-2630-5. LCCN QA76.5 .S894 1992. IEEE catalog no. 92CH3216-9.

Joisha:2001:ECO

[JB01] Pramod G. Joisha and Prithviraj Banerjee. The efficient computation of ownership sets in HPF. IEEE Transactions on Parallel and Distributed Systems, 12(8):769–788, August 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td2001/pdf/l0769.
pdf;   http://www.computer.
org/tpds/td2001/l0769abs.
htm.

Joubert:1995:FAH

[Jou95] A. Joubert. Financial applications and HPF. In Dongarra [Don95], pages 317–336. ISBN 0-444-82163-5. ISSN 0927-5452. LCCN QA76.88 .H54 1995.

Karin:1995:PAI

[Kar95] Sidney Karin, editor. Proceedings of the 1995 ACM/IEEE Supercomputing Conference, December 3–8, 1995, San Diego Convention Center, San Diego, CA, USA. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-89791-816-9. URL

                                        http://www.supercomp.org/
                                        sc95/proceedings/. These pro-

ceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Kasahara:1993:SSP

[Kas93] H. Kasahara. System software for parallel processing. Joho-Shori (J. Information Processing Soc. Japan), 34(9):1134– 1138, September 1993. CODEN JOSHA4. ISSN 0447-8053.

Kohn:1994:RPP

[KB94] S. R. Kohn and S. B. Baden. A robust parallel programming model for dynamic nonuniform scientific computations. In IEEE [IEE94c], pages 509– 517. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog no. 94TH0637-9.

Kennedy:1994:IIF

[KBKT94] J. G. Kennedy, M. Behr, V. Kalro, and T. E. Tezduyar. Implementation of implicit finite element methods for incompressible flows on the CM-5. In Anonymous [Ano94l], pages 95– 111. CODEN CMMECC. ISBN ???? ISSN 0045-7825, 0374-2830. LCCN ????

Klinker:1994:PPV

[KC94] Gudrun J. Klinker and I-Yu Chen. PerfVisS: a performance visualizer for high performance Fortran programs on workstation farms. Technical report series / Cambridge Research Laboratory CRL 94/4, Digital Equipment Corp., Cambridge Research Laboratory, Cambridge, MA, USA, June 30, 1994. 18. pp.

Kennedy:1994:CTM

[Ken94a] K. Kennedy. Compiler technology for machine-independent parallel programming. International Journal of Parallel Programming, 22(1):79–98, February 1994. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Kennedy:1994:PPS

[Ken94b]  K. Kennedy.  Parallel pro-

gramming support in Fortran D and High Performance Fortran. In Anonymous [Ano94b], pages 427–432.

Koppler:1997:VDD

[KGV97] Rainer Koppler, Siegfried Grabner, and Jens Volkert. Visualization of distributed data structures for High Performance Fortran-like languages. Scientific Programming, 6(1):115– 126, Spring 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kaushik:1994:ACD

[KHJS94] S. D. Kaushik, C.-H. Huang, R. W. Johnson, and P. Sadayappan. An approach to communication-efficient data redistribution. In ACM [ACM94], pages 364–373. ISBN 0-89791-665-4. LCCN ????

Kaushik:1995:MAR

[KHRS95] S. D. Kaushik, C.-H. Huang, J. Ramanujam, and P. Sadayappan. Multi-phase array redistribution: modeling and evaluation. In IEEE [IEE95a], pages 441–445. ISBN 0-8186-7074-6. ISSN 1063-7133. LCCN QA76.58 .I58 1995.

Kaushik:1995:IGI

[KHS95] S. D. Kaushik, C.-H. Huang, and P. Sadayappan. Incremental generation of index sets for array statement execution on distributed-memory machines. In Pingali et al. [PBG^{+}95], pages 251–265. ISBN 3-540-58868-X. LCCN QA76.58 .W656 1994.

Kaushik:1996:EIS

[KHS96] S. D. Kaushik, C.-H. Huang, and P. Sadayappan. Efficient index set generation for compiling HPF array statements on distributed-memory machines. Journal of Parallel and Distributed Computing, 38(2):237– 247, November 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic).  URL http://www.
idealibrary.com/links/doi/

10.1006/jpdc.1996.0144/production;
 http://www.idealibrary.com/
links/doi/10.1006/jpdc.1996.
0144/production/pdf.

Kwon:2012:HAO

[KJEM12] Okwan Kwon, Fahed Jubair, Rudolf Eigenmann, and Samuel Midkiff. A hybrid approach of OpenMP for clusters. ACM SIGPLAN Notices, 47(8):75– 84, August 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPOPP ’12 conference proceedings.

Kornkven:1994:EIH

[KK94] E. Kornkven and L. V. Kale. Efficient implementation of High Performance Fortran via adaptive scheduling — an overview. In Kumar [Kum94], pages 136–

141. ISBN 0-07-462332-X. LCCN QA 76.58 I587 1994.

Kennedy:1995:ADL

[KK95a] Ken Kennedy and Ulrich Kremer. Automatic data layout for High Performance Fortran. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/580_UKRE/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Kornkven:1995:EIH

[KK95b] E. Kornkven and L. V. Kale. Efficient implementation of High Performance Fortran via adaptive scheduling — an overview. In Prasanna [Pra95], pages 136– 141. ISBN 0-07-462332-X. LCCN ????

Kennedy:1998:ADL

[KK98] Ken Kennedy and Ulrich Kremer. Automatic data layout for distributed-memory machines. ACM Transactions on Programming Languages and Systems, 20(4):869–916, July 1998. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic).  URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/toplas/1998-20-4/
                                        p869-kennedy/.

Kennedy:2001:CHP

[KK01] Ken Kennedy and Charles Koelbel. Chapter 1. High Performance Fortran 2.0. Lecture Notes in Computer Science, 1808:3–??, 2001. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

                                        URL http://link.springer-
                                        ny.com/link/service/series/
                                        0558/bibs/1808/18080003.htm;
                                         http://link.springer-ny.
                                        com/link/service/series/0558/
                                        papers/1808/18080003.pdf.

Konovalav:1995:FDL

[KKMP95] N. A. Konovalav, V. A. Kryukov, S. N. Mikhailov, and A. A. Pogrebtsov. Fortran DVM: a language for portable parallel program development. Programming and Computer Software; translation of Programmirovaniye (Moscow, USSR) Plenum, 21(1):35–38, January–February 1995. CO-DEN PROGD3. ISSN 0132-3474, 0361-7688.

Kamachi:1995:HCP

[KKS^{+}95] Tsunehiko Kamachi, Kazuhiro Kusano, Kenji Suehiro, Yoshiki Seo, Masanori Tamura, Shoichi Sakon, Yukimitsu Watanabe, and Yukimasa Shiroto. HPF compiler for parallel computers: implementation and performance evaluation on Cenju-3. Nippon Electric Company research and development, 36(2): 325–334, April 1995. CODEN NECRAU. ISSN 0048-0436.

Kennedy:2007:RFH

[KKZ07] K. Kennedy, C. Koelbel, and H. Zima. The rise and fall of High Performance Fortran: An historical object lesson. In ACM [ACM07], pages 7.1–7.22. ISBN 1-59593-766-8. LCCN QA76.7 .H56 2007. URL http://

portal.acm.org/toc.cfm?id=
1238844.

Kennedy:2011:RFH

[KKZ11] Ken Kennedy, Charles Koelbel, and Hans Zima. The rise and fall of High Performance Fortran. Communications of the ACM, 54(11):74–82, November 2011. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kim:1996:PSS

[KMBK96] Taegeun Kim, Kyeongdeok Moon, Nanjoo Ban, and Jungkwon Kim. PPTran: Source to source translator for High Performance Fortran. Parallel Algorithms and Applications, 9(3–4):213–225, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a778707980.

Kamachi:1997:KPH

[KMR^{+}97] T. Kamachi, A. Muller, R. Ruhl, Y. Seo, K. Suehiro, and M. Tamura. Kemari: a portable High Performance Fortran system for distributed memory parallel processors. Scientific Programming, 6(1):41–58, Spring 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kimelman:1995:VEH

[KMS^{+}95] D. Kimelman, P. Mittal, E. Schonberg, P. F. Sweeney, Ko-Yang Wang, and D. Zernik. Visualizing the execution of High Performance Fortran (HPF) programs. In IEEE [IEE95a], pages 750– 759. ISBN 0-8186-7074-6. ISSN 1063-7133. LCCN QA76.58 .I58 1995.

Kalns:1995:DPD

[KN95a] E. T. Kalns and L. M. Ni. DaReL: a portable data redistribution library for distributedmemory machines. In IEEE [IEE95b], pages 78–87. ISBN 0-8186-6895-4. LCCN QA76.58 .S34 1994.

Kalns:1995:PMT

[KN95b] Edgar T. Kalns and Lionel M.

Ni. Processor mapping techniques toward efficient data redistribution. IEEE Transactions on Parallel and Distributed Systems, 6(12):1234–1247, December 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kennedy:1995:LAC

[KNS95a] K. Kennedy, N. Nedeljkovic, and A. Sethi. A linear-time algorithm for computing the memory access sequence in dataparallel programs. In Anonymous [Ano95b], pages 102–111. CODEN SINODQ. ISBN ???? ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN ????

Kennedy:1995:LTA

[KNS95b] Ken Kennedy, Nenad Nedeljkovic, and Ajay Sethi. A linear-time algorithm for computing the memory access sequence in dataparallel programs. ACM SIG-PLAN Notices, 30(8):102–111, August 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Koelbel:1992:OHP

[Koe92] Charles Koelbel. An overview of High Performance Fortran. ACM Fortran Forum, 11(4):9– ??, December 1, 1992. ISSN 1061-7264 (print), 1931-1311 (electronic).

Koelbel:1994:HPF

[Koe94] Charles H. Koelbel. The High Performance Fortran handbook. Scientific and engineering computation. MIT Press, Cambridge, MA, USA, 1994. ISBN 0-262-11185-3, 0-262-61094-9 (paperback). xiv + 329 pp.

Knies:1993:HPF

[KOM93] A. Knies, Michael (Michael A.) O’Keefe, and T. MacDonald. High Performance Fortran: a practical analysis. AHPCRC preprint 93-107, Army High Performance Computing Research Center, Minneapolis, MN, USA, 1993. ii + 18 pp.

Knies:1994:HPF

[KOM94] Allan Knies, Matthew O’Keefe, and Tom MacDonald. High Performance Fortran: A practical analysis. Scientific Programming, 3(3):187–199, Fall 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Koniges:2000:ISP

[Kon00] Alice E. Koniges, editor. Industrial Strength Parallel Computing. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2000. ISBN 1-55860-540-1. xxv + 597 pp. LCCN QA76.58 .I483

Kremer:1994:COR

[KR94] Ulrich Kremer and Marcelo Ramé. Compositional oil reservoir simulation in Fortran D: A feasibility study on Intel iPSC/

860. International Journal of Supercomputer Applications, 8 (2):119–128, Summer 1994. CO-DEN IJSAE9. ISSN 0890-2720. See erratum [KR95].

Kremer:1995:ECO

[KR95] U. Kremer and M. Ramé. Erratum: Compositional Oil Reservoir Simulation in Fortran D: A Feasibility Study on Intel iPSC/

860. International Journal of Supercomputer Applications and High Performance Computing, 9 (1):71, Spring 1995. CODEN IJSAE9. ISSN 0890-2720. See [KR94].

Kennedy:2002:SIH

[KS02] K. Kennedy and Y. Seo. Special issue: High Performance Fortran. Concurrency and Computation: Practice and Experience, 14(8–9):551–553, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/95016134/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=95016134{\&}PLACEBO=IE.
pdf.

Keppens:2000:UHP

[KT00] R. Keppens and G. Toth. Using High Performance Fortran for magnetohydrodynamic simulations. Parallel Computing, 26(6):705–722, May 2000. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kumar:1994:PPI

[Kum94] V. K. Prasanna Kumar, editor. Parallel processing: 1st IWWP: proceedings of the First International Workshop on Parallel Processing (IWPP-94), December 26–31, 1994, Bangalore, India. Tata McGraw-Hill Pub. Co, New Delhi, India, 1994. ISBN 0-07-462332-X. LCCN QA 76.58 I587 1994.

Kim:1994:CAM

[KW94] I. Kim and M. Wolfe. Communication analysis for multicomputer compilers. In Cosnard et al. [CGS94], pages 101–109. CODEN ITATEC. ISBN 0-444-81926-6. ISSN 0926-5473. LCCN QA76.58 .I46 1994.

Kolte:1995:ERA

[KW95] Priyadarshan Kolte and Michael Wolfe. Elimination of redundant array subscript range checks. ACM SIGPLAN Notices, 30(6):270–278, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
proceedings/pldi/207110/p270-
kolte/.

Luecke:1997:HPF

[LC97] G. R. Luecke and J. J. Coyle. High Performance Fortran versus explicit message passing on the IBM SP-2 for the parallel LU, QR, and Cholesky factorizations. Supercomputer, 13(2): 4–14, ???? 1997. CODEN SP-COEL. ISSN 0168-7875.

Lefebvre:1998:VEH

[Lef98] C. Lefebvre. Visual edition of HPF mappings with HPF-Builder. Lecture Notes in Computer Science, 1401:943– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levesque:1994:APR

[Lev94] John M. Levesque. Applied Parallel Research’s xHPF system. IEEE parallel and distributed technology: systems and applications, 2(3):71, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

LaFrance-Linden:1998:CDH

[LL98] David C. P. LaFrance-Linden.

Challenges in designing an HPF debugger. Digital Technical Journal, 9(3):50–??, ???? 1998. CODEN DTJOEL. ISSN 0898-901X. URL http:

//www.digital.com:80/info/
DTJR04/DTJR04AH.HTM;   http:
//www.digital.com:80/info/
DTJR04/DTJR04HM.HTM;   http:
//www.digital.com:80/info/
DTJR04/DTJR04P8.PS;   http:
//www.digital.com:80/info/
DTJR04/DTJR04PF.PDF;   http:
//www.digital.com:80/info/
DTJR04/DTJR04SC.TXT.

Lorenzo:1996:HPF

[LMMW96] P. A. R. Lorenzo, A. Mueller, Y. Murakami, and B. J. N. Wylie. High Performance Fortran interfacing to ScaLAPACK. Lecture Notes in Computer Science, 1184:457–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luksch:1997:SSE

[LMR^{+}97] Peter Luksch, Ursula Maier, Sabine Rathmayer, Matthias Weidmann, and Friedemann Unger. Sempa: Software engineering for parallel scientific computing. IEEE Concurrency, 5(3):64–72, July/ September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

                                        URL http://dlib.computer.
                                        org/pd/books/pd1997/pdf/p3064.
                                        pdf;   http://www.computer.
                                        org/concurrency/pd1997/p3064abs.
                                        htm.

Loveman:1993:HPF

[Lov93] D. B. Loveman. High performance Fortran. IEEE parallel and distributed technology: systems and applications, 1(1): 25–42, February 1993. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Loveman:1994:DHP

[Lov94] D. B. Loveman. The DEC High Performance Fortran 90 compiler front end. In IEEE [IEE94a], pages 46–53. ISBN 0-8186-6965-9. LCCN QA76.58.S95 1994. IEEE catalog no. 95TH8024.

Li:1993:ANL

[LP93] Wei Li and K. Pingali. Access normalization: loop restructuring for NUMA computers. ACM Transactions on Computer Systems, 11(4):353–375, November 1993. CODEN AC-SYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

LeFur:1995:APA

[LPA95] M. Le Fur, J.-L. Pazat, and F. Andre. An array partitioning analysis for parallel loop distribution. In Haridi et al. [HAM95], pages 351–364. ISBN 3-540-60247-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I553 1995.

Li:1997:EHC

[LZ97] Konming Gary Li and Nabil M.

Zamel. An evaluation of HPF compilers and the implementation of a parallel linear equation solver using HPF and MPI. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ???? URL

http://www.supercomp.org/
sc97/proceedings/TECH/LI/

INDEX.HTM. ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Murai:2002:IEH

[MAH^{+}02] Hitoshi Murai, Takuya Araki, Yasuharu Hayashi, Kenji Suehiro, and Yoshiki Seo. Implementation and evaluation of HPF/SX V2. Concurrency and Computation: Practice and Experience, 14(8–9): 603–629, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/95016132/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=95016132{\&}PLACEBO=IE.
                                        pdf.

Marquet:1993:LED

[Mar93] P. Marquet. Languages and expressions of data parallelism. Technique et Science Informatiques, 12(6):685–714, 1993. CO-DEN TTSIDJ. ISSN 0752-4072, 0264-7419.

Machacek:1995:HPF

[MB95] M. Machacek and E. Bertschinger. A High Performance Fortran PM code for cosmological Nbody simulations. In Anonymous [Ano95a], page 2.07. ISSN 0002-7537.

Merlin:1998:MDP

[MBFC98] J. H. Merlin, S. B. Baden, S. J. Fink, and B. M. Chapman. Multiple data parallelism with HPF and KeLP. Lecture Notes in Computer Science, 1401:828– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Merlin:1999:MDP

[MBFC99] John Merlin, Scott Baden, Stephen Fink, and Barbara Chapman. Multiple data parallelism with HPF and KeLP. Future Generation Computer Systems, 15(3):393–405, April 1, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Mellor-Crummey:2002:AOS

[MCAB^{+}02] J. Mellor-Crummey, V. Adve, B. Broom, D. Chavarría-Miranda, R. Fowler, G. Jin, K. Kennedy, and Q. Yi. Advanced optimization strategies in the Rice dHPF compiler. Concurrency and Computation: Practice and Experience, 14(8–9):741–767, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/95016130/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=95016130{\&}PLACEBO=IE.
pdf.

Merlin:1996:SSH

[MCH96] J. H. Merlin, D. B. Carpenter, and A. J. G. Hey. SHPF: a Subset High Performance Fortran compilation system. Fortran Journal, ??(??):??, March/ April 1996. ISSN 1060-0221.

Moss:1995:CNF

[MCL^{+}95] Leonard J. Moss, William B.

Clodius, Dave Lloyd, Michael Metcalf, Bernd R. Eggen, Chuck Ritz, Cliver Page, Marty Cohen, Emilio C. Lopes, and Jerrold L. Wagener. Captured on the Net: Fortran is alive and well and living in London; where to find Fortran 90 varying string module; formal grammars for Fortran 77 and Fortran 90; linking modules without MAKE; Fortran 90 efficient storage for triangular matrices; email group welcomes participants; Fortran Journal wants more articles; downloadable Fortran 77 textbook (free); pass values, not variables (objects) as arguments; ELF linker — not to be confused with Essential Lahey Fortran; High Performance Fortran Forum meeting; latest Fortran 95 draft available on line. ACM Fortran Forum, 14 (4):1–3, December 1995. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Mehrotra:1993:DPP

[Meh93a] P. Mehrotra. Data parallel programming: the promises and limitations of High Performance Fortran. In Volkert [Vol93], page 114. ISBN 3-540-57314-3, 0-387-57314-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I5 1993. DM58.00.

Mehrotra:1993:ILD

[Meh93b] P. Mehrotra. Invited lecture: Data parallel programming: The promises and limitations of High Performance Fortran. In Volkert [Vol93], pages 114–?? ISBN 3-540-57314-3, 0-387-57314-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I5 1993. DM58.00.

Mehrotra:1994:HPF

[Meh94] P. Mehrotra. High Performance FORTRAN. In Wieseman [Wie94], pages 546–561. ISSN 0191-7811.

Meissner:1994:EFF

[Mei94] Loren P. Meissner. From the Editor: Fortran Forum schedule change; HPFF 1995 meeting. ACM Fortran Forum, 13 (4):1, December 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:1995:HPF

[Met95] M. Metcalf. High Performance Fortran. In Vandoni [Van95], pages 193–197. ISBN 92-9083-076-X. ISSN 0007-8328. LCCN QC770 .E83 v.95, no. 5.

Metcalf:1999:FH

[Met99a] Michael Metcalf. Fortran 90/95/ HPF. ACM SIGPLAN Notices, 34(12):24–29, December 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Metcalf:1999:FHIa

[Met99b] Michael Metcalf. Fortran 90/ 95/HPF information file (Part 1, compilers). ACM Fortran Forum, 18(1):22–23, April 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:1999:FHIb

[Met99c] Michael Metcalf. Fortran 90/ 95/HPF information file (Part 2, utilities). ACM Fortran Forum, 18(2):26–27, August 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:1999:FHIc

[Met99d] Michael Metcalf. Fortran 90/ 95/HPF information file (Part 3, books in English). ACM Fortran Forum, 18(3):31, December 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2000:FHIb

[Met00a] Michael Metcalf. Fortran 90/ 95/HPF information file (Part 1, compilers). ACM Fortran Forum, 19(2):29–30, August 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2000:FHIc

[Met00b] Michael Metcalf. Fortran 90/ 95/HPF information file (Part 2, tools). ACM Fortran Forum, 19 (3):27–28, December 2000. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2000:FHIa

[Met00c] Michael Metcalf. Fortran 90/ 95/HPF information file (Part 4, courses etc.). ACM Fortran Forum, 19(1):28–29, April 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2001:FHIc

[Met01a] Michael Metcalf. Fortran 90/ 95/HPF information file (Part 1, compilers). ACM Fortran Forum, 20(3):29–30, December 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2001:FHIa

[Met01b] Michael Metcalf. Fortran 90/ 95/HPF information file (Part 3, books). ACM Fortran Forum, 20 (1):27–30, April 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2001:FHIb

[Met01c] Michael Metcalf. Fortran 90/ 95/HPF information file (Part 4, courses, etc.). ACM Fortran Forum, 20(2):10, August 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2002:FHIa

[Met02a] Michael Metcalf. Fortran 90/ 95/HPF information file (part 2, tools, etc.). ACM Fortran Forum, 21(1):30–31, April 2002. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2002:FHIb

[Met02b] Michael Metcalf. Fortran 90/ 95/HPF information file (part 3, books). ACM Fortran Forum, 21 (3):27–29, December 2002. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Merlin:1995:IHP

[MH95] John Merlin and Anthony Hey.

An introduction to High Performance Fortran. Scientific Programming, 4(2):87–113, Summer 1995. CODEN SCIPEV.

ISSN 1058-9244 (print), 1875-919X (electronic).

Mochizuki:1995:WML

[MIN^{+}95] M. Mochizuki, Y. Itoh, M. Nii, T. Takagi, and Y. Mitsui. A wideband monolithic lossy match power amplifier having an LPF/HPF-combined interstage network. IEICE Transactions on Electronics, E78-C(9):1252– 1254, September 1995. CODEN IELEEJ. ISSN 0916-8524.

McShan:1995:AIP

[MKF95] D. L. McShan, M. L. Kessler, and B. A. Fraass. Advanced interactive planning techniques for conformal therapy: high level beam descriptions and volumetric mapping techniques. International Journal of Radiation Oncology, Biology, Physics, 33(5): 1061–1072, December 1, 1995. CODEN IOBPD3. ISSN 0360-3016.

Moore:1994:MPP

[MKS94] M. S. Moore, G. Karsai, and J. Sztipanovits. Model-based programming for parallel image processing. In IEEE [IEE94b], pages 811–815 (vol. 3). ISBN 0-8186-6952-7 (casebound), 0-8186-6950-0 (paperback), 0-8186-6951-9 (microfiche). LCCN TA1637.I25 1994. Three volumes. IEEE catalog no. 94CH35708.

Meadows:1994:MCF

[MM94] L. Meadows and D. Miles. Migrating CM Fortran applications to HPF. In IEEE [IEE94a], pages 37–40. ISBN 0-8186-6965-9. LCCN QA76.58.S95 1994. IEEE catalog no. 95TH8024.

Middleton:1995:EDS

[MMV95] D. Middleton, P. Mehrotra, and J. Van Rosendale. Expressing direct simulation Monte Carlo methods in High Performance Fortran. In Bailey et al. [BBG^{+}95], pages 698– 703. ISBN 0-89871-344-7. LCCN QA76.58.S55 1995.

Meadows:1995:PRS

[MMY95a] L. F. Meadows, D. Miles, and M. Young. Performance results of several High Performance Fortran benchmarks. In IEEE [IEE95a], pages 516–517. ISBN 0-8186-7074-6. ISSN 1063-7133. LCCN QA76.58 .I58 1995.

Miles:1995:PRS

[MMY95b] D. Miles, L. Meadows, and M. Young. Performance results of several High Performance Fortran benchmarks. In IEEE [IEE95a], pages 516–517. ISBN 0-8186-7074-6. ISSN 1063-7133. LCCN QA76.58 .I58 1995.

Mueller:1995:EHP

[MR95] A. Mueller and R. Ruehl. Extending High Performance Fortran for the support of unstructured computations. In ACM [ACM95], pages 127–136. ISBN 0-89791-728-6. LCCN ????

Muller:1996:CDI

[MR96] A. Muller and R. Ruhl. Communicationbuffers for data-parallel, irregular computations. In Szymanski and Sinharoy [SS96], pages 295– 298. ISBN 0-7923-9635-9. LCCN QA76.58.L37 1996.

Mehrotra:1998:HPF

[MVZ98a] P. Mehrotra, J. Van Rosendale, and H. Zima. High Performance Fortran: History, status and future. Parallel Computing, 24 (3):325–354, May 1998. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mehrotra:1998:HPFb

[MVZ98b] P. Mehrotra, J. Van Rosendale, and H. Zima. High Performance Fortran: Status and prospects. Lecture Notes in Computer Science, 1541:345–356, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mehrotra:1998:HPFa

[MVZ98c] Piyush Mehrotra, John Van Rosendale, and Hans Zima. High Performance Fortran: History, status and future. Parallel Computing, 24(3–4):325–354, May 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/cas/tree/
store/parco/sub/1998/24/3-
4/1293.pdf.

Mehrotra:2000:HPF

[MZ00] Piyush Mehrotra and Hans Zima. High Performance Fortran for aerospace applications. NASA contractor report NASA/CR-2000-210321, National Technical Information Service, Washington, DC, USA, 2000. ???? pp. Shipping list number 2001-0362-M.

Mehrotra:2001:HPF

[MZ01] Piyush Mehrotra and Hans Zima. High Performance Fortran for aerospace applications. Parallel Computing, 27(4):477– 501, March 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/35/21/47/28/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/35/21/47/28/30/article.
pdf.

Nakatani:1995:SIHb

[Nak95a] T. Nakatani. Status and issues of HPF compilers. Transactions of the Institute of Electronics, Information and Communication Engineers D-I, J78D-I(2): 142–148, February 1995. CO-DEN DTRDES.

Nakatani:1995:SIHa

[Nak95b] Toshio Nakatani. Status and issues of HPF compilers. Systems and computers in Japan, 26(12): 1–??, ???? 1995. CODEN SC-JAEP. ISSN 0882-1666 (print), 1520-684X (electronic).

Nelson:1996:COR

[Nel96] C. A. Nelson. Compiler optimizations for red-black HPF codes. Lecture Notes in Computer Science, 1184:534–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nicolau:1991:ALC

[Nic91] Alexandru Nicolau. Advances in languages and compilers for parallel processing. Research monographs in parallel and distributed computing. MIT Press, Cambridge, MA, USA, 1991. ISBN 0-262-64028-7. ISSN 0953-7767. 467 pp. LCCN QA76.7 .A38 1991.

Niewels:1994:SDA

[NJ94] J. Niewels and W. Jorden. Systematical development of an autonomous HPF driven and controlled inspection robot. In Chung et al. [CKMU94], pages 324–329. ISBN 1-880653-12-5, 1-880653-10-9. LCCN ????

Nucciarone:1997:NLD

[NOL97] Jeffrey J. Nucciarone, Yusuf Ozyoruk, and Lyle N. Long. New life in dusty decks: Results of porting a CM Fortranbased aeroacoustic model to High Performance Fortran. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

                                        URL http://www.supercomp.
                                        org/sc97/proceedings/TECH/
                                        NUCCIARO/INDEX.HTM. ACM

SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Ning:1994:ADE

[NVG94] Qi Ning, V. Van Dongen, and G. R. Gao. Automatic decomposition in EPPP compiler. In Botsford et al. [BGG^{+}94], pages 283–291. ISBN ???? LCCN ????

Okuda:2002:OEE

[OA02] Hiroshi Okuda and Norihisa Anan. Optimization of elementby-element FEM in HPF 1.1. Concurrency and Computation: Practice and Experience, 14(8–9):647–663, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/95016128/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=95016128{\&}PLACEBO=IE.
pdf.

Offner:1994:DSM

[Off94] C. D. Offner. A data structure for managing parallel operations. In Hesham and Shriver [HS94], pages 33–42. ISBN 0-8186-5060-5. ISSN 1060-3425. LCCN ???? IEEE catalog no. 94TH0607-2.

Offner:1998:PBH

[Off98] Carl D. Offner. Per Brinch Hansen’s concerns about High Performance Fortran. ACM SIGPLAN Notices, 33(8):34– 39, August 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ogino:2002:TDG

[Ogi02] Tatsuki Ogino. Three-dimensional global MHD simulation code for the Earth’s magnetosphere using HPF/JA. Concurrency and Computation: Practice and Experience, 14(8–9):

631–646, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/95016136/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=95016136{\&}PLACEBO=IE.
                                        pdf.

Orlando:1998:CLE

[OP98a] S. Orlando and R. Perego. A coordination layer for exploiting task parallelism with HPF. Lecture Notes in Computer Science, 1511:386–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orlando:1998:MBR

[OP98b] S. Orlando and R. Perego. An MPI-based run-time support to coordinate HPF tasks. Lecture Notes in Computer Science, 1497:289–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orlando:1998:MRS

[OP98c] S. Orlando and R. Perego. An MPI-based run-time support to coordinate HPF tasks. Lecture Notes in Computer Science, 1497:289–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orlando:1999:CRT

[OP99] Salvatore Orlando and Raffaele Perego. COLT_{H}P F, a run-time support for the high-level co-ordination of HPF tasks. Concurrency: Practice and Experience, 11(8):407–434, July 1999. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=63000994;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=63000994&PLACEBO=IE.pdf.

Orlando:2000:MDT

[OPP00] S. Orlando, P. Palmerini, and R. Perego. Mixed data and task parallelism with HPF and PVM. Cluster Computing, 3(3): 201–213, 2000. CODEN ???? ISSN 1386-7857.

Pazat:1996:THP

[Paz96] J.-L Pazat. Tools for High Performance Fortran: A survey. Lecture Notes in Computer Science, 1132:134–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pingali:1995:LCP

[PBG^{+}95] K. Pingali, U. Banerjee, D. Gelernter, A. Nicolau, and D. Padua, editors. Languages and compilers for parallel computing: 7th International Workshop, Ithaca, NY, USA, August 8–10, 1994: proceedings, volume 892 of Lecture notes in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-58868-X. LCCN QA76.58 .W656 1994.

Perrin:1996:DPP

[PD96] Guy-Rene Perrin and Alain Darte. The data parallel programming model: foundations, HPF realization, and scientific applications, volume 1132 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNCSD9. ISBN 3-540-61736-1. ISSN 0302-9743 (print), 1611-3349 (electronic). xv + 284 pp. LCCN QA76.642 .D38 1996.

Ponnusamy:1993:UCD

[PDS^{+}93] R. Ponnusamy, R. Das, J. Saltz, D. Mavriplis, and A. Choudhary. Unstructured computations and the Dybbuk runtime system. In Tentner [Ten93], pages 170–175. ISBN 1-56555-052-8. LCCN ????

Perrin:1994:SSA

[Per94] G.-R. Perrin. Some synthesis aspects for data parallel programming. In Gentzsch and Harms [GH94a], pages 469–474. ISBN 3-540-57980-X (Berlin), 0-387-57980-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1–2 (c1994).

Pais:2004:UHP

[PFS^{+}04] V. A. Pais, N. Fournier, M. A. Sutton, K. J. Weston, and U. Dragosits. Using High Performance Fortran to parallelise a multi-layer atmospheric transport model. Parallel Computing, 30(1):21–33, January 2004. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Parashar:1996:CPP

[PH96a] Manish Parashar and Salim Hariri. Compile-time performance prediction of HPF/ Fortran 90D. IEEE parallel and distributed technology: systems and applications, 4(1):57–??, ???? 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Parashar:1996:CTP

[PH96b] Manish Parashar and Salim Hariri. Compile-time performance prediction of HPF/ Fortran 90D. IEEE parallel and distributed technology: systems and applications, 4(1): 57–73, Spring 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

URL http://dlib.computer.
org/pd/books/pd1996/pdf/p1057.
pdf;   http://www.computer.
org/concurrency/pd1996/p1057abs.
htm.

Ponnusamy:1995:SID

[PHD^{+}95] Ravi Ponnusamy, Yuan-Shin Hwang, Raja Das, Joel H. Saltz, Alok Choudhary, and Geoffrey Fox. Supporting irregular distributions using dataparallel languages. IEEE parallel and distributed technology: systems and applications, 3(1):12–24, Spring 1995. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

URL http://dlib.computer.
org/pd/books/pd1995/pdf/h10012.
pdf;   http://www.computer.

                                        org/concurrency/pd1995/p1012abs.
                                        htm.

Parashar:1994:DAD

[PHHF94a] M. Parashar, S. Hariri, T. Haupt, and G. C. Fox. Design of an application development toolkit for HPF/Fortran 90D. In Kumar [Kum94], pages 482–487. ISBN 0-07-462332-X. LCCN QA 76.58 I587 1994.

Parashar:1994:IPH

[PHHF94b] Manish Parashar, Salim Hariri, Tomasz Haupt, and Geoffrey C. Fox. Interpreting the performance of HPF/Fortran 90D. In IEEE [IEE94e], pages 743–752. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994.

                                        URL http://sc94.ameslab.
                                        gov/AP/contents.html. IEEE

catalog number 94CH34819.

Parashar:1995:DAD

[PHHF95] M. Parashar, S. Hariri, T. Haupt, and G. C. Fox. Design of an application development toolkit for HPF/Fortran 90D. In Prasanna [Pra95], pages 482–487. ISBN 0-07-462332-X. LCCN ????

Ponnusamy:1994:SIDa

[Pon94a]  Ravi Ponnusamy.  Support-

ing irregular distributions in FORTRAN 90D/HPF compilers. Computer science technical report series CS-TR-3268, University of Maryland, College Park, MD, USA, May 1994. 20 pp.

Ponnusamy:1994:SIDb

[Pon94b]  Ravi Ponnusamy.  Support-

ing irregular distributions in FORTRAN 90D/HPF compilers. Computer science technical report series CS-TR-3268.1, University of Maryland, College Park, MD, USA, November 1994. 25 pp.

Palmer:1994:WND

[PPW94] D. W. Palmer, J. F. Prins, and S. Westfold. Work-efficient nested data-parallelism. In IEEE [IEE94a], pages 186–193. ISBN 0-8186-6965-9. LCCN QA76.58.S95 1994. IEEE catalog no. 95TH8024.

Parsons:1994:RRT

[PQ94] R. Parsons and D. Quinlan.

Run-time recognition of task parallelism within the P++ parallel array class library. In IEEE [IEE94d], pages 77–86. ISBN 0-8186-4980-1. LCCN QA76.58.S34 1993.

Prasanna:1995:PPI

[Pra95] V. K. Prasanna, editor. Parallel processing: 1st International workshop — December 1994, Bangalore, India, Parallel Processing — International Workshop. Tata McGraw-Hill Pub. Co., New Delhi, India, 1995. ISBN 0-07-462332-X. LCCN ????

Ponnusamy:1993:DRS

[PSC^{+}93a] R. Ponnusamy, J. Saltz, A. Choudhary, R. Das, and D. Mavriplis. The Dybbuk runtime system.

In IEEE [IEE93a], pages 205– 212. ISBN 0-8186-3400-6. LCCN QA75.5.C58 1993. IEEE catalog no. 93CH3251-6.

Ponnusamy:1993:RCT

[PSC93b] Ravi Ponnusamy, Joel Saltz, and Alok Choudhary. Runtime compilation techniques for data partitioning and communication schedule reuse. In IEEE [IEE93c], pages 361–370. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Ponnusamy:1995:RSC

[PSC^{+}95] R. Ponnusamy, J. Saltz, A. Choudhary, Yuan-Shin Hwang, and G. Fox. Runtime support and compilation methods for user-specified irregular data distributions. IEEE Transactions on Parallel and Distributed Systems, 6(8):815–831, August 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pan:2003:SHI

[PSG03] Yi Pan, Joseph J. S. Shang, and Minyi Guo. A scalable HPF implementation of a finite-volume computational electromagnetics application on a CRAY T3E parallel system. Concurrency and Computation: Practice and Experience, 15(6): 607–621, May 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Perrott:1993:LPD

[PZA93] R. H. Perrott and A. Zarea-Aliabadi. Languages for programming distributed memory systems. Computing and Control Engineering Journal, 45(6): 269–276, December 1993. CO-DEN CCEJEL. ISSN 0956-3385 (print), 1741-0460 (electronic).

Qiang:2000:FIO

[QRH00] J. Qiang, R. D. Ryne, and S. Habib. Fortran implementation of object-oriented design in parallel beam dynamics simulations. Computer Physics Communications, 133(1):18–33, December 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Raghavachari:1995:BRH

[Rag95] Mukund Raghavachari. Book review: The High Performance Fortran Handbook by Charles Koelbel, David Loveman, Robert Schreiber, Guy Steele Jr., and Mary Zosel. ACM SIGPLAN Notices, 30(7):5, July 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rodriguez:1996:POW

[RHH96] Bernardo Rodriguez, Leslie Hart, and Tom Henderson. Parallelizing operational weather forecast models for portable and fast execution. Journal of Parallel and Distributed Computing, 37(2):159–170, September 15, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

idealibrary.com/links/doi/
10.1006/jpdc.1996.0116/production;
 http://www.idealibrary.com/
links/doi/10.1006/jpdc.1996.
0116/production/pdf.

Roth:1997:CSH

[RMCKB97] Gerald Roth, John Mellor-Crummey, Ken Kennedy, and R. Gregg Brickner. Compiling stencils in High Performance Fortran. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

URL http://www.supercomp.
org/sc97/proceedings/TECH/
ROTH/INDEX.HTM. ACM SIGARCH

order number 415972. IEEE Computer Society Press order number RS00160.

Ramaswamy:1997:FET

[RSB97] S. Ramaswamy, S. Sapatnekar, and P. Banerjee. A framework for exploiting task and data parallelism on distributed memory multicomputers. IEEE Transactions on Parallel and Distributed Systems, 8(11):1098–1116, November 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l1098.
pdf;   http://www.computer.
org/tpds/td1997/l1098abs.
htm.

Sabot:1995:HPC

[Sab95] Gary Sabot. High performance computing: problem solving with parallel and vector architectures. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-59830-2. xvi + 246 pp. LCCN QA76.5 .H4817 1995.

Saini:1995:NEP

[Sai95] Subhash Saini. NAS experiences of porting CM Fortran codes to HPF on IBM SP2 and SGI Power Challenge. Technical report NAS-95-010, NASA-CR 203382, National Technical Information Service, Washington, DC, USA, 1995. ?? pp.

Schnabel:1993:WLC

[Sch93] Robert B. Schnabel, editor.

Workshop on Languages, Compilers, and Run-Time Environments for Distributed Memory Multiprocessors: Boulder, Colorado, September 30–October 2, 1992: extended abstracts of talks and poster presentations, volume 28(1) of ACM SIGPLAN Notices. ACM Press, New York, NY 10036, USA, January 1993. CODEN SINODQ. ISBN ???? ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.7 .S54 v.28:1.

Schuster:1994:PPG

[Sch94] Vincent J. Schuster. PGHPF from the Portland Group. IEEE parallel and distributed technology: systems and applications, 2(3):72, Fall 1994. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Schreiber:1996:SIC

[Sch96a] R. Schreiber. Support for irregular computation in High Performance Fortran. Lecture Notes in Computer Science, 1117:285–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schreiber:1996:IH

[Sch96b] R. S. Schreiber. An introduction to HPF. Lecture Notes in Computer Science, 1132:27–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schreiber:1997:HPF

[Sch97] Robert Schreiber. High Performance Fortran, Version 2. Parallel Processing Letters, 7(4): 437–??, December 1997. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Sips:1998:ALE

[SDv98] Henk J. Sips, Will Denissen, and Kees van Reeuwijk. Analysis of local enumeration and storage schemes in HPF. Parallel Computing, 24(3–4):355–382, May 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/parco/sub/1998/24/3-
                                        4/1294.pdf.

Sakagami:2002:PEJ

[SF02] Hitoshi Sakagami and Shingo Furubayashi. Performance evaluation for Japanese HPF compilers with special benchmark suite. Lecture Notes in Computer Science, 2327:491– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/
service/series/0558/bibs/
2327/23270491.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/
2327/23270491.pdf.

Senkov:2013:HPF

[SHZ13] R. A. Sen’kov, M. Horoi, and V. G. Zelevinsky. A highperformance Fortran code to calculate spin- and paritydependent nuclear level densities. Computer Physics Communications, 184(1):215– 221, January 2013. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465512002925.

Shindo:1995:HCA

[SIDH95] T. Shindo, H. Iwashita, T. Doi, and J. Hagiwara. HPF compiler for the AP1000. In ACM [ACM95], pages 190–194. ISBN 0-89791-728-6. LCCN ????

Siegal:1994:PEI

[Sie94a] Howard Jay Siegal, editor. Proceedings / Eighth International Parallel Processing Symposium, April 26–29, 1994, Cancun, Mexico. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5602-6. LCCN QA76.58.I58 1994. IEEE catalog no. 94CH34819.

Siegel:1994:PEI

[Sie94b] Howard Jay Siegel, editor. Proceedings / Eighth International Parallel Processing Symposium, April 26–29, 1994, Cancun, Mexico. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5602-6, 0-8186-5601-8. ISSN 1063-7133. LCCN QA 76.58 I56 1994.

Seo:2002:HJE

[SIOS02] Yoshiki Seo, Hidetoshi Iwashita, Hiroshi Ohta, and Hitoshi Sakagami. HPF/JA: extensions of High Performance Fortran for accelerating real-world applications. Concurrency and Computation: Practice and Experience, 14(8–9):555–573, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract/95016127/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=95016127{\&}PLACEBO=IE.
                                        pdf.

Sakagami:2002:CCP

[SM02a] H. Sakagami and T. Mizuno. Compatibility comparison and performance evaluation for Japanese HPF compilers using scientific applications. Concurrency and Computation: Practice and Experience, 14(8–9):

679–689, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/95016123/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=95016123{\&}PLACEBO=IE.
pdf.

Shires:2002:EHM

[SM02b] D. Shires and R. Mohan. An evaluation of HPF and MPI approaches and performance in unstructured finite element simulations. Journal of Mathematical Modelling and Algorithms, 1 (3):153–167, 2002. CODEN ???? ISSN 1570-1166.

Smith:1995:WHA

[Smi95] Brian T. Smith. Workshop:

HPF: A user’s perspective. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/MISC/WORKSHOP.

HTM#HPF. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Sakagami:2002:TTD

[SMSY02] Hitoshi Sakagami, Hitoshi Murai, Yoshiki Seo, and Mitsuo Yokokawa. 14.9 TFLOPS three-dimensional fluid simulation for fusion science with HPF on the Earth Simulator. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN

                                        ????  URL http://www.
                                        sc-2002.org/paperpdfs/pap.
                                        pap147.pdf.

Sumana:1994:PRA

[SN94] S. Sumana and U. Nagaraj Shenoy. Porting realistic applications to HPF — A case study. In Kumar [Kum94], pages 370–375. ISBN 0-07-462332-X. LCCN QA 76.58 I587 1994.

Sumana:1995:PRA

[SN95] S. Sumana and U. Nagaraj Shenoy. Porting realistic applications to HPF — A case study. In Prasanna [Pra95], pages 370–375. ISBN 0-07-462332-X. LCCN ????

Snir:1992:PI

[Sni92] Marc Snir. Proposal for IO. Posted to HPFF I/O Forum, August 31, 1992. Second Draft.

Satoh:2006:ATL

[SNK06] Makoto Satoh, Kiyoshi Negishi, and Atsushi Kobayashi. Analysis of two-level data mapping in an HPF compiler for distributed-memory machines. Parallel Computing, 32(4):280– 300, April 2006. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schuster:1993:RHP

[SNMC93] V. J. Schuster, S. Nakamoto, L. Meadows, and A. Choudhary. Retargetable High Performance Fortran compilers. In Anonymous [Ano93l], pages 111–123. ISBN ???? LCCN ????

Stichnoth:1994:GCA

[SOG94] James M. Stichnoth, David O’Hallaron, and Thomas R. Gross. Generating communication for array statements: Design, implementation, and evaluation. Journal of Parallel and Distributed Computing, 21 (1):150–159, April 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic).  URL http://www.
idealibrary.com/links/doi/
10.1006/jpdc.1994.1048/production;
 http://www.idealibrary.com/
links/doi/10.1006/jpdc.1994.
1048/production/pdf.

Sharma:1994:RCS

[SPM^{+}94] S. D. Sharma, R. Ponnusamy, B. Moon, Yuan-Shin Hwang, R. Das, and J. Saltz. Runtime and compile-time support for adaptive irregular problems. In IEEE [IEE94e], pages 97–106. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994.

URL http://sc94.ameslab.
gov/AP/contents.html. IEEE

catalog number 94CH34819.

Spoerl:1994:SHC

[Spo94] James R. Spoerl. SofTech’s HPF compiler. IEEE parallel and distributed technology: systems and applications, 2(3):75, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Sivaraman:1995:PSP

[SR95] H. Sivaraman and C. S. Raghavendra. Parallelizing sequential programs to a cluster of workstations. In Agrawal [Agr95], pages 38–41. ISBN 0-8493-2618-4. LCCN QA76.58.I34 1995.

Sivaraman:1996:AAD

[SR96] H. Sivaraman and C. S. Raghavendra. ADDT: Automatic data distribution tool for porting programs to PVM. In El-Rewini and Shriver [ERS96], pages 557–564. ISBN 0-8186-7324-9. ISSN 1060-3425. LCCN ???? Five volumes.

Souravlas:2004:PTD

[SR04] Stavros Souravlas and Manos Roumeliotis. A pipeline technique for dynamic data transfer on a multiprocessor Grid. International Journal of Parallel Programming, 32(5):361– 388, October 2004. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/openurl.
                                        asp?genre=article&issn=0885-
                                        7458&volume=32&issue=5&spage=
                                        361.

Szymanski:1996:LCR

[SS96] Boleslaw K. Szymanski and Balaram Sinharoy, editors. Languages, Compilers and Run-Time Systems for Scalable Computers, 22–24 May 1995, Troy, NY, USA. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 0-7923-9635-9. LCCN QA76.58.L37 1996.

Sturler:1997:SPH

[SS97] E. De Sturler and V. Strumpen.

Scientific programming with High Performance Fortran: A case study using the xHPF compiler. Scientific Programming, 6(1):127–152, 1997. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Subhlok:2000:APM

[SS00] J. Subhlok and P. Steenkiste.

Airshed pollution modeling in an HPF style environment. Journal of Parallel and Distributed Computing, 60(6):690–715, June 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shih:2000:EAG

[SSC00] K-P. Shih, J-P. Sheu, and C-Y.

Chang. Efficient address generation for affine subscripts in dataparallel programs. The Journal of Supercomputing, 17(2):205– 227, January 2000. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Suzuoka:1994:PDB

[SSG94] T. Suzuoka, J. Subhlok, and T. Gross. Performance debugging based on scalability analysis. In IEEE [IEE94a], pages 406–413. ISBN 0-8186-6965-9. LCCN QA76.58.S95 1994. IEEE catalog no. 95TH8024.

Suzuoka:1997:PDT

[SSG97] Takashi Suzuoka, Jaspal Subhlok, and Thomas Gross. A performance debugging tool for high performance Fortran programs. Concurrency: Practice and Experience, 9(10):927–945, October 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13809;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13809&PLACEBO=IE.pdf.

Stanford:1994:PQC

[Sta94] N. Stanford. Portable QCD codes for massively parallel processors. In Anonymous [Ano94g], pages 817–819. CO-DEN NPBSE7. ISBN ???? ISSN 0920-5632 (print), 1873-3832 (electronic). LCCN ????

Steele:1993:HPF

[Ste93] Guy L. Steele, Jr. High Performance Fortran: status report. In Schnabel [Sch93], pages 1–4. CODEN SINODQ. ISBN ???? ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.7 .S54 v.28:1.

Stewart:1995:RAD

[Ste95] A. Stewart. Reasoning about data-parallel array assignment. Journal of Parallel and Distributed Computing, 27(1):

79–85, May 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic).  URL http://www.
idealibrary.com/links/doi/
10.1006/jpdc.1995.1073/production;
 http://www.idealibrary.com/
links/doi/10.1006/jpdc.1995.
1073/production/pdf.

Strok:1994:NJI

[Str94] Dale C. Strok. In the news:

Jupiter impacts: Resolution makes a big difference. supercomputer farming down under. HPF Forum welcomes comments. Smithsonian Awards honor computational scientists. low-life computer viruses. PVM developers get R&D-100 award. the eyes have it. neural nets detect breast cancer. better cars through cooperation. parallel version of global climate model. Lockheed to run Idaho National Engineering Lab. public-private partners: new drugs, new software. IEEE Computational Science & Engineering, 1(3):88–90, Fall 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Sullivan:1988:RHU

[Sul88] Kevin M. (Kevin Maurice) Sullivan. The redefinition of HP-FORTRAN and its use in the translation from VAX-11 FOR-TRAN to Microsoft FORTRAN

77. Thesis (m.s.), Auburn University, Auburn, AL, USA, 1988. vii + 67 pp.

Subhlok:1995:OMS

[SV95] Jaspal Subhlok and Gary Vondran. Optimal mapping of sequences of data parallel tasks. In Anonymous [Ano95b], pages 134–143. CODEN SINODQ. ISBN ???? ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN ????

Sips:1996:ALE

[SVD96] H. J. Sips, K. Van Reeuwijk, and W. Denissen. Analysis of local enumeration and storage schemes in HPF. In ACM [ACM96a], pages 10–17. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Schulz-Ziemer:1995:HIP

[SZG95] G. Schulz-Ziemer and A. Geiger. HPF on Intel Paragon and CRAFT on CRAY T3D: basic performance measurements and experiments with a block-sparse CG-algorithm. In Hertzberger and Serazzi [HS95], pages 618– 625. ISBN 3-540-59393-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Sarma:1998:UHP

[SZM98] G. Sarma, T. Zacharia, and D. Miles. Using High Performance Fortran for parallel programming. Computers and Mathematics with Applications, 35(12):41–57, June 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.

sciencedirect.com/science/
article/pii/S0898122198000959.

Thakur:1994:COD

[TBC94a] R. Thakur, R. Bordawekar, and A. Choudhary. Compilation of out-of-core data, parallel programs for distributed memory machines. ACM SIGARCH Computer Architecture News, 22 (4):23–28, September 1994. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Thakur:1994:CRS

[TBC94b] R. Thakur, R. Bordawekar, and A. Choudhury (or Choudhary??). Compiler and runtime support for out-of-core HPF programs. In ACM [ACM94], pages 382–391. ISBN 0-89791-665-4. LCCN ????

Thakur:1994:RAR

[TCF94] R. Thakur, A. Choudhary, and G. Fox. Runtime array redistribution in HPF programs. In IEEE [IEE94c], pages 309– 316. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog no. 94TH0637-9.

Thakur:1996:EAA

[TCR96] Rajeev Thakur, Alok Choudhary, and J. Ramanujam. Efficient algorithms for array redistribution. IEEE Transactions on Parallel and Distributed Systems, 7(6):587–594, June 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic).  URL http://www.

                                        computer.org/tpds/td1996/
                                        l0587abs.htm.

Tentner:1993:HPC

[Ten93] Adrien Michel Tentner, editor. High Performance Computing Symposium 1993. Grand Challenges in Computer Simulation. Proceedings of the 1993 Simulation Multiconference on the High Performance Computing Symposium. SCS, San Diego, CA, USA, 1993. ISBN 1-56555-052-

8. LCCN ????

Thole:1993:HPF

[Tho93] C. A. Thole. High Performance Fortran. In Anonymous [Ano93j], pages 885–892 (or 885– 891??). ISBN ???? ISSN 0254-6213. LCCN ????

Thirumalai:1996:ECA

[TR96] Ashwath Thirumalai and J. Ramanujam. Efficient computation of address sequences in data parallel programs using closed forms for basis vectors. Journal of Parallel and Distributed Computing, 38(2):188– 203, November 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

                                        tronic).  URL http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jpdc.1996.0140/production;
                                         http://www.idealibrary.com/
                                        links/doi/10.1006/jpdc.1996.
                                        0140/production/pdf.

Thirumalai:1996:CGO

[TRV96] A. Thirumalai, J. Ramanujam, and A. Venkatachar. Communication generation and optimization for HPF. In Szymanski and Sinharoy [SS96], pages 311– 316. ISBN 0-7923-9635-9. LCCN QA76.58.L37 1996.

Ulberg:1994:BRC

[UMM94] Dima E. Ulberg, A. John Mallinckrodt, and Susan McKay. Book review: Charles H. Koelbel, David B. Loveman, Robert S. Schreiber, Guy L. Steele, Jr., and Mary E. Zosel, The High Performance Fortran Handbook. Computers in Physics, 8(4): 428–??, July 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.
org/doi/10.1063/1.4823319.

Ujaldon:1995:NDL

[UZCZ95] M. Ujaldon, E. L. Zapata, B. M.

Chapman, and H. P. Zima. New data-parallel language features for sparse matrix computations. In IEEE [IEE95a], pages 742– 749. ISBN 0-8186-7074-6. ISSN 1063-7133. LCCN QA76.58 .I58 1995.

Ujaldon:1996:DLF

[UZCZ96] M. Ujaldon, E. L. Zapata, B. M.

Chapman, and H. P. Zima. Data-parallel language features for sparse codes. In Szymanski and Sinharoy [SS96], pages 253– 264. ISBN 0-7923-9635-9. LCCN QA76.58.L37 1996.

Ujaldon:1997:VFH

[UZCZ97] M. Ujaldon, E. L. Zapata, B. M. Chapman, and H. P. Zima. Vienna-Fortran/HPF extensions for sparse and irregular problems and their compilation. IEEE Transactions on Parallel and Distributed Systems, 8(10):1068–1083, October 1997. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://dlib.computer.org/
td/books/td1997/pdf/l1068.
pdf;   http://www.computer.
org/tpds/td1997/l1068abs.
htm.

vanWaveren:1994:HPF

[van94a] G. M. van Waveren. High Performance Fortran. In Gentzsch and Harms [GH94a], pages 429–433. ISBN 3-540-57980-X (Berlin), 0-387-57980-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1–2 (c1994).

Vanderlip:1994:PSV

[Van94b] Jeff Vanderlip. Pacific Sierra’s VAST-HPF and VAST/77toHPF. IEEE parallel and distributed technology: systems and applications, 2(3):74, Fall 1994. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Vandoni:1995:SCA

[Van95] C. E. Vandoni, editor. 1995 CERN School of Computing: Arles, France, 20 August – 2 September 1995: proceedings, number 5 in CERN European Organization for Nuclear Research — Reports. CERN, Geneva, Switzerland, 1995. ISBN 92-9083-076-X. ISSN 0007-8328. LCCN QC770 .E83 v.95, no. 5.

VanderPloeg:1998:PBT

[Van98] A. Van der Ploeg. Parallelization of a block tridiagonal solver in HPF on an IBM SP2. Lecture Notes in Computer Science, 1401:242–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanReeuwijk:1996:IFH

[vDSP96] Kees van Reeuwijk, Will Denissen, Henk J. Sips, and Edwin M. R. M. Paalvast. An implementation framework for HPF distributed arrays on messagepassing parallel computer systems. IEEE Transactions on Parallel and Distributed Systems, 7(9):897–914, September 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://www.

computer.org/tpds/td1996/
l0897abs.htm.

Veen:1994:PHP

[Vee94] Arthur H. Veen. The Prepare HPF programming environment. IEEE parallel and distributed technology: systems and applications, 2(3):73, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

vonHanxleden:2000:BCP

[vHK00] Reinhard von Hanxleden and Ken Kennedy. A balanced code placement framework.

ACM Transactions on Programming Languages and Systems, 22(5):816–860, 2000. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

www.acm.org/pubs/articles/
journals/toplas/2000-22-5/
p816-von_hanxleden/p816-von_
hanxleden.pdf;   http://www.
acm.org/pubs/citations/journals/
toplas/2000-22-5/p816-von_
hanxleden/.

vonHanxleden:1994:VDA

[vHKS94] R. von Hanxleden, K. Kennedy, and J. Saltz. Value-based distributions and alignments in Fortran D. Journal of Programming Languages, 2(3):259– 282, September 1994. CODEN JPLAER. ISSN 0963-9306.

Volkert:1993:PCS

[Vol93]  Jens Volkert, editor.  Par-

allel computation: Second International ACPC Conference, Gmunden, Austria, October 4– 6, 1993: proceedings, volume 734 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-57314-3, 0-387-57314-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I5 1993. DM58.00.

Venkatachar:1997:CGB

[VRT97] A. Venkatachar, J. Ramanujam, and A. Thirumalai. Communication generation for blockcyclic distributions. Parallel Processing Letters, 7(2):195– 202, June 1997. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

vanWaveren:2002:CGH

[vWAH^{+}02] Matthijs van Waveren, Cliff Addison, Peter Harrison, Dave Orange, Norman Brown, and Hidetoshi Iwashita. Code generator for the HPF Library and Fortran 95 transformational functions. Concurrency and Computation: Practice and Experience, 14(8–9): 589–602, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/95016135/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=95016135{\&}PLACEBO=IE.
pdf.

Wagenbreth:1994:AAH

[Wag94] G. Wagenbreth. APR’s approach to High Performance Fortran for distributed memory multiprocessor systems. In IEEE [IEE94a], pages 41–45. ISBN 0-8186-6965-9. LCCN QA76.58.S95 1994. IEEE catalog no. 95TH8024.

Wu:1999:CFF

[WCC99] Jan-Jan Wu, Marina Chen, and James Cowie. CRAFT: a framework for F90/HPF compiler optimizations. Concurrency: Practice and Experience, 11(10):529–569, August 25, 1999. CODEN CPEXEI. ISSN 1040-3108. URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=63003393;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=63003393&PLACEBO=IE.pdf.

Wolfe:1994:AAA

[WI94] M. Wolfe and M. Ikei. Automatic array alignment for distributed memory multicomputers. In Hesham and Shriver [HS94], pages 23–32. ISBN 0-8186-5060-5. ISSN 1060-3425. LCCN ???? IEEE catalog no. 94TH0607-2.

Wieseman:1994:RCR

[Wie94] C. D. Wieseman, editor. Role of computers in research and development at Langley Research Center: Workshop entitled “The role of computers in LARC R and D” — June 1994, Hampton, VA, number 10159 in NASA Conference Publication. National Aeronautics and Space Administration, Washington, DC, USA, 1994. ISSN 0191-7811.

Walker:1996:RBC

[WO96] D. W. Walker and S. W. Otto. Redistribution of block-cyclic data distributions using MPI. Concurrency: Practice and Experience, 8(9):707–728, November 1996. CODEN CPEXEI. ISSN 1040-3108. URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=23305.

Wholey:1994:TMP

[WSL94] Skef Wholey, Richard Shapiro, and David Loshin. Thinking Machines’ plans for HPF. IEEE parallel and distributed technology: systems and applications, 2(3):76, Fall 1994. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Wakatani:1994:NAA

[WW94] A. Wakatani and M. Wolfe. A new approach to array redistribution: strip mining redistribution. In Halatsis et al. [HMPT94], pages 323–335. ISBN 3-540-58184-7. LCCN QA76.58 .I564 1994.

Wakatani:1995:OAR

[WW95] Akiyoshi Wakatani and Michael Wolfe. Optimization of array redistribution for distributed memory multicomputers. Parallel Computing, 21(9):1485– 1490, September 12, 1995. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/parco/cas_sub/
browse/browse.cgi?year=1995&
volume=21&issue=9&aid=1006.

Yau:1997:EHP

[YFH97] H. W. Yau, G. C. Fox, and K. A.

Hawick. Evaluation of High Performance Fortran through application kernels. Lecture Notes in Computer Science, 1225:772– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1994:HPF

[YGS^{+}94] S. X. Yang, D. Gannon, S. Srinivas, F. Bodin, and P. Bode. High Performance Fortran interface to the parallel C++. In IEEE [IEE94c], pages 301– 308. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog no. 94TH0637-9.

Yang:1995:PCM

[YO95] Bwolen Yang and D. R. O’Hallaron. Procedure call models for distributed parameters in data parallel programs. In IEEE [IEE95b], pages 157–164. ISBN 0-8186-6895-4. LCCN QA76.58 .S34 1994.

Yang:1994:DIP

[YWS^{+}94] B. Yang, J. Webb, J. M. Stichnoth, D. R. O’Hallaron, and T. Gross. Do and Merge: integrating parallel loops and reductions. In Banerjee et al. [BGNP94], pages 169–183. ISBN 3-540-57659-2. LCCN QA76.58 .W656 1993.

Zarea-Aliabadi:1993:LPD

[ZA93]  Adib Zarea-Aliabadi.  Lan-

guages for programming distributed memory systems. Computing and Control Engineering Journal, 4(6):269–??, December 1, 1993. CODEN CCEJEL. ISSN 0956-3385 (print), 1741-0460 (electronic).

Zima:1994:SVF

[ZBC94] H. P. Zima, P. Brezany, and B. M. Chapman. SUPERB and Vienna Fortran. Parallel Computing, 20(10–11):1487– 1517, November 3, 1994. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/parco/cas_sub/
browse/browse.cgi?year=1994&
volume=20&issue=10-11&aid=
909.

Zhang:1998:PBH

[ZCFL98] G. Zhang, B. Carpenter, G. Fox, and X. Li. PCRC-based HPF compilation. Lecture Notes in Computer Science, 1366:204– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zima:1995:CTS

[ZCP95] H. P. Zima, B. M. Chapman, and M. Pantan. Compiler technology for scalable parallel architectures — a short overview. In Hoffmann and Kreitz [HK95], pages 240–249. ISBN 981-02-2211-4. LCCN QC866.E26 1994.

Zima:1999:IHP

[Zim99] H. P. Zima. An introduction to HPF+ project. Lecture Notes in Computer Science, 1615:9– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zima:2002:HPF

[Zim02] Hans P. Zima. High Performance Fortran — history, status and future (HiWEP invited talk). Lecture Notes in Computer Science, 2327:490– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/series/0558/bibs/
                                        2327/23270490.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/series/0558/papers/
                                        2327/23270490.pdf.

Zosel:1993:HPF

[Zos93] M. E. Zosel. High performance Fortran: an overview. In IEEE [IEE93a], pages 132– 136. ISBN 0-8186-3400-6. LCCN QA75.5.C58 1993. IEEE catalog no. 93CH3251-6.